A Troop, 3rd Squadron, 4th Cavalry 25th ID - Vietnam

Counteroffensive Phase VI, 2 Nov 1968 - 22 Feb 1969
Last Updated on: Tuesday, February 17, 2015 08:08:04 PM

    We are building timelines for the Troop's service in Vietnam. If you have award citations, SITREP's, AAR's or other documents, they can provide dates that can add to the timelines.  Mel Moss, Trp Cdr from Sep. 68 to Feb. 69, is our volunteer coordinator for the Phase V timeline.
     Note 1: Entries that include an exact time or grid coordinates have been extracted from offical documents.  If no time is provided it is a recollection from a unit member.
     Note 2 Email us and provide additions or corrections for this timeline.

Nov 1
Summary: Outpost & secure MSR operating from Tay Ninh base camp.
0553    A Trp will outpost 1st plt from CP03 – CP022
0628    A Trp departing Main Gate of Tay Ninh
0724    A Trp reports that the tanker (which was at CP02) which was laying on its side in                       
the rice paddy is now back on the road.
0740    A Trp, foad clear in sector, beginning sweep of rubber, MSR clear  Hoc Mon to Tay Ninh                http://home.comcast.net/~ATrp3-4Cav/Photos/index.html
0937    Sqdn CO inroute to B Trp’s location, then to A Trp’s, location then on to Tay Ninh.                       
1351    A Trp elements at CP23
1403    A Trp elements at CP02
1408    A Trp elements at CP23
1420    A Trp elements at CP01
1435    A Trp element trail at CP31
1437    A Trp  lead at CP86
1446    A Trp at CP87                                   
1453    A Trp elements closed Tay Ninh
1745    Ref to log 165, relayed to A Trp
1822    A & B Trps security checks neg;                    
Nov 2
0055    3-4 Cav sitrep no contacthttp://home.comcast.net/~ATrp3-4Cav/Photos/index.html
0548    A Trp entering net
0632    A Trp sweep at CP87
0648    A Trp at CP86
0705    A Trp approaching CP04
0708    A Trp at CP24
0728    A Trp at CP23
0731    Convoy crossed bridge
0740    A Trp sector clear
0745    A Trp outpost:  A30 from CP22 – CP03, A10  from CP03 - CP23
0808    A Trp has cleared & outposted rubber
0924    A Trp Reports an Engr Convoy of 10 veh coming south on MSR, presently at CP03                        
1320    A Trp, 3rd plt rolling up, securing Engrs at CP23
1338    A trp at CP02
1347    A Trp at CP24
1405    A Trp at CP31
1417    A Trp at CP87
1419    A Trp closed Tay Ninh, wpns cleared, all personnel accounted for
1426    Lead at main gate and trail at CP09
1726    A Trp left net
Nov 3
0015    A Trp neg sitrep
0637    A Trp outpost: A30 from23 – CP03 & A20 from CP03 – CP22
0644    A Trp at CP86
0649    A Trp at CP31
0704    A Trp at CP01
0709    A Trp at CP24
0719    A Trp elements at CP02
0730    A Trp elements at CP23
0735    A Trp at CP03
0743    A Trp sweep at CP22, MSR is open
1323    Maj Jordan is enroute to A Trp’s position then going to their night lager
1336    Maj Jordan will arrive at A Trp’s position in 2 min. if anyone wants him he will be at their position
1355    SPOTREP A Trp 3rd platoon Booby Trap at XT316355, 1 US Grenade & trip flare, cont mission, will 
          blow in place
1358    Maj Jordan airborne to Tay Ninh at this time
1405    A Trp reports 4 MU in vic of CP01
1450    A Trp MU 3 is at eastern portion of big rubber, two wreckers, 1MP Jeep, also consist of trail, there
          is a quad 50 downed at CP23.  3 bolts on right front wheel of Veh, request someone come down
          from Tay Ninh to tow the downed veh.
1458    A Trp reports there is a wrecker there to tow the downed veh at CP23 back to Tay Ninh
1531    A Trp lead at CP02
1542    A Trp lead at CP24
1600    A Trp lead at CP86, trail at CP31
1628    A Trp all elements have entered main gate Tay Ninh at this time.
Nov 4
0144    A Trp Sitrep N/C
0606    A Trp entered net
0631    A Trp cleared main gate at Tay Ninh
0632    A Trp 10 will outpost from CP03 – Cp22 and 20 from CP23 – CP03
0648    A Trp lead at CP86
0710    A Trp lead at CP24
0722    A Trp lead at CP02
0734    A Trp lead at CP23
0738    A Trp lead at CP03
0751    A Trp  reports road swept to CP22
1328    Maj Jordan, Sqdn XO airborne at this time to check out the Aerorifles and then on to check on the
          rest of the 3-4 Cav elements
1340    CO A Trp given permission to roll up as soon as convoy passes, Engrs are working at CP23, secured
          by the PHILCAV, Bailly Bridge is now open and in use (CP23)
1344    Trail passing eastern edge of big rubber, A Trp outposts rolling up.
1359    A Trp located at CP23
1445    A Trp lead at CP 31
1446    A Trp lead at CP 86
1457    A Trp lead at CP87
1500    A Trp closing Tay Ninh, all wpns cleared and elevated, all personnel accounted for
1312    Maj Jordan, Sqd XO, departing Tay Ninh for Cu Chi
1655    A Trp leaving net until tomorrow morning
1743    A & B Trps internal security ok
Nov 5
0604    A Trp entered the net.
0626    A Trp beginning sweep:  A Trp at CP87, outposts will be: 3rd platoon from CP3 to CP22, 1st platoon
          from CP3 to CP23
0650    A Trp lead at CP31
0906    A Trp reports an ARVN  convoy, vic big rubber moving east on MSR, approx35-40 veh with a few US
          veh w/them
1020    SQDN XO (MAJ Jordan) airborne enroute to A Trp CO’s Location
1131    SQDN XO off Tay Ninh enroute to Cu Chi
1436    Late entry:  A Trp rolling up, now located at CP23
1445    A Trp at CP02,  Air strikes are continuing at XT2833
1538    A Trp closed Tay Ninh
1743    A Trp Checked area, cleared of all personnel
2100    All 3-4 Cav Sitrep N/C
Nov 6
0052    3-4 Cav Sitrep N/C
0158    DISCOM: reports believed incoming mortars & rockets on Cu Chi Base Camp, informed to get all
          personnel into bunkers, all Trps notified (HQ, A, B, C, D, & LRPs
0225    To DISCOM & 554 Engrs. Neg mortar rds impacted in 3-4 Cav area, neg casualties and neg damage
0231    CPT Hallanck: stand down complete at this time, all Trps notified, (HQ, A, B, C, D, & LRRPs)
0251    3-4 Cav Sitrep N/C
0615    A Trp enroute to Main Gate Tay Ninh
0630    A Trp at CP86
0647    A Trp at CP31
0703    A Trp at CP01http://home.comcast.net/~ATrp3-4Cav/Photos/index.html
0717    A Trp at CP02
0737    A Trp at CP03
0759    A Trp Sector clear
0802    ETA of convoy at Hoc Mon, 0830H acknowledged by A,B,C Trps
1436    CO A Trp: trail of convoy at CP03, MPs inform him that there were not any stragglers.
1451    Trail at CP22
1518    A Trp at CP31
1544    A Trp lead at CP87
1548    A Trp closed gate at Tay Ninh
1653    A Trp leaving net until tomorrow
1707    A Trp Security check neg results
Nov 7
0603    A Trp entered the net
0625    A Trp started sweep outpost:  A10 from CP22- CP03, A20 from CP03 – CP23
0626    A Trp at CP87
0648    A Trp at CP31
0711    A Trp lead at CP24
0720    A Trp lead at CP02
0732    A Trp lead at CP23
0739    A Trp lead at CP03
0753    A Trp swept to CP22
0801    A Trp reports that ARVN convoy at CP03, heading east consists of APCs
0813    Visual reconnaissance on the way
1042    MAJ Jordan airborne to A Trp’s location in Tay Ninh
1231    A Trp: last MU at CP06, trail at eastern edge of Big Rubber
1430    CO A Trp: lead MP at CP03, also an ARVN convoy directly behind the lead MP
1431    Convoy lead at CP01
1550    Trail cleared by A Trp’s area.  A Trp outposts rolling up at this time
1620    A Trp trail at CP02
1622    A Trp lead at CP24
1625    A Trp lead at CP01
1647    A Trp lead at CP86
1707    A Trp closed Tay Ninh
1742    A Trp internal security ok
1855    Sp4 Niles, Cu Chi Subsector – no aval – VC will try to launch a winter-spring offensive 68-69 end of
           Nov attacking Inf Trps in force
2300    3-4 Cav sitrep N/C
Nov 8
0559    A Trp entered Net
0632    A Trp  lead at CP87
0645    A Trp lead at CP86
0652    A Trp leat at CP31
0736    A Trp outposting status: CP03 – CP22 3rd platoon, CP03 – CP23 1st platoon
0746    A Trp sector clear; MSR cleared and outposted in ¾ Cav area
0820    CPT Kitchin (Maint Officer) going out on MSR today w/1 APC to check on tanks & maintenance
0959    A Trp reports hearing some shots, believed to be ARVN test firing, will check
1443    Trail at CP03, A Trp will roll up when convoy clears area
1507    A Trp at CP02
1521    A Trp at CP24
1526    A Trp at CP01
1527    A Trp to leave sweep team at Tay NinhThank you for all you have done on this project.
1543    A Trp at CP31
1545    A Trp At CP86
1705    A Trp closed Tay Ninh approx 15 min ago
2123    A Trp security check: all clear
Nov 9
Summary: Escort division resupply convoys, clear & outpost MSR, & strongpoint MSR at night.
0620    A Trp entered the net
0628    A Trp road swept to CP08
0655    Convoy ETA Phu Cuong 0740H; 3MU for Dau Tieng, 8MU for Tay Ninh
0744    A Trp lead at CP03, trains and HQ elements at CP23
0745    A Trp sector clear; MSR open and outposted
0752    A Trp status: 1st plt from CP03-CP22,3rd Plt from CP03 to CP23, 2nd Plt will become CPCON to C
          Trp, and OP from CP09-CP11
0756    Convoy now located at Phu Cuong Bridge
0759    A Trp 2nd plt lead at CP22
0806    A Trp 2nd plt lead at CP04
0814    A Trp 2nd plt at CP21
0821    A Trp 2nd plt at CP05
0827    A Trp  lead at CP19
0842    A Trp 2nd plt at CP19
0905    A Trp 1st plt at CP08
0948    Trail of convoy at CP09
1006    Lead of Convoy at CP05
1010    A Trp 20 at CP09
1018    A Trp 20 going up on C Trp push, will be under C Trp control for the rest of today
1024    A Trp HQ elements cleared Cu Chi main gate, all wpns cleared & elevated, personnel accounted for
1024    Lead of Convoy at CP21, trail at CP08
1027    Lead of Convoy at CP04
1054    A Trp 2nd Plt now outposted crom CP09-CP11
1100    Lead of Convoy at CP02 trail at CP20
1413    Convoy closed Tay Ninh at 1315H
1444    A Trp reports a downed veh at CP03,  A Trp will leave security and begin rolling up
1500    A Trp at CP22
1513    A Trp at CP04
1519    A Trp at CP21
1523    A Trp reports trail of trail at CP22
1526    A Trp at CP05
1534    A Trp trail of trail at CP22
1536    A Trp reports lowboy at CP20
1543    A Trp at CP20
1548    Trail Party at CP06
1553    Trail Party at CP20
1601    DISCOM has people taking care of downed lowboy at CP20, will pull it into FSB Stewart
1604    Trail party at CP07
1610    A Trp: Trail of trail at CP07
1626    A Trp at CP17, starting to outpost at An Duc
1630    Trail at CP16
1646    A10 outposted from An Duc to CP09
1656    Convoy closed CP16
1704    A Trp line up north to south 30-10-20
1816    A Trp 1st & 3rd plts moving to night laagers
1828    Col request: LOH to fly cover for COL Riley and General Long visit to outposting elements
1834    A Trp 1st & 3rd plts in night laagers
1839    Lt Gillard gave A Trp permission to test fire at 1930H
1843    A Trp security check clear
1848    A Trp vehicle status rept: 24 tracks, and 9 tracks operational, 3 downed tracks
1920    A Trp 2nd plt in night lager
1928    A Trp 1st & 3rd plts test firing at this time
1946    A Trp 3rd plt AP Departing for night laager
1949    A Trp 1st plt test fire complete
1955    A Trp 3rd plt radar report: L 153 deg, R 268 deg, range L1000m,  R 10m, C3500m
2004    A Trp 1st & 3rd plts AP’s in position
2010    A Trp 1st plt sniper team departing for position
2021    A Trp 2nd plt to test fire on a 25deg azimuth
2033    ¾ Cav sitrep N/C   
2035    A Trp 3rd plt sniper team in position
2039    A Trp 3rd plt sniper team and AP’s in position
2106    A Trp: All elements in position
2140    A Trp Neg Sitrep
2359    Change and check Secure
Nov 10
0026    A Trp and LRP teams neg Sitrep
0620    A Trp 3rd plt on move to northwest
0628    A Trp 2nd plt on move to south, will outpost from CP09-CP11,  A Trp 1st plt from night laager to
          CP17 then into base-camp for stand down; 3rd plt outpost from northern boundary to CP09
0639    A Trp 20 in VP16
0645    A Trp 10 at CP09, A30 at An Duc
0700    SPOTREP A Trp 30 2 mines at XT531183, 30lb mines, pressure type, homemade, 1 on north and 1
          south side of road, blown in place, crater 4’x3’ / LE cont mission
0715    A Trp clear CP11 to CP17, from CP08 to An Duc clear
0730    A Trp sector clear and outposted
0743    A Trp reports he sent an element (1st Plt) in Ambush Gate.  He was not informed it was closed to
          cav vehicles, The MP’s gave the entire platoon a DR.
0750    Convoy from Cu Chi: 12 MU to Go Dau Ha, 10MU to Tay Ninh, 2 to Dau Tieng
0945    LTC Riley is airborne starting at CP11 will move up MSR to Tay Ninh
1314    A Trp reports 554th Engr elements will be at CP17 for the rest of today
1700    A Trp 1SG Barney, security check rear area at Cu Chi: no mines or booby traps
1725    A Trp trail at CP17
1735    A Trp 1st plt departing main gate
1740    A Trp trail at CP16,  also A Trp rolling up
1751    A Trp line up from north to south: 10-20-30
1753    A Trp  trail clear CP09
1756    A Trp 1st plt at CP17
1759    A Trp has 4 APC’s and 16 personnel for bunker line reaction
1806    A Trp 2nd plt at CP09
1814    A Trp 2nd plt at CP17
1845    A Trp 1st plt in night laager, also 3rd plt at CP17
1848    A Trp 3rd plt in night laager
1939    A Trp 3rd plt AP moving out at this time
1955    A Trp 2nd plt AP and sniper team now departing
2014    A Trp 3rd plt reports AP in position
2026    A Trp radar report: 1st plt Left 100deg, Right 280deg, Center 190deg, Range Left 300m, Right
          900m, Center 1000m; 3rd plt, Left 60deg, Right 140deg, Center 104deg, range 9000m
2047    A Trp 3rd plt Sniper team in position
2113    A trp neg sitrep
Nov 11
0323    A Trp Sitrep N/C
0432    A Trp: D Trp AP will remain in position and sweep the area at first light
0503    A Trp 1st and 3rd plts stand-to complete
0505    A Trp 2nd plt stand-to complete
0618    A Trp 1st and 3rd plts beginning road sweeps
0624    A Trp 2nd plt beginning road sweep
0645    A Trp 20 at 200m w/o CP17, A30 at CP10
0652    A Trp  20 at CP17
0653    Aerorifles swept area of A Trp contact w/neg findings
0659    A Trp 30 at CP15, A10 at CP17, A20 at CP09
0710    A 20 at CP16
0721    A Trp sector clear from CP11 to An Duc and down to Mui Lon
0734    A Trp 20 entering Cu Chi, all wpns clear
0745    To A,B,C Trps; Convoy crossed bridge at Phu Cuong
1145    A Trp 10 reports vehicles at CP17 receiving stray rds, the 2/12 is in contact at  
1315    Convoy closed Tay Ninh 1255 hrs
1512    A Trp reports Engrs have left CP08
1545    A Trp southern elements at CP11 to CP16 starting to roll up
1600    A Trp 3rd plt at main gate Cu Chi, all wpns cleared, all personnel accounted for
1603    A Trp rolling up, leaving 2 vehs to secure Engrs at CP17 and CP08
1732    A Trp all elements on move also up from north to south: 20-30-10
1735    A Trp 2nd plt departing main gate
1738    A Trp 3rd plt departing main gate
1739    A Trp security check clear
1740    A Trp 2nd plt at CP16
1748    A Trp 3rd plt at CP09
1749    A Trp permission to fire: 2nd plt 180deg max range 700m, 3rd plt can fire due south
1758    A Trp 2nd plt at CP17
1803    A Trp 1st plt at CP19
1808    A Trp 2nd plt in night laager test firing
1832    A Trp 3rd plt request to fire toward south
1844    A Trp 3rd plt test firing
1937    A Trp 2nd plt sniper team in position, also 3rd plt radar report: 114deg, Right 229deg,      
          Center 172deg, range Left 1000m, Right 1000m, Center 3500m
1939    A Trp AP element they have working w/them have departed for night laager
2001    A Trp 2nd plt radar report; L 20deg, R120deg, C70deg range 500m
2025    A Trp 1st plt AP departing at this time also neg sitrep
2035    A Trp 1st plt AP in position
2056    A Trp: AP that is working with them is in position
2120    A Trp Sitrep N/C
2223    A Trp neg Sitrep
Nov 12
0506    A Trp stand-to complete
0613    A Trp sweep and outpost status:  2nd plt sweep from middle of An Duc to eastern edge of Bao Tre,
          3rd plt sweep from night laager to CP17, then back to CP16 and down to Mui Lon and into base
          camp for stand down, 1st plt sweep from night laager to CP11 and outpost from CP11 to CP17
0614    A Trp 1st plt beginning sweep
0626    SPOTREP A Trp 1st plt (15)  /LB road obstacle / LC Xt614133 time 0615/ LD road obstacle consists
          of tree branches from rubber trees, 3’ high x 3’deep, completely across road/ LE clearing road
0631    A Trp 30 has made turn around, A10 is on move again
0651    A Trp 10 at CP15,
0655    A Trp has swept to Mui Lon
0659    A Trp 10 finished sweep now outposted
0700    A Trp 30 at CP16
0708    A Trp 30 at Main Gate, all wpns clear
0709    A Trp sector clear
0900    Convoy lead at main gate
1527    A Trp asked permission to roll up southern elements, 2nd plt will remain outposting from An Duc to
          CP16 to secure Engrs and convoy
1602    A Trp 1st plt at CP17, HQ101 at CP17
1607    A Trp closed main gate, all wpns elevated, all personnel accounted for
1724    Convoy has closed
1725    A Trp is rolling up
1728    A Trp 2nd plt securing downed Engr veh at XT567164
1758    A Trp 3rd plt at CP09
1803    A Trp 3rd plt at CP17, 1st plt at CP09
1811    A Trp 1st plt at CP17
1812    A Trp security check clear
1819    A Trp 3rd plt at CP08
1908    A Trp 3rd plt in night laager, also 2nd plt in night laager, 1st plt will wait until darker to move out.
1927    A Trp 1st plt reports their D Trp AP that is working with them has departed, also 3rd plt sniper team
          has departed
1947    A Trp 3rd plt radar report: L 100deg, R180deg, C 140deg, range L 400m, R700m, C3000m
1948    A Trp 2nd plt has departed
1956    A Trp report: D Trp AP that is working with them is in position, also sniper team is in position
2013    A Trp Sitrep N/C, also new location: XT528185
2056    A Trp 2nd plt radar report:  Left 34deg, Right 76deg, Center 56deg, range 700m
2126    SPOTREP A Trp, 3rd plt sniper team/  LB XT531184?  LC 3rd plt sniper team spotted 1 light flickering
          on and off on north side of road/  LE sniper team will engage next time light goes on.
Nov 13
0519    A Trp 10 & 30 stand-to complete
0611    A Trp 10 sweep south to Mui Lon then stand down. A Trp 20 sweep south  & outpost from CP11 to
          CP09. A Trp 30 sweep south & outpost from An Duc to CP09
0615    A Trp 10 on the move
0621    A Trp 10 at CP17, A30 on move
0626    A Trp 20 on the move
0629    A Trp 20 at CP09
0653    A Trp:  road obstacle at XT673100, spot report to follow
0711    SPOTREP: LA A Trp 2nd plt/  LB road obstacle/  LCXT673101 time 0650/ LD consists of 25 dirt
          mounds 16” in dia. X 10” high stretching across the road for 30m; directly behind this was one solid
          row of dirt 1’ high w/pal, leaves stuck in mound; behind the mound another 30 dirt mounds, 16”  in
          dia and 10” high, stretching 20m across the road/ LE continuing mission
0716    A Trp: sweep to Mui Lon complete
0741    Ref Log 31: Late Entry: at 0700 A Trp 2nd plt also reports that the spot where the road obstacle
          was is also a spot where OP veh pulled in;  personnel noticed a lot of small foot prints ( child sized)
          believe possibly VC had children make road obstacle.
0754    A Trp: road clear in Sector III, MSR clear and outposted in ¾ CAV area
1555    A Trp 2nd plt except 23 & 26 at main gate, all wpns elevated all personnel accounted for
1724    A Trp lineup from north to south: 10-20-30
1728    A Trp 1st plt departing Cu Chi
1735    A Trp 1st plt at CP16
1740    CO A Trp stopped by Cu Chi Sub, informed them of Aerorifles AP
1740    A Trp 2nd plt departing motor pool
1745    A Trp 1st plt at CP09
1746    A Trp 2nd plt at CP16
1750    A Trp 1st plt at CP17
1819    A Trp reports a down tank from 65th Engrs at CP17
1825    A Trp 3rd plt at CP09
1830    A Trp informs us an ARVN convoy of 14 veh passed his location, also the broken down tank at CP17
          also passed his location; also 1st & 2nd plts in vic of night laagers, 3rd plt east of CP09
1852    A Trp have RF/PF with them
1902    A Trp 3rd plt in night laager
1918    A Trp 1st plt in night laager, AP on way to night laager
1937    A Trp 2nd plt AP departed, also D trp AP that is with 3rd plt departed
1940    A Trp 2nd plt  radar report: Left 120deg, Right 300deg, Center 210deg range 3000m
1948    A Trp 1st plt AP sniper team departing from night laager
1950    msg from G2, VC squad is supposed to infiltrate into Cay Trom, XT585155 to collect taxes,  intel
           information from Cu Chi Sub – A Trp notified
2005    A Trp 2nd plt AP in position
2028    A Trp neg Sitrep
2030    A Trp 2nd plt radar report:  L 3900mills, R 4820mills, C 4630mills, range 400m also 2nd plt AP in
2038    SPOTREP LA 2nd plt / LB 2 moving lights/ LC XT566154 time 2038/ LD 2nd plt observed lights
          blinking in wood line, moving west to east lights blinking on sporatically/ LE requesting arty
2050    A Trp all elements in position
2053    in ref to log 214 at 2038 Arty has been brought in
2140    SPOTREP LA A Trp 2nd plt AP/ LB  1 light/  LC XT583158 time 2140/  LD light blinking on & off/  LE
          moving east to west
Nov 14
0525    A Trp Stand-to complete
0620    A10 will sweep north & outpost from An Duc to CP09,  A20 will sweep N to CP08, then to base
          camp,  A30 will sweep south & outpost from CP09 to CP11
0622    A10 & A20 on the move
0625    A30 on the move
0630    A30 at CP16
0649    A Trp swept from CP08 to An Duc, making turn
0652    A Trp 1st plt at CP 08
0653    A Trp 3rd plt at CP15, 2nd plt at CP08, Turning around
0658    A Trp 3rd plt at CP11
0659    A Trp 3rd plt at CP17
0710    A Trp 1st plt at CP09 outposting
0718    A trp at CP16, road swept in A Trp’s sector
0737    A Trp 2nd plt at main gate Cu Chi, all wpns elevated, all personnel accounted for
0751    A Trp reported lead MP jeep at CP16
0806    advanced of convoy at CP08
1508    A Trp elements between CP11 and CP16 rolling up
1603    A Trp given permission to roll up their elements
1720    A Trp 2nd plt departing motor pool
1740    A Trp 20 at CP09
1747    A Trp 30 departing Main Gate, A Trp 20 at CP17,  A Trp will only have 1 radar tonight at PPS 5
1758    A Trp 30 at CP16
1803    A Trp 30 at CP09, A20 holding in position
1808    A 30 at CP17
1822    A Trp line-up N-S: 20-30-10
1839    A Trp internal security check OK
1920    SPOTREP: A Trp 20/ LB 1 mortar rd/ LC XT527194, time 1919/ LD 75m to south of road/  LE
          continue mission
1951    A Trp 30 AP departing
2004    A Trp 10’s AP’s and sniper team departing
2013    Ref log 212 (1920): LA A Trp 20/ LB lights and movement/ LC XT527194, time 2000 / LD received 1
          mortar rd, returned with Arty, fire is burning, also spotting lights & movement, Arty is continuing, A
          Trp Firing into area/  LE continue mission
2033    Ref log 217, & 212:  Arty was fired into area, then 3-4 lights were spotted moving west to east, will
          engage with main gun
2037    A Trp 10 & 30 AP’s in position
2050    A Trp reports an ARVN AP (XT567164), A trp will move their AP back towards their plt.
2115    CPT Skiff gave A Trp permission to roll back their AP that was in conflict with ARVN AP
2122    A Trp 20 radar boundaries: L112deg, R 227deg, Range 3000m
2400    ¾ CAV Sitrep N/C
Nov 15
0508    A Trp all platoons stand-to complete
0623    A Trp: 2nd platoon will sweep from An Duc to east of  Bao Trai, 3rd plt will sweep from night
          location to CP19 and Mui Lon and then go back to Cu Chi; 1st plt will sweep from night location to
          CP11 east of Bao Trai; 3rd platoon on the move at this time.
0628    A Trp 1st & 2nd Plts on the move at this time.
0634    A Trp 1st Plt at CP09, 2nd Plt at western edge of  An Duc, 3rd Plt  at CP17
0644    SPOTREP: LA A Trp 2nd plt/ LB road obstacle/ LC XT531383 Time 0635/ LD debris from hootch,
          table, wood, roofing, bricks, completely covers road, 50m long x 1m high at highest point/ LE
          continue missionhttp://home.comcast.net/~ATrp3-4Cav/Photos/index.html
0710    A Trp sector clear except for 200m north of CP09, from CP09 to southern sector outposted
0715    CO A Trp calls Dust Off for 1 VN (male) who caught scrap metal in chest when road obstacle was
          cleared ( ref log 52)
0729    CO A Trp is taking injured VN civilian out of built up area on stretcher.  Dust off 158 will pick up in
          10-15 min.
0738    A Trp: Dust off complete at 0737Hrs
0745    A Trp stand down 3rd plt is at CP16
0757    A Trp  3rd plt entering main gate, all wpns clear, all personnel accounted for
1600    A Trp 1st plt in Cu Chi
1602    A Trp 2nd plt coming in except 1 tank and 1 APC at An Duc w/Engrs and 2 APC’s at CP17 w/
1615    A Trp 2nd plt at main gate, ref log 156 Engrs at An Duc finished, veh moving to CP17
1715    A Trp:  leaving net at this time
1807    A Trp internal security check neg, also has 21 APC’s  & 8 Tanks for reaction force.
Nov 16
1901    A Trp internal security check neg: bunker line defense status: 21 APC, 8 Tanks
Nov 17
1817    A Trp internal security check neg.
Nov 18
          No entries for A Trp in staff journal.
Nov 19
0655    A Trp 10 on the move
0659    A Trp 30 on the move
0707    A Trp elements 10,20,30, have cleared Chu Chi main gate
0712    A Trp 2nd plt at CP16
0716    A Trp 30 at CP16 & A Trp 10 at CP09
0720    A Trp 20 at CP09
0729    A Trp 20 at CP08.
0735    A30 in position. A10 spotted some propaganda leaflets nailed to stakes on the right side of MSR at
0821    A Trp day time outposted north to south. A10, A20, A30.  2 APCs at Trang Bang Bridge & 1 at
          Laterite Pit.
0950    A Trp trains departing Cu Chi for Trang Bang.
0953    A Trp had 2 APCs sweep to CP11 and return to CP10.
1035    A Trp trains closed FSB Stewart.
1210    A10 will conduct RIF today.  A20 will oupost at CP08 & A30 will outpost at CP07.
1318    A Trp Cdr (Cpt Moss) request info on the Engrs that are bing flown to his position.
1355    A30 is at CP09.
1403    A30 is at CP08.
1414    A10 has started the RIF.
1417    A10 is at CP L.
1427    A10 is at CP X.
1431    A10 is at CP Z and starting RIF at this time.
1510    A10 is at CP T.
1515    A10 is at CPB.
1543    A30 is rolling up behind the convoy.
1550    A10 has 2 tanks stuck 1000 meters east of CP08.  No assistance needed.
1640    A20 is securing Engrs at An Duc & CP08.
1658    A Trp still has one tank stuck.  His VTR is being escorted out to the position now.
1714    SPOTREP: LA A10; LB APC hit mine; LC XT465238 at 1712 hrs; LD A13 hit mine, negative casualties;
          LE enroute to determine damages to vehicle. (photo)
1743    Follow-up to SPOTREP: crater was 4' wide & 3' deep, north side of road, damage to vehicle: final
          drive housing and roadwheels on right side, don't know if combat loss or not. Crew Phil Hicks, Alton
          Philips, Paul Brockman, & Warren Yeagley.
1802    A Trp has damaged APC hooked up to a tank.  VTR enroute to stuck tank.
1843    A20 departing FSB Stewart enroute to night laager.
1852    A20 & A30 in night laagers.
1858    A Trp line-up north to south: A20, A30, A10.  A10 will be located at XT442269 due to the tank
          being stuck.
2001    A30 ambush patrol is in position.
2002    SPOTREP: LA A20; LB campfires; LC XT532192 at 1959hrs; LD campfires in the woodline observed
          with a starlight scope; LE preparing to engage with 50 cal and 90mm.
2040    A Trp & HHT report CS gas in troop area in Cu Chi.
2118    A Trp radar scans:  A30 L 250 mills, R 750 mills, C 500 mills & range 1000 meters; A20 L 2400 mills, R
          4300 mills, C 3350 mills & range 1000 meters.
Nov 20
0508    A Trp Stand-to complete.
0620    A Trp sweeps: A30 south of FSB Stewart, 2 APCs sent to escort VTR; A20 sweep to A30's night
          laager, do a 180 turn at CP20 and outpost at CP08.  A10 works on stuck tank.
0625    A20 is on the move.
0648    SPOTREP: LA A10; LB road obstacle; LC XT535181 at 0640hrs; LD wood racks & straw 25 meters
          long & 5 meters wide & 1.5 meters high across road; LE clearing and continuing mission.
0717    Cdr Trp A: A20 ouposting Vic CP07, A30 outposting vic CP08, & A20 will conduct RIF.
0756    A10 tank is unstuck, returning to the MSR.
0825    SPOTREP: LA A10; LB mine; LC XT435272 at 0820 hrs; LD Engr found 1 Soviet P-27 AT mine in mud
          hole, had been run over several times; LE engrs blew in place, was a dud.
1110    A20 has started his RIF.
1211    A20 location at XT534186.
1236    SITREP: A20 is 75 meters from XT534186, moving north, found 10 old foxholes that have not been
          used recently.
1249    A20 RIF is 800 meters north of MSR.
1302    A20 is complete to XT533192.
1308    A20 RIF is 50 meters from XT533192 and found 1 M17 protective mask.
1330    A20 is at northern most extent of RIF, turning around to sweep back south.
1334    A20 has completed turn-around.
1420    SPOTREP: LA A20; LB tunnel; LC XT534196 at 1416hrs; LD tunnel of left of road going south parallel
          to road, just big enough for one person to crawl through; LE blew tunnel & continued mission.
1435    A20 RIF completed.
1448    A20 enroute to FSB where Cpt Moss will debrief.
1520    A Trp rolling up elements from the north.
1620    A30 rolling up his elements from the south, leaving some elements to secure Engrs at An Duc &
1804    A Trp elements departing FSB Stewart enroute to night laagers, A30 elements will stay at FSB.
1807    A20 is at CP08.
1812    A20 is at CP17.
1819    A10 is at CP08.
1823    A20 is in NDP.
1833    A20 has been given permission to test fire.
1932    A20 ambush patrol is departing night laager.
1948    A20 radar scans: L 2800 mills, R 4740 mills, C 3770 mills & range 100 meters.
2007    A20 ambush patrol is in position.
2015    A10 radar scans: L 3800 mills, R 5400 mills, C 3200 mills & range 800 meters.
2027    A10 sniper team is in position.
Nov 21
0020    A Trp negative SITREP
0035    Message to A Trp from Maj Harrison, S3, at first light or ASAP A Trp is to send 1 platoon to sweep
          to the FSB at Go Dau Ha (XT443238) and relieve A/2/22 IN and secure the Arty battery at FSB.
0535    3rd PLT will relieve A/2/22 IN at FSB near CP20.  A20 will outpost at CP08 and A10 will outpost at
0618    A30 departing FSB.
0627    A30 at CP07.
0628    A20 is on the move.
0629    A10 is on the move.
0653    Message from Cpt Johnson, 3rd Bde 25th ID, A30 is to pickup the Arty battery on MSR between CP5
          & CP21.
0715    A30 has picked up 2/77 Arty and moving to CP20 for security mission.
0731    Cdr A Trp: new location for 2/77 Arty FSB changed from CP20 to XT385254.
0751    A10 and A20 in day positions.  A20 enroute to CP09.
0800    A20 in day positions.
1602    A Trp mission for tonight:  secure Arty at Go Dau Ha with 1 platoon, have 1 platoon at FSB
          Stewart, & 1 platoon on standdown at CP08 w/no ambush patrol.
1722    A20 is at FSB vic CP05.  A30 is at CP20 headed to FSB Stewart.
1814    A20 enroute to night laager.
1825    A20 holding just short of CP08.
1853    A30 is in night position.
1924    Cpt Moss reports that someone is firing out of Thai My again just like they did last night.
2020    A10 radar status: left 3000 mills, right 5200 mills, center 4100 mills, & range 800 meters.
2030    SPOTREP: LA A30; LB 2 lights; LC XT526180 at 2030 hrs; LD seen with starlight scope what appears
           to be 2 cigarettes or pen lights; LE did not engage because they went out.
2045    A30 ambush patrol is in positon.
2247    SPOTREP: LA A30; LB 1 individual; LC XT525186 at 2239 hrs; LD recieved AK sniper fire from north
          side and south side of MSR, returned fire with all weapons; LE require dustoff for 1 US wounded
          w/head wound, urgent but ambulatory, and Arty being called in.
2315    Cu Chi Dustoff: ETA 15 min callsign Dustoff 158.
Nov 22
0008    Dust-off completed.
0031    A Trp has a negative SITREP.
0516    A Trp stand-to is complete.
0623    A Trp sweep status: A10 from night laager south to CP15 then to FSB Meade; A20 will sweep south
          to FSB Lincoln to CP07, remain at CP07 until rest of plt is released from FSB Lincoln, then outpost at
          CP08; A30 will sweep through An Duc then back to CP07.
0626    A30 beginning his sweep.
0645    SPOTREP: LA A30; LB road obstacle; LC at XT531183 at 0640hrs; LD boards, branches, and trash
          spread across the road 30 meters in depth; LE being cleared by engineers.
0658    A10 & A30 will conduct plt RIFs.  A20 will outpost CP08.
0731    A10 is in day position.
0740    A Trp notified of makeup of convoy units.
0900    A Trp Cdr rpts his element at CP08 has some sporadic automatic wpns fire near CP17.
0905    2nd Bde 25th ID rpts ambush of convoy vic CP17.  A Trp sending 4 APCs to secure convoy.  More
          to follow.  Fire seems to be coming from XT562188. 1/5 INF also at the scene.
0931    Cdr A Trp reports firing stopped at 0920hrs.  Negative reports on casulaties or damages.
0934    Cdr A Trp reports convoy was ambushed at XT562168.  Drew automatic weapons fire from the north
          side of the road, one explosion from possible RPG, 2 plts from 1/5 INF at scene along with 2 plts
          from A Trp.  Firing stopped at 0920hrs, no casualties or damages.
0950    A20 elements are now OPCON to 2nd Bde 25th ID for sweep of ambush site at CP17.
1010    A20 and plts from 1/5 INF being sweep.  Have a possible contact and firing maingun.
1020    A30 has crossed his LD and is putting dismounts out.
1100    A6 reports that A10 has found 3 fighting positions and tunnel at XT611139.  Very old & unused,
          blown in place.
1106    A10 & A30 have crossed Line Tango on RIFs.
1111    A10 is at Phase Line Pear.
1132    A20's sweep for 2nd Bde is complete. found 2 detainees and some spider holes.  Returning to MSR.
1135    SITREP: A6 rpts A10 & A30 are just short of search area.  Only 3 bunkers found so far, all 
1145    A6 reports A10 & RF dismounts making sweep of area.
1328    SPOTREP: LA A10; LB attached RF's claim one body count; LC XT605152 at 1315hrs; LD 1 VC in
          tunnel; LE threw grenade in tunnel, 1 body count, continuing mission.
1352    A20 has 4 vehicles stuck 400 meters north of CP08, will probably need a Chinook.
1359    A20 released from OPCON to 2nd Bde.
1402    RF's departing from A Trp, A10 is back on the MSR.
1404    A Trp rpts 1 ARVN element heading back to Cu Chi and 1 will stay with A Trp.
1425    A Trp in open heading into rubbers & returning to the MSR.
1445    A Trp A20 elements stuck at XT555176.
1446    2nd Bde in contact at XT560190 receiving sniper fire.  Will not interfere with A20 recovery.
1456    A Trp on Ambush Rd retuning ARVNs to Cu Chi.
1502    A30 is at CP16, A10 is at CP08 & will stay in Horseshoe until convoy passes.
1655    A Trp has commo with 1st plt of B Trp, rest of B Trp is at Tay Ninh, 1st Plt B Trp OPCON to A Trp.
1745    Late entry A20 has 1 more APC out for a total of 2, working on getting 3 more out.
1800    A10 elements received 1 mortar round, calling in gunships.
1816    A20 has 1 more APC out, 1 more still stuck.
1836    A20 has all APCs unstuck but one threw a track at the same location.
1838    SPOTREP: LA A10; LB 1 mortar round; LC XT555176 at 1800 hrs; LD 1 mortar round approximately
          20 meters in front of their lead vehicle; LE negative casualites or damage, put suppressive fire out.
1840    A20 has track back on vehicle and element is on the move.
2034    A Trp has negative SITREP.
Nov 23
0100    A Trp has negative SITREP.
0610    A Trp sweeps: A20 south to CP15 and return to CP07; A30 north to CP05; A10 to FSB Stewart,
          south down RT 8A to FSB Meade.
0624    A30 is on the move.
0644    1st Plt B Trp released from OPCON to A Trp.
0647    SPOTREP:  LA A30;  LB road obstacle; LC XT538180 at 0653hrs; LD consists of branches, schrubs,
           & a box, no mines or booby traps; LE cleared by engineers.
0706    SPOTREP: LA A10; LB road obstacle; LC XT400243 at 0705hrs; 75 dirt mounds 50 to 75 meters in
          depth across road, mounds 12" in diameter & 6" high; LE engineers cleared.
0750    A20 completed sweep of Mui Lon.
0750    A30 in outpost positions.
0825    A10 will conduct RIF today; A20 will outpost at CP07; A30 will outpost at CP08.
0832    A20 in outposting position; A10 is at the Horseshoe (CP16).
1014    A10 made coordination with Cu Chi Sub-sector.  RF force will be ready.  RIF will start at 1100hrs.
1054    A10 & RF's moving to RIF location.
1114    A10 began RIF at 1100hrs.
1145    A10 at RIF location (XT635155); negative SITREP.
1228    A10 rpts RIF complete with negative results.
1250    A10 is at CP A.
1253    A Trp is at XT640150; negative SITREP.
1303    A10 is back at Ambush Gate.  He will give a SITREP on the RIF with the RFs.
1332    A10 reports the road from Wolfhound Gate to the 62 north/south gridline is passable for tracks &
          wheels; from the 62 to the 61 for tracks only; west of the 61 is impassable due to bomb craters and
          brush 3' high.  On the RIF with the RFs at XT615170 was a built up area with hooches that are
          shown on the map but no longer there; some foot trails but negative findings.
1337    A30 rpts that a US jeep pulled up to their position at 1320hrs at CP08 and said they received 2 or 3
          rds of sniper fire in An Duc.  A30 sent 2 APCs to check with negative results.
1811    A30 enroute to night strong point position.
1835    A dustoff helicopter landed on MSR at CP08; ship had an oil leak & was carrying VC POWs; A Trp
          will secure until relieved or recovered.
1856    A Trp Cdr reports POWs have been picked up by another helicopter.  Helicopter vic of XT5420 being
          shot at, no return fire.
1901    G3 to LTC Riley directing that the 3/4 Cav will begin working in the Filhol Woods AO.
1930    Change of mission to A Trp & C Trp for squadron operations in Filhol Woods AO.  Info to LTC
          McGowen & 3rd Bde.
1936    Cdr A Trp is securing downed Dustoff ship.  Will be pipe-smoked out on 24 Nov.
2002    A30 ambush patrol is in position.
2320    A Trp negative SITREP.
Nov 24
Summary:  A & C Trps begin squadron operations in Filhol Woods.
0023    A Trp Negative SITREP.
0543    A Trp sweep status:  A10 to sweep thru An Duc then back to FSB Steart & outpost at CP08; A30
          sweep from NDP to CP15 then back to CP16 & to FSB Meade at Mui Lon; A20 sweep from FSB
          Stewart to CP05 and return to the FSB.
0620    A30 is on the move.
0632    A10 & A20 are on the move.
0721    A30 is heading back to base camp.
0723    A20 is at CP19 and heading back to Cu Chi.
0727    Down chopper north of An Duc is blocking traffic.  Will be 30 minutes before removed.
0746    A30 is in Cu Chi.
0751    A Trp is monitoring heavy volume of fire at XT522197, civilians in large numbers leaving area.
0753    Chinook is about 2 min out of A Trp Cdr's location.
0756    Chinook is landing at this time.
0800    Corrected entry was not a Chinook that landed.  Was a slick with straps for the Chinook.
0822    Chinook and downed chopper are off of the MSR.
0836    A20 is at the south edge of An Duc headed for Cu Chi.
0838    A10 is at CP17.
0840    A Trp trail element is at CP Oscar.
0845    A Trp lead is at CP09, trail is at CP17.
0851    A Trp trail at CP09, MU 1 almost to Phuou Hiep.
0853    A10 is held up at the Horseshoe.
0857    Rest of A Trp is held up at CP16.
0919    A10 is at the main gate at Cu Chi.
0926    A Trp trains at CP 16 headed towards main gate.
0932    A20 is in Cu Chi at this time.
1345    A Trp departed motor pool.
1359    A Trp departs Ann Margaret Gate for AO Blue (Filhol Woods).
1415    A Trp reaches AO Blue and putting out dismounts.
1441    A Trp has cleared their SP.
1500    SPOTREP: LA A20; LB spider hole; LC XT675174 at 1455hrs; LD scout track ran over fallen tree
          which disclosed the hole; LE destroyed with a grenade.
1537    A Trp located at XT682179.
1550    A Trp moving into its NDP.
1742    A Trp internal security check (base camp) is negative.
1750    A Trp is at XT690184.
1830    SPOTREP: LA A20; LB tunnel; LC XT690184 at 1810hrs; LD tunnel spotted while setting up trip flare,
          7' deep & 30' long, fresh prints in area; LE blown in place with C4 & CS.
1915    A Trp request dustoff, 1 US ambulatory at XT690184.  LZ is secure, shrapnel in right arm.
1923    A Trp ambush patrol and listening posts in position.  AP is at XT671182, LP at XT685182.
1945    A Trp dustoff is complete.
2000    A10 & A30 ambush patrols in position.  A10 ambush patrol at XT692175 & A30 ambush patrol at
2352    A Trp has a negative SITREP.
Nov 25
0033    Message from Col Taylor, G3, at 0730hrs 3/4 Cav is OPCON to 2nd Bde 25ID.
0536    A Trp standto is complete.
0733    A10 & A30 on RIFs.  A20 checking tunnel found last night.
0754    A Trp starting RIFs at this time.
0938    A Trp is at XT692178.
0944    A Trp tank has sheared a sprocket.  Being towed back to base camp.
1025    A Trp APC having trouble bringing up security, 15 minute wait.  VTR with down tank is in Cu Chi.
1055    SPOTREP:  LA A10; LB tunnel complex; LC XT692182 at 1030hrs; LD tunnel leds to room 4'x5'x6',
          fresh ashes, 1 booby trap (Chicom hand grenade), mud wall with wood wall behind it; LE tearing
          down wall.
1135    SPOTREP:  LA A6; LB hedgerow w/bunkers; LC XT695177 at 1135; LD fresh bunkers & spider holes
          heavily booby trapped; LE calling in Arty.
1141    A Trp leaving bunker area and calling in Arty.
1246    A Trp still at XT695177
1329    A Trp on move to investigate bunker complex that Arty fire was called on.
1353    A Trp knocking down bunkers with tanks, dismounts checking area.
1357    A10 securing hedgerow, A30 continuing RIF.
1410    SPOTREP:  LA A6; LB recently setup area; LC XT698177 at 1410hrs; LD several booby traps, tunnel
          complex, fighting positions 2'x3', spider holes 5'x2', VN clothes, rat traps with rats in them, bamboo
          stakes and barbed wire for a perimeter.
1421    A Trp: clum of trees w/path leading into it.  Recon by fire into it with M60.
1430    A30 has destroyed 15 fighting positions, now on the move.
1450    SPOTREP: LA A6; LB foxholes; LC XT695177 at 1450 hrs; LD 20 to 25 foxholes, 5'x1', 1 week old,
          tunnel 8 to 10 feet long a couple of days old, ground recently broken up, 2 VC grenades, 4 bottles
          of lotion, 1 VC hat, 1 bottle of pills; LE destroying.
1455    A Trp found a cave 1.5' deep leading into a thicket, found some old bandages.
1526    SPOTREP: LA A10; LB entrenchment; LC XT697181 at 1510; LD L-shaped entrenchment 16'x2'x4',
          trail leading to it, spider holes containing rice canisters, bamboo baskets, 2 Chicom grenades on
          trail make unknown; LE bringing one grenade in for ordnance and destroying one.
1540    A Trp moving back to road where they started.
1553    A Trp: HQ6 & 7 tracks with C Trp at this time.
1622    A Trp night position is at XT695186.
1900    A Trp internal security check (base camp) complete.
1908    A Trp ambush patrol locations: A10 XT697189, A20 XT693185, A30 XT694188.
1927    A20 ambush patrol and listening post departing for positions.
1940    A20 AP & LP in position.
2021    A Trp requests resupply of watter & Class I.
2222    A Trp has negative SITREP.
2307    LTC McGowen rpts 3/4 Cav DEFCONS for A & C Trp are being fired at this time.
Nov 26
Summary: Troop moves North of the Filhol into the Iron Triangle NDP on the bluffs above the Saigon River.
          1st Platoon tanks engaged sampans on the Saigon River.
0025    A Trp has negative SITREP.
0536    A Trp stand-to is complete.
0718    A Trp reports that the ARVNs are at his location.  Sending one element to C Trp.
0753    A Trp requests that they test fire 1 50cal due north on the ground.  Request granted.
0818    A10 & A20 on RIFs.  A30 securing the trains and will follow later.
0827    A Trp has 45 ARVNs on the sweep with them.
0848    A Trp location is XT695195.
0857    A Trp rpts old bunkers that have been destroyed by air strikes/napalm.
0909    A Trp location XT694197.
0938    A & C Trps request permission to reconn by fire into hedge row.  Denied by 2nd Bde.
0956    A Trp location is XT688200.  From this location they will sweep southwest to AO boundary then
          turn around & sweep to new night location.
1033    A Trp starting sweep to the southwest.
1121    A Trp is at wood line on a 250 degree azimuth.  Putting dismounts out at this time.
1150    SPOTREP:  LA A20; LB spider hole;  LC XT687200 at 1140hrs; LD 5'x2'x1.5'; LE negative findings,
          being destroyed.
1200    A Trp location XT679202.
1321    A Trp location XT679205, starting turn around on sweep.
1407    A Trp has completed turn and is moving out.
1410    SPOTREP:  LA A20; LB 2 objects;  LC XT682202 at 1400hrs; LD 2 objects shaped like football
          gridlines, made of bamboo strips, 3' off ground, 5' long facing southeast & one northeast, with
          hole 2'x2'x2' between them; LE negative findings,
1424    A Trp reports objects are ARVN special forces setup.
1456    A Trp position is XT693204.
1617    A Trp reports a camouflaged barge floating down the river, request permission to engage. 
          Permission granted by 2nd Bde if they do not hit far bank of Saigon River.  A15 engages barge.
1627    A Trp NDP is at XT689213.
1635    2nd Bde has given clearance to fire on anything on river if the opposite bank is not hit.
1649    SPOTREP:  LA A10; LB 2 bardges on river;  LC XT696213 at 1625hrs; LD 1st barge left the bank,
          paddled to middle of river, is camouflaged, a 2nd barge drifted out; LE 1st barge out of range,
          engaged 2nd barge, hit it and saw secondary explosion.
1910    A Trp ambush patrol locations: XT689217, XT689200, XT687212; Listening posts at XT689211,
          XT688212, & XT700212.
1932    A Trp radar: left 1000 mills, right 2400 mills, center 1200 mills, & range 1000 meters.
2018    Major Harrison informed A & C Trps that ARVNs must be returned to compound by 1000hrs.
2020    A & C Trp told to disregard message on ARVNs.
2025    ARVNs will be available to trps for next 2 days.  Pick up at XT705195 hrs tomorrow.
2400    A Trp has negative SITREP.
Nov 27
Summary:  LTC McGowan assumes command of squadron.
0100    A Trp has negative SITREP.
0549    A Trp stand-to complete.
0621    Cpt Moss rpts that A Trp's DEFCONs did not get fired last night.
0855    B Trp received RPG fire, request dustoff for 3 litter & 4 ambulatory at XT695210.
0925    B Trp will establish blocking position at current location.  A Trp will sweep into that area from the
0940    S3 & Trp Cdrs enroute to location of contact.
1055    A Trp has their element on the north sweeping SW to NE.  Sweeps moving at this time.
1114    Maj Harrison to LTC McGowan:  The sweeps with A Trp & B Trp are going slowly, encountering many
          spider holes with recent activity & bunkers.
1120    SPOTREP:  LA A30; LB spider holes; LC XT695210 time now; LD spider hole 2x3x5', also 2 saws,
          rain jackets, shower shoes, fresh meat, & rice nearby; LE blown in place.
1121    SPOTREP:  LA A30; LB spider holes; LC XT695210 time now; LD spider hole 2x3x5', recent activity;
          LE blown in place.
1141    SPOTREP:  LA A30; LB well; LC XT695210 at 1132hrs; LD well w/soap & paths, brush for overhead
          cover; LE blown in place.
1150    A Trp resupply on the way into Cu Chi.
1203    SPOTREP:  LA A30; LB pole; LC XT695210 at 1146hrs; LD poll flattened on one end, possible paddle
          for sampan.
1210    A Trp resupply at Ann Margaret gate.
1211    SPOTREP:  LA A30; LB ammo; LC XT695210 at 1200hrs; LD spider hole containing 6 claymore mines,
          1 81mm mortar round, 1 can of 50cal ammo, 3 Chicom grenades, 1 can of AK47 ammo, 4 M79 rds,
          2 to 3 lbs of rice, entrance 2' in diameter.
1239    Followup to last SPOTREP: spider hole entrance leads to room 7'x15'x7'.
1331    SPOTREP:  LA A20; LB 2 sampans; LC XT697213 at 1310hrs; LD 2 well camouflaged sampans in
          brush 50 meters apart moving from center of Saigon River to south bank; LE continuing to observe.
1404    LTC McGowan enroute to A Trp's location.
1440    A Trp located at XT695208.
1547    A Trp located at XT678219.
1736    A Trp night position is at XT687218.
1842    A Trp ambush patrol #1 at XT688219, listening post #1 at XT689216, & listening post #2 at
1955    A Trp LPs in position.
2011    A Trp LPs in position, AP departing.
2027    A Trp AP in position.
2055    SPOTREP:  LA A Trp Radar; LB engine on river; LC XT687223 time now; LD engine noise on radar,
          moving down river; LE will attempt arty then maingun after arty ceases.
2155    A Trp Arty mission complete.
Nov 28 (Thanksgiving 1968)
Summary: Special Thanksgiving meal with all the trimming from mess section. (view menu) Col McGowan
           lands in NDP and scoops out ice cream for the troops. 1st Platoon tank fell into tank trap in
           camouflaged bomb crater.  1st Platoon and Troop Hqs laagered around tank for the night in thick
           brush. 1st PLT 60mm mortar periodically fired throughout the night. (View annotated map of area)
0527    A Trp negative SITREP.
0702    LTC McGowan to Maj Harrison:  Setup NLT 1500hrs to set up properly.  A Trp to secure squadron
          trains in the field.
0802    SPOTREP:  LA A65; LB RVN Marine fighting positions; LC XT686215 at 0755hrs; LD moved to
          set-up for today's operation, moved through RVN Marine positions, found c-ration cans; LE
          destroyed by running tracks over them.
0925    Maj. Harrison to all Trps; Change of Mission: instead of a detailed search, recon the entire Filhol,
          moving as far and as rapidly as feels safe.  RIF area for A Trp from XT678215 to XT665234 to
          XT650215 to NDP at XT676225.
1000    A Trp located at XT588216.
1118    A Trp located at XT674214.
1300    SPOTREP: LA A10; LB spider hole; LC XT656224 now; LD 1 spider hole 3'x4'x4', cigarette ashes
          about a week old; LE blowing in place.
1415    A30 is braking track on a tank, rest of trp proceeding to NDP.
1505    A10 & A20 in NDP at XT676255.
1645    A30 elements are now in the NDP.
1649    A Trp (rear area Cu Chi base camp) internal security check Negative.
1713    A Trp escort for RVN Marines arrived at C Trp NDP.
1756    A Trp is at XT676277.  Trains have reached their location.
1818    SPOTREP:  LA A30; LB bunker & tunnel complex; LC XT677277 at 1810hrs; LD bunker 8'x4'x4',
          entrance is 2x'x3'x3, tunnel length not known; LE blown in place.
1823    A Trp listening post positions: XT676224, XT678225, & XT676226.
1941    A Trp will recon by fire on the river occasionally throughout the night.
2010    A30's ambush patrol has departed NDP.
2024    A Trp's LPs are in position.
2056    SPOTREP: LA A30; LB movement; LC XT676228 at 2045hrs; LD spotted movement, reconned area;
          LE continuing to observe.
2105    A Trp radar status: Left 0000 mills, right 1995 mills, range 800 meters.
2230    A Trp has negative SITREP.
Nov 29
Summary: VTR stuck after trying to extracting tank. Finally, several vehicles daisy chain a together to pull
           VTR out.  Ramp pulled off of APC. 1st platoon crew unbeknown to Plt Ldr & Trp Cdr shoots large
           hawk during recovery operation. (photo)
0530    A Trp stand-to is complete.  Negative SITREP.
0805    A Trp lined up and moving out on RIF.
0806    B Trp has 1 tank with a blown pack.  Being towed to A Trp's location so that VTR can bring it into
          Cu Chi.
0857    A Trp location is XT669209.
1002    A Trp location is XT669209.
1200    A Trp has VTR stuck, working on it now at XT660213.
1253    A Trp is linking 3 APCs & 1 tank to VTR to pull it out.
1320    A Trp has the stuck VTR out.  Is now working on the stuck tank.
1430    A Trp has tank partially clear, will wench him out with VTR.  1 APC down with final drive failure & 1
          1 APC with track off.
1519    A Trp tank unstuck, 2 APCs are bieng lined up.
1630    A10 is at XT683200.  The VTR has run out of gas.
1640    A Trp needs MOGAS.
1657    A Trp has 1/2 tank of gas in VTR but can't get it started.
1702    A10 element is at a halt.  Tank has thrown a track.
1704    A Trp has VTR started.
1735    A30 w/resupply is at XT700190.
1757    A Trp closing into NDP.
1855    A20 has closed into NDP.
1945    A30 listening post departing NDP for XT697188.
2009    LTC McGowan requests Engr Stakes & PSP in large quatity tomorrow for all troops.
2011    A30 LP in position; A20 LP at XT699190; A10 LP at XT699188.  Radar operational: left 0000 mills,
          right 2400 mills, center 1200 mills, & range 800 meters.
2100    A10 LP is in position.
2145    From LTC McGowan: trains will be escorted by C Trp to A Trp; A Trp will escort trains to B Trp;  A
          Trp will pickup B Trp trains and escort to Cu Chi.  Return trains will be escorted by HHT Radar.
2220    A Trp negative SITREP.
2345    Late Entry:  At 2310 hrs an explosion occurred in the A Trp area in base camp.  DISCOM has sent
          EOD to location.  EOD states it was a grenade explosion.  A Trp will hold police call in AM to find
Nov 30
Summary: Troop re-enters Ann Margaret Gate, passes through base camp, and proceeds to MSR vicinity of
          Trang Bang.  Operations center out of FSB Stuart.  Lt Ayres assigned to A Trp. Colonel McGowan
          directed A Troop to conduct operations in the Ho Bo Woods. Troop minus the tank sections moved
          north of FSB Patton and Trung Lap into the Ho Bo Woods. Terrain too soft for the tanks. Late
          afternoon, light small arms fire & A10, Lt Moore's APC, hits a mine. A10 driver (Kurt Pearson) and TC
          (Carroll Hawley) medevaced by helicopter. Troop continued movement North to enter Ho Bo Woods
          with only APCs. SA fire received and returned.  Tanks called forward. Movement continued till
          directed to move to NDP and link up with tanks and colonel's track. Troop proceeds to NDP and
          about to link up with Lead 3 Tanks get stuck up to sponson boxes just as they leave dirt road in
          soft ground. (photo) Remainder of tanks redirected to FSB Patton.  Troop sets up NDP around 3
          stuck tanks. A10 recovered and pulled into NDP. (photo) Two 1st platoon tracks go back South and
          escort 1SG Barney up to NDP with hot meal and resupply. (photo of 1SG Barney) Col. McGowan's
          helicopter just about to touch down in NDP, hit by RPG after dark. RPG entered fuel tank and blew
          through and out the other side. Word received Colonel McGowan, 2 pilots, door gunners and S3 Air
          all secured. Troop returned fire. (view annotated map of area)
0100    A Trp Negative SITREP.
0101    CPT Allstatt of the EOD Team reports that the explosion in A Trp's base camp area yesterday was
          a M79 rd based on his findings.  All that was left was small pieces of shrapnel & no crater.
0722    A Trp closing on Cu Chi base camp.
0724    A Trp at Ann Margaret gate, all personnel accounted for & all weapons clear.
0734    A Trp in base camp, all personnel accounted for & all weapons clear.
0825    A Trp clear Ambush Gate.
0852    A Trp at CP16.
0900    A Trp lead is at CP17.
0901    A Trp is at CP17, sweep is down and are far enough to the side to let traffic pass.
0920    A Trp is clear CP 17.
0933    A Trp is at XT680181
0941    A Trp location is XT573175
0950    A Trp has meet other sweep team.
1011    A Trp location is XT585200 on Hwy 7A.
1033    A Trp location is XT599222.  They are leaving the tanks back.
1045    A Trp reports that  at XT595220 the culvert is blown out, rice paddies are deep and 3.5' of water
          on the trail.
1106    A Trp reports that the A30 element is the only one across at this time.
1120    SPOTREP:  A Trp; LB explosion; LC XT608223 now; LD sounds like blowing explosives; LE could have
          been arty.
1128:   SPOTREP:  A Trp; LB 1 woman; LC XT609221 at 1121hrs; LD came out of hooch yelling "beau coup
          VC, beau coup mines"; LE no interpreter, continuing mission.
1145    A Trp reports vehicle hit a command detonated mine, receiving small arms fire, developing
          situation, need dustoff for 2 personnel on litters, at coordinates XT610220.
1150    Dustoff 155 ETA 10-15 mins.
1213    Dustoff complete.
1229    Sqdn Cdr requests someon to come out to pickup A Trp's APC.
1240    A Trp to send it's VTR up to pick up it's combat loss APC by order of MAJ Harrison.
1320    A Trp is presently at XT616232.
1428    A Trp is located at XT603248.
1429    SPOTREP: LA A Trp; LB camouflaged hooch; LC XT601248 at 1422hrs; LD old hooch, one blue shirt,
          some khakis on top; LE ran over and destroyed hooch, continuing mission.
1432    A Trp reports A30 elements towing the down APC hit a mine and are receiving small arms fire at
1444    A Trp reports sniper fire has stopped and are hooking up downed APC.
1445    2nd Bde 25ID is working on clearance for air strikes around XT598221.
1506    LTC McGowan rpts that the road only passable to APCs.  The subelemnts are spread to far apart to
          conduct RIF effectively, request they be able to regroup at XT616227.  The sub elements are at
          Trung Lap, XT602248, & XT613255.
1520    2nd Bde 25ID directs A Trp regroup and setup NDP at XT616227.
1705    A Trp trains at CP17.
1747    A Trp has permission to recon by at XT602221.  2nd Bde 25ID requests they mak max us of M79 &
          do not fire at houses.
1850    A Trp has 1 tank stuck and bringing up another to pull it out.
1940    A10 elements at XT613227.
1950    Everything south of the 21 grid cleared by CPT Skiff from 010800hrs to 011800hrs for DIV Chemical
          for defoliation.
1952    A Trp: 4 vehicles stuck at XT613227.  Have APC starting to work on them.
1955    A Trp night location is northeast of Trung Lap.
2025    LTC McGowan's C&C helicopter hit by ground fire.  He is alright, firing has stopped.  A Trp returning
          fire at this time,  A Trp perimeter receiving mortar rounds.
2030    Ref: 2025hrs helicopter was hit by rpg while landing, crew and LTC alright, gunships called,
          XT615226, Arty coming in.
2037    Gunships on their way at this time.
2040    Heavy Fire Team on station at this time.
2045    Arty in check-fire due to HFT on station.
2049    A Trp not receiving fire at this time, have moved to location of stuck tanks.
2135    SPOTREP: LA A Trp; LB 1 RPG & AK47 fire; LC XT618224 at 2045hrs; LD C&C ship landing in NDP
          hit by 1 RPG, crew alright, C&C destroyed, returned fire with organic weapons, gunships put in fire
          on XT618224 & 617230, small amount of AK fire; LE have regrouped, also received 5 mortar rds,
          negative damage to vehicles.
2215    SPOTREP: LA A10; LB mined section of Rt 7A; LC XT599222 at 1730hrs; LD 4 mines, 1 booby
          trapped grenade, 20 meters apart in a 100 meter section; LE blown/craters 4'x6', continuing mission.
2215    A10 radar status: left 2800 mills, right 4740 mills, center 3770 mills at 1000 meters.
2300    LTC McGowan said A Trp's mission for tomorrow is altered; 1st get stuck tanks unstuck, then make
          sweep of area where C&C was hit, then continue sweep to north.
Dec 1










Dec 2






Dec 3








Dec 4






Dec 5


Dec 6






Dec 7






Dec 8










Dec 9











Dec 10







Dec 11











Dec 12


















Summary: Troop continues NDP defense in Hobo Woods around burnt helicopter & stuck tanks. During
day A Trp buried Huey remains, rotor blades, and left front door. VTR arrives in the afternoon &
retrieves 3 stuck tanks.
LTC McGowan requests that we check w/appropriated personnel to insure that A Trp's ammo
resupply will be brought in this morning.  LT See reports that it is laid on w/Chinook for 0700hrs.
Stand-to complete.
A Trp reports civilians coming toward their location carrying a stretcher.  CPT Moss will move them
back to the compound.
A Trp will have problems with the soft ground in this area, destroying them.
A Trp at XT618228.
A Trp sweep at XT620228.
A Trp sweep at XT616226.
2nd Bde 25th ID Cdr is at A Trp's location.
1st PLT A/3/4 Cav is getting their stuck tanks out.  2nd PLT A/3/4 Cav will conduct their sweep. 
3rd PLT A/3/4 Cav will sweep road down to get the VTR.
A Trp is in the vicinity of Trung Lap.
3rd PLT A/3/4 Cav starting on sweep.
3rd PLT A/3/4 Cav at XT610223.
A Trp at XT610223.
SPOTREP A36 has a booby trap & spider hole at XT609221 consisting of concertina wire w/Chicom grenade & 155mm rd.  Wire across road. Spider hole freshly dug 2' x 5'.  Blown in place.
3rd PLT A/3/4 Cav is at XT604222.
A Trp is at XT604249.  Has a tank stuck again.
A Trp tank is unstuck, ground is soft.
3rd PLT A/3/4 Cav is at XT592228.
A Trp is turning around and starting back.
A Trp reports that the ground is to soft to move any farther at XT608253.  Will conduct
reconnaissance in force in local area.  Got stuck vehicles out.
2nd PLT A/3/4 Cav closed into NDP.  VTR  is on way to XT608253 to get stuck vehicle out.
A Trp reports VTR is at XT608253 to pull out stuck vehicles.
Resupply is at CP 17 the location of A Trp NDP.
A Trp reports that 2 tanks are unstuck but the other tank will require the track to be broken.
A Trp reports that they still have 1 tank stuck.  1 tank they have the track laid out & are preparing
to fix it.
Resupply to A Trp is leaving at 1530hrs.
A Trp reports all vehicles are out and will make the night location at XT612232.
A Trp location is XT625229.
A Trp reports resupply & the A30 elements are closing into the NDP.  Holding the POL vehicle at
the FSB.
SP4 Messler putting radar equipment on C&C ship to be flown out to A Trp.
All A Trp trains have closed the NDP at this time.
Dustoff for A Trp, urgent priority.  Litter patient at XT614234.  Man stepped into a punji pit,
has wounds in leg & thighs.
Dustoff complete, patient ambulatory (not litter).  Cdr's C&C ship is bringing in.
A Trp radar scan: right 400 mils, left 4600 mils, center at 5500 mils at 1000m.
A Trp's ambush patrol departing perimeter for XT611231.
Ambush patrol is in position.
A Trp FO reports only time he can fire DEFCONs is when movement is spotted.
A Trp radar new scan: right 3840 mils, left 2870 mils at 1000m.
SPOTREP 1st PLT A/3/4 Cav reports movement at XT616230 observed with starlight scope &
spotted on radar.  130m from NDP, engaged w 50cal fire.  Negative return fire & results.

Summary: Squadron directed A Troop to return to Cu Chi base camp for maintenance. 1st Platoon
tank accidentally fired off a 50 cal round inside the gate.  On arrival in A Troop motor pool, Squadron wanted to know who fired 50 cal round?  Maintenance conducted and resupply. Day ends with
barbeque in Troop area of base camp. (photo) Lt See assigned to Division.
A Trp, Squadron, & LTC McGowan all on new frequencies & callsigns.
A Trp stand-to is complete
A Trp reports engrs moving out.
SPOTREP A20 reports mine at XT612224 Russian type pressure detonated, 30 lb & wrapped in green plastic on side of road.  Underwater, blown in place leaving crater 4.5' in diameter.
A Trp reports road is impassable due to mud at XT612224.  Rice paddies on each side.  Will use route
used by resupply last night instead.
SPOTREP A Trp has 4 suspected mortar rds at XT615227 from 60mm mortar at a safe distance.  Can
not determine the source.
SPOTREP A Trp has a mine at XT595221 in water, estimated 40lb, blown in place.
SPOTREP A Trp reports a shape charge at XT595221 estimated 50lb, blown in place.
SPOTREP A30 has 2 tunnel entrances at XT644217 camouflaged 2.5' in diameter & covered with large
tree limbs.  1 entrance is 4' diameter.  75' along tunnel is concrete water pipe.
SPOTREP A36 reports trench & tunnel at XT642219. Trench is 15' long & 3' wide.  3 sheets of steel
plate 4' x 3' covering tunnel entrance.  Tunnel is 17' long.
A Trp is moving out.
A Trp is lining up and picking up the VTR.
A Trp reports civilians moving back into village at XT605224.  A Trp is at XT593205.
2nd PLT A/3/4 Cav is at XT576186 (intersection of QL7a & Hwy 1).
A Trp elements at CP 17.
A Trp is at CP 16.
A Trp elements entered Ambush Gate.  All weapons cleared & all personnel accounted 4.
All of A Trp has closed base camp.
A Trp internal security check complete. 18 APCs & 6 tanks for reaction force.
A Trp directed to hand receipt a PPS-5 radar to B Trp.
A Trp FO, LT Klar, reports operation will start at 0630hrs tomorrow. 

Summary: A10 elements start at XT6917 sweep to XT684197 northern limit XT686182.  A20 elements start at XT6817 weep from XT668188 northern limit XT673175. A30 elements starts at XT684171
sweep to XT675195 northern limit XT679179. Reconnaissance-in-force mission Northeast of Cu Chi
 in Filhol Plantation.
Lt Ayres in field as 3d PLT Leader. Lt Hogue reassigned to Troop XO.
A Trp departs the motor pool.
1st PLT A/3/4 Cav clear Ann Margaret gate.
3rd PLT A/3/4 Cav clear Ann Margaret gate.
2nd PLT A/3/4 Cav clear Ann Margaret gate.
SPOTREP A20 reports 1 60mm mortar rd at XT675174 US type buried in ground.  Blown in place.
A Trp reports A10 elements at XT683188, A30 elements at XT682178, & A20 elements at XT678172.
SPOTREP A10 reports tunnel at XT682170 with 2 openings in gully.
1st PLT A/3/4 Cav at XT684198 (Hwy TL8a).
2nd PLT A/3/4 Cav at XT683167.
A20 elements at XT682198 & A30 elements at XT685174.
SPOTREP A20 has 2 tunnels at XT675174 with 2 openings 30m apart.  Burned area, old tunnel.
SPOTREP A Trp reports area of an old camp fire at XT686180.  Several days old, 1 can of Vietnamese food, maybe 2 people used site.  No tracks or trail & 2 pieces of paper with writing.
SPOTREP A Trp reports punji pit at XT686181 with 20 stakes, 1 month old.  Blown in place.
SPOTREP A20 found dud 4.2" HE rd on ground at XT678177.  Blown in place
A10 elements at XT678177, A20 elements at XT685182, & A30 elements at XT687183.
A10 elements at XT687179 & A20 elements at XT676179.
A30 is holding his element.  Checking a tank roadwheel.
SPOTREP A38 has tunnel complex at XT681188.  About 1 month old not recently used.
A30 has started his reconnaissance-in-force again.
A30 elements are at XT679194.  Having trouble with tank, will leave it secure & continue mission.
A20 elements at XT677179 & A10 elements at XT686175.
SPOTREP A34 has a footpath at XT676204 lined with holes 2' x 4' x 3'.  Destroyed holes.
A30 elements are at XT675196 and has a stuck tank.
SPOTREP A Trp has a tunnel at XT688172 that is 4' x2' x 4'.  Has been caved in by a bomb.  Fresh footprints in area.
A Trp reports A30 elements still holding with down tank at XT675196.  A20 elements at XT688170,
& A10 elements at XT672186.
SPOTREP A30 has a tunnel complex at XT677197 about 1 week old and not completed.  Destroyed
with a tank.
A10 elements at XT681187, A30 elements in night laager at XT681179 pulling maintenance on
vehicles & weapons, A20 elements still with stuck tank.
A10 elements in NDP at XT681179.
A Trp reports all A20 elements except for 2 vehicles have closed NDP location.
A20 reports tank is pulled out and is headed to their NDP.  A30 elements will be in NDP at XT683181.
A55 (VTR) is on its way to pickup resupply.
A20 elements NDP is at XT673175, Trp Hq will be at XT681180, A10 & A30 elements closing.
A Trp internal security check in base camp is negative.
2nd Bde 25ID directs that all Trps must complete RIF missions east of the 68 grid by 1200hrs tomorrow.  Must be west of the 68 grid line due to Arty & ARVN show of strength.
A Trp radar scan: left 4800 mils, right 3800 mils, center 5770 mils, at 1000m.
A Trp given permission to test fire 50cal & M60/coax on azimuth of 290 degrees but must keep it low.

1st PLT & 2nd PLT A/3/4 Cav enroute to meet 3rd PLT A/3/4 Cav.
A10 elements starting RIF.
A20 elements starting RIF.
SPOTREP A20 found 3 Chicom grenades at XT688188 in a triangle pattern on the ground.  Blown in place.
SPOTREP A10 found anti-personnel mine at XT685173 on an old trail, mine was deteriorating. Blown in place.
A10 at XT686175 & A20 at XT688189.
A Trp combat trains coming in.
SPOTREP A13 has 5 ox carts pulled by 2 oxen each at XT681170 with 8 personnel.  Are local wood gatherers.
A10 is at XT680168 & A20 is at XT681189.
SPOTREP A15 has a tunnel complex at XT679166 running southeast to northwest 20m long.  Several tunnels are caved in.  Continuing search.
A Trp at XT673176, breaking for chow.
A10 element has a tank that hit a mine.
A Trp starting RIF.
SPOTREP A20 found Chicom grenade at XT678175 on the ground. Blown in place.
A10 is at XT677187 & A20 is at XT678175.  A30 in motor pool coming out w/resupply run.
A20 is at XT679172 & A10 is at XT676184.
A10 & A20 elements located at XT678177 pulling maintenance and checking weapons.  A30 is still
in motor pool.
A Trp locations for tonight are: Squadron HQ, Trp Hq, A10 & A20 at XT688184, & A30 at XT682178.
A Trp has 4 APCs, 1 VTR, & 1 POL truck for base security.
A20 elements in night laager.
A Trp base camp security check is negative.
A Trp is in night laagers.
A Trp radar scan: right 1000 mils, center 1970 mils, left 2940 mils at 1000m.
A Trp has 2 LPs in position.
Trp (-) in NDP in the Fihol Rubber Plantation with Squadron HQ at grid XT688184. After daybreak
Troop conducts 2 platoon size searches at grids XT680200 and XT687179. Troop establishes
3 NDPs in the southern Fihol.

DISCOM reports incoming mortar fire at Cu Chi.
A Trp has negative incoming in troop area & negative damage.
A Trp reported seeing a few flashes on a 280 degree azimuth from his NDP.  Stand-to complete.
A10 elements & A30 elements on the move.
A Trp made it to Cu Chi with the down truck.
A20 elements on the move.
A10 & A30 are at XT689194. A20 is at XT680176.
A Trp NDPs will be A10 & Trp Hq at XT680182.  A20 at XT671173 & A30 at XT688184.
A Trp locations are A10 & A30 at XT670216. A20 is at Xt683197.
SPOTREP A Trp has a dud 155mm rd on ground at XT688172.  Blown in place.
SPOTREP A20 has a live 105 Arty rd on ground at XT676187.  Not booby trapped, blown in place.
SPOTREP A30 has a spider hole at XT687167, 2' x 2' with no recent activity.  Blown in place.
SPOTREP A20 has a 105 Arty rd on ground at XT678173 with shipping plug in it. Blown in place.
A Trp is at the A6's location to eat chow at XT671178.  After chow will move to NDP.
A Trp night laagers will be A10 at XT680182.  A20 at XT671173 & A30 at XT688184.
A Trp line-up east to west is A30, A20, & A10.
A10 & A30 are in their NDPs.
A30's LP is on a 220 degree azimuth 50 m from his NDP.
A20's LP is on a 120 degree azimuth 400 m from his NDP.
A20's LP is in position.
A30's LP is in position.
A Trp radar scan: left 0 mils, center 970 mils, right 1990 mils at 1000 m.
All LPs & ambush patrols are in position.
INTELREP from 2nd Bde 25th ID: combined VC/NVA forces expected to make a main offensive tonight
or tomorrow morning.  Be watchful for lights and movement in AO.

Summary: A Trp in 3 NDPs in southern Filhol Plantation.  Will conduct platoon RIFs.
A Trp elements are on the move.
SPOTREP A6 found chinaware at XT678176 consisting of 10 to 15 plates.  No activity, appears to be
old village.  Picked up some plates.
A10 elements are at XT667173. A20 elements are at XT690195. A30 elements are at XT677165.
Late entry: A Trp RIFs are A10 XT667173, XT668178, XT674178, XT670190, to XT690196; A20 XT674190, XT690200, XT676180, XT685185, to XT690196; & A30 XT672178, XT675166, XT682166, XTXT675180, XT690190, XT683190, to 670180.
SPOTREP A16 has tunnel at XT669169 2.5' x 20' looks fresh.
SPOTREP A30 has spider hole at XT680171 2' x 3' overhead cover, looking east.  Destroyed.
SPOTREP A30 has 2 fighting positions at XT676177 2' x 3' w/small overhead covers.
SPOTREP A Trp found M26 US hand grenade at XT677176 in a burnt out area, destroyed.
SPOTREP A30 has 3 old fighting positions at XT676177 2' x 2' w/overhead cover.  Old bunkers.
SPOTREP A Trp has dud 105mm Arty round at XT676177 laying on ground. Blown in place.
SPOTREP A30 has dud 155mm HE Arty round at XT675184, blown in place.
A Trp reports all elements on the move.
SPOTREP A6 has old ARVN positions at XT678185 consisting of 10 foxholes, destroyed with APCs.
A Trp is returning to base camp.  They have to hook up down APC before returning.
A Trp entering base camp.  All personnel accounted for and all weapons clear.
A10 elements and A20 elements have closed into the motor pool.
A30 elements have entered Cu Chi.  All A Trp elements have closed.
A Trp elements have all closed into the motor pool.
A Trp has a fire in the washroom next to the mess hall.  It is out at this time.
A Trp has left the net.
A Trp status is 115 personnel and 28 vehicles.  Internal security check is negative.
Troop located in Cu Chi. Troop moved from Cu Chi to south of An Duc vicinity grid XT544150 and establishes NDP.

A Trp & B Trp given tentative SP time of 0745hrs with B Trp leading.
A Trp departing main gate.
A Trp elements clear CP 16. Trail element approaching CP 16.
A Trp lead element has closed up to B Trp's trail.
A Trp lead at CP 9.
A Trp has pulled in at CP 8.
A Trp is starting to move south down Rt 7A.  Road has been swept to 1st outpost by another element.
A Trp is down to the old laager site.  No one is there, they are continuing sweep.
A Trp sweep is 1500m from CP 17.  Sweep is moving slow.
A Trp is 4500m down Rt 7A.
A Trp is 500m from FSB Stewart.
SPOTREP A30 has bunker complex at XT502154 consisting of 5 bunkers 4' x 6', 2 covered with logs. 
Signs of recent use, blowing in place.
A Trp located at XT504160 (intersection of RT 6A & RT 7A).
A Trp reports 2 of his platoons will go back to night laager.  One platoon is working on a track and will arrive later.
A Trp (-) & Trp HQ at XT542151.  PLT strong point at XT550176.  Ambush patrol at XT555170.
A Trp reports an ambush patrol fro Cu Chi Sub-sector will be at the location of his A10's night laager. 
It was cancelled.
A Trp ambush patrol is at XT542148.
A10 elements at CP 17.
A10 element is in position.  CPT Skiff cleared H&I fires from 55 to 62 grid north of E/W grid 30.
A Trp will be test firing on zero degree azimuth.
A Trp listening post and ambush patrol in position.
A20 element is in position.
Troop A (-) is located vicinity XT542151. At first light, 1st platoon sweeps Hwy 1 South and West to vicinity XT503153 and returns to NDP.

Stand-to complete.
A30 elements will sweep from their night laager to XT503161.  A10 elements will sweep from CP 8 to
CP 16.  A20 elements will be held at their NDP.
A10 elements at CP 17.
A10 elements at CP 9.
A10 elements at CP 16
A10 elements at CP 9.
A 10 elements are 300m off of the MSR.  A30 elements are 300m outside of the main gate.
A10 elements have completed their sweep and are going back to the FSB to eat chow.
A30 elements are at XT533150 and will conduct a RIF.
SPOTREP A30 reports booby trap at XT528150 on entrance to hooch with wire across entrance. 
Threw a grenade into hooch.  No secondary explosion.
SPOTREP A30 has punji pits at XT528149 consisting of 7 pits 1.5' x 1'.  Look new, about 1 week old.  Destroyed with grenades.
SPOTREP A30 has a dud 60mm mortar illum rd on ground at XT528153.  Possible booby trap, blowing in place.
A30 elements at XT505150.
SPOTREP A30 found propaganda leaflets at XT521152 on southern edge of RT 7A.  5 leaflets on sticks.  Was not booby trapped, A30 has the leaflets.
SPOTREP A30 found plastic mine at XT514153 12" in diameter.  Found by probing, was under 2" of dirt.  Blown in place.
SPOTREP A30 has a mine made of RPG round at XT514153 standing upright, blew in place.
A Trp requests dustoff.  Track hit a mine.  Has 1 WIA.  Dustoff ETA is 10 to 15 mins.
SPOTREP A30 track hit a mine at XT513156. 40 to 50lb mine, track is combat loss, one WIA with slight head wound.  Crater was 4' x 4', continuing sweep.
SPOTREP A30 found AT mine at XT514154 12" in diameter, blown in place. Crater is 5' x 5'.
A Trp is escorting dump trucks up RT 19.
Dustoff is complete.
A Trp at XT506154.
A Trp taking a break for chow for Engrs.  Permission given by SFC Kief.
A Trp is at XT506154 waiting for Rome plow to come up.
SPOTREP A30 has tree stumps in road at XT504154 blocking traffic.  Checked by Engrs, no booby traps.  Cleared road & continuing mission.
A Trp request to recon by fire at XT507153 & XT503157.  OK from MAJ Harrison.
SPOTREP A30 has 1 mine, brush, & 3 dirt mounds at XT503158 with sign in middle of each mound. Engrs blew mine and cleared brush.
A Trp was pulling a downed tank and hit a pedal bicycle/cart.  No injuries but minor damage to bicycle.
A Trp burning dried grass before departing for NDP.  Rome plow going back.
A Trp in trp laager at this time.
A20 elements in night laager.
A Trp internal security check base camp.  18 EM for reaction force.
A20's ambush patrol in position.  A10's ambush patrol enroute to position.
A Trp radar scan: left 1400 mils, right 5940 mils at 1000m.
A10's ambush patrol is in position.

Summary: Troop NDP located at grid XT542151. 2nd platoon & RF/PF unit establishes blocking position
vic XT560155. 1 platoon RF/PF and 1st platoon sweep through Thai My to vicinity grids XT553168 and XT570163. Troop  (-) sweeps Hwy 7A fromm junction at Hwy 1 to vic XT503161 and returns to NDP.
2nd Bde 25th ID directs that: 1) Each FSB in grassy areas will start a control burn program to clear
fields of fire.  2) Every man on a mine sweep team or on mine sweep security will wear a flak jacket.  Point man  on all operations will wear a flak jacket.  Point men on sweep security will move off the
road on both sides a such a distance that they will be able to detect command detonated wires. 
They will carry wire cutters and be at a distance that the VC normally set out command detonated claymores.
Incoming rds into base camp.  Negative incoming rds in the A Trp area.
A Trp stand-to complete.
A30 elements will sweep from night laager to XT503161.  A20 elements will sweep from CP 8 to CP 16
and then back to the night laager.  A10 elements will conduct a RIF with 1 Co RF/PF from Cu Chi Sub Sector.
A Trp needs urgent dustoff.  1 US litter, man woke up with stiff back.  Dustoff 162 ETA in 10 to 15 mins.
A30 elements have left gate and started on their RIF.
SPOTREP A20 reports road obstacle at XT559168 consisting of shrubs & grass covering 50 meters of
the road stacked 2' high.  Clearing & continuing mission.
A30 elements at XT540152
A Trp has swept from night laager to CP 17 on RT 7A and CP 8 to CP 16 is swept on Hwy 1.
A Trp requests dustoff for 1 US who stepped on booby trap at XT545153.  Has frag wounds to leg and foot.  ETA is 10 to 15 mins.
A Trp dustoff is complete.
SPOTREP ARVNs attached to A Trp have 1 detainee at XT552143.  Eligible male 36 years old taken back to ARVN compound. Continuing mission.
A Trp has completed RIFs with RF/PF forces.
A30 element sweep is at XT510154.
A10 elements on the move.
SPOTREP A30 has RPG rd buried as a mine at XT509154 pressured detonated & wrapped in plastic. 
Blown in place.
A Trp locations are A10 & A30 elements at XT505154.  A20 elements outposting & fixing laager.
SPOTREP A30 has a tunnel & spider hole at XT505155.  Hole is 2' x 2' x 2', threw grenade in which exposed the tunnel.  Tunnel too small to explore.
A10 elements at XT510150 will recon by fire to his left flank.
SPOTREP A17 has tunnel at XT500350 4' x'3' x5' with recent digging and sandal prints in area.
SPOTREP A30 has AT mine at XT530151 that 2 vehicles had passed over.  3rd vehicle spotted the
mine in the dust.  Planted after earlier sweep, 12" in diameter, Chicom type blown in place.
A10 & A30 elements have closed last night's laager position.
A Trp ambush patrol will be at XT541150.
A30 elements on move to night laager.
A30 element at CP 17.
A30 element in night laager.
A30's night laager is at XT549171.  Ambush patrol is on 165 degree azimuth 145m out.
A20's ambush patrol is departing for his ambush position.  A30's ambush patrol is in position. A30 radar scan: left 4000 mils, right 5946 mils, center 4970 mils at 1000m.
A20's ambush patrol is in position
SPOTREP A30 reports movement at XT552175 near ambush patrol.  Will fire a flare &recon by fire.
SPOTREP A30 ambush patrol spotted movement at XT552170 20m north of their position with the
starlight scope. Returned w/all organic weapons & sweeping area.
A Trp reports swept area of contact with negative results.  Will sweep again at first light.

Summary: Troop NDP vic XT543151 south of An Duc. 1 platoon established night strong point vicinity XT550193. At first light, 1 platoon sweeps Hwy 1 from XT543177 to vic XT503161. Troop (-) conducts Reconnaissance-in-Force vic of Thai My. 1 platoon Trp A conducts combined Reconnaissance-in-Force and search with 636th RF vic XT551177. Troop secures FSB STUART.
A Trp stand-to complete
A Trp has dismounts down at this time.
A Trp has started RIFs at this time.
A10 elements at XT540150.
SPOTREP A30 found propaganda leaflets at XT557168 attached to stakes, will collect to bring in.
A30 elements at CP 17.
A30 elements at CP 9.  A10 is at XT538150.
A20 elements have completed RIF.
A30 is at CP 9.
A10 elements are at XT533150.  A30 elements are at CP 17 south to the night laager.
A30's night laager is at XT575163.  Ambush patrols are at XT570164 & XT570160.
SPOTREP A10 has anti-personnel mine at XT530150 Chicom type 10 to 15m off of road, pressure detonated.  Will blow in place.
A10 elements at XT471150.
A Trp moving to pickup ARVNs for the RIF.
A10 elements at XT527151.
A Trp located at XT515152.
A Trp requests dustoff for 3 VN ambulatory at XT555195 multiple frag wounds.  Dustoff 156
ETA is 10 mins.
A Trp is located at XT562175.
A Trp requests dustoff for 1 RF/PF tripped a booby trap is a litter patient.
Dustoff 156 is completed.
Dustoff 155 ETA is 10 to 15 mins.
Dustoff for A Trp is completed for 1 ARVN litter patient.
A10 elements are at XT506154.
A Trp feels that the RIF is nearing completion.  CPT Moss feels action was result of RF blocking force moving in prematurely.  They had the VC in the open at XT563177.  Result was 2 VC body count, 2
AK47s & 10 detainees.  RFs' had 7 WIA & 1 KIA.  Preparing to pickup ARVNs and take back to their compound.
A10 elements at XT504154.
A Trp is dropping ARVNs at their compound.
SPOTREP A35 hit AT mine at XT535151 blowing 1 roadwheel and some suspension damage, no injuries
& towing to NDP.
Reference A35's report, mine was 40 to 50lb.  Tank can be repaired.
A Trp will knock down & burn hooches that are not being lived in.
A Trp ambush patrol at XT537153.
A10 requests test fire on 190 degree azimuth.
SPOTREP A10 spotted 1 person on road at XT535151 engaged with M60 & M79 fire.  Negative results.
A Trp given neg clearance to fire down road by LTC McGowan.
A20's ambush patrol departing for AP position.
A Trp reports that the attached engr sweep team needs a mine detector.
SPOTREP A6 3 lights mvoing north to south 50 to 100 meters apart at XT532150.  Request permission
to engage.  Approved by LTC McGowan.
Lights went out before A Trp could engage.  A10's ambush patrol is in position.
A Trp radar scan: left 4000 mils, center 4920 mils, right 5940 mils at 1000m.
LTC McGowan informs A Trp that they will have a secondary mission to sweep Hwy 19 w/1 platoon and escort VTR to tow vehicles at XT462245 back to Cu Chi.
INTELSUM II Field Force & ARVN 3rd Corps reports VC will launch a general offensive tonight.

Summary: Troop (-) NDP is located vicinity XT543151. 1st platoon night strong point vicinity
XT575163. At first light, 1st platoon sweeps Hwy 1 from vicinity XT543177 to Cu Chi and conducts local RIFs vicinity of Thai My. 2nd platoon Troop A conducts combined RIF with 1 platoon 636th RF Co to vicinity grids XT525150 and XT502160 and returns to NDP.  1st PLT conducts secondary mission
to pickup
VTR & recover the burned APC & tank from B Trp.
A20's ambush patrol returning to perimeter.
A20's ambush patrol inside perimeter.
A Trp stand-to complete.
A10 elements will sweep from night laager on main road to CP 16 and then to the main gate to pickup VTR.
A Trp night positions for tonight are XT550173 with ambush patrols at XT555169 & XT538141.
A10 element is on the move and at CP 9.
A Trp sweep is at CP16.  2 vehicles have been sent to the main gate.
A10 element is at CP 9.
VTR is now with A Trp and has departed.
A Trp elements with VTR is at CP17.
VTR is now at FSB Stewart.
A10 element is at XT471230 (Trang Bang).
A10 with the VTR is at B Trp's tank that hit mine.
A Trp will be able to use a lowboy from FSB Stewart to pick up APC.  CPT Lavine gave permission
to use it.
A Trp escort veh enroute to FSB Stewart to pickup lowboy.
A Trp sent 2 APCs to check on tank heading for Cu Chi.  Found 2 Engr dump trucks at An Duc that
had collided.  2 APCs securing them.
A10 elements securing VTR.  A10 eports finding parts of 105 or 155mm round underneath tank.
 Needs a lowboy.
A10 elements at CP 8.
A Trp starting on his RIF at this time.
A Trp RIF at XT537148.
A Trp told to move one of his platoons to CP 8 and to stand-by for a mission.
A10 elements at CP 17.
A10 element is at CP 8.
A6 is moving to CP 8.
A Trp new mission is to check out 3 hooches at CP 8 believed to contain 8 possible VC.
A20 elements at CP 17.
A20 & A30 elements are in the NDP.
SPOTREP A Trp reports his elements have reached objective at XT560170.  Has ARVNs and is
checking the houses.  Negative contact, negative mines or booby traps.
A Trp has checked out the 3 hooches.  Locals say VC have left area.  A Trp moving back down
road to NDP.
2nd Bde 25th ID directs that 3/4 Cav provide security for downed lowboy and dozer outside of
main gate at XT615135.  MAJ Harrison, S3, directs A Trp to set-up NDP at this position.
A Trp requests dustoff for one VN child at XT543150 with frag wound to head.  Disregard the child's father took her back.
A20 elements departed for new NDP location.
A20 reports his element is approaching the new NDP.  Sees no lowboy but a dozer is in the ARVN compound.
A Trp is to secure the dozer at the ARVN compound.
A20's NDP is at XT618133 and his ambush patrol is at XT612135.
A20's ambush patrol is in position.
A10's ambush patrol is in postion.

Summary: Troop NDP located vicinity XT543151. 1st platoon sweeps Hwy 1 from vicinity XT544177 to vicinity XT628126. Troop also sweeps Hwy 7A from NDP to Hwy 1; then West to vicinity XT503161.
Upon completion of sweep, conducts RIFs vicinity Thai My and returns to NDP. Troop and ARVNs detonated an unknown type booby trap, resulting in 1 ARVN KIA, 2 ARVNs WIA (dustoff), 1 US
WIA (dustoff). (photo)
2nd Bde 25th ID directs that all troops will carry protective masks on all operations.
A Trp stand-to is complete.
A20 element is at CP 16.
A20 element is at CP 9
A20 element is at CP 17.
A Trp has picked up Engrs at FSB.
A20 elements are at XT484210.
A20 elements are at XT468232.
A Trp is at CP 20.
A20 elements are at XT467236.
A10 elements are at CP 17.
A20 elements have started sweep at XT462246.  A Trp's combined RIF is at XT591150.
SPOTREP A20 found homemade mine at XT460250 consisting of 2 sticks of bamboo filled w/explosives.  Blown in place.
A Trp combined RIF at XT539150.
Dustoff need at XT541151 2 litter & 1 ambulatory all ARVN from booby trap.  Dustoff 155 ETA in 10 min.
SPOTREP A20 booby trap at XT538151 triggered by ARVN's checking hooch.  1 ARVN KIA, 2 WIA
& dustoff called.
Dustoff 155 is complete.
Engrs checked area around mine that A20 reported at 0940hrs and found a 50lb mine made from a
green powder in a sandbag.  Also had a bandoleer w/white powder wrapped in plastic with wires
 running into powder.  Was positioned to blow under left tracks of vehicle on road.
Reference the booby trap that wounded the ARVN, booby trap was a grenade sitting on a 18" pole
in pile of dirt outside of hooch.  Triggered by a trip wire.
A Trp will not secure FSB Stewart but will have night laager on MSR.
A Trp will establish a platoon strong point at the Football field, another platoon strong point at FSB Stuart, & another platoon strong point at CP 8.
A30 elements are on the last leg of their move.
A30 elements are at CP 17.
SPOTREP A30 found spider hole at XT514193 relatively old 2' x 6' deep & empty.  Destroyed by running over it.
A30 elements are at XT510190.  A10 elements are on the move again.
SPOTREP A30 has 2 spider holes at XT514195 2' x 2' x3' with nothing inside.
SPOTREP A30 has punji pit w/stakes & VC stretchers at XT514195 consisting of 8" rusty spikes on the stakes & stretchers of cord hammocks and 2 pieces of bamboo.  In hooch covered with trash. 
Destroying hooch and bringing in articles.
SPOTREP 3rd PLT A/3/4 Cav found 5 punji pits at XT514179 2' x 5' consisting of stakes covered by
straw w/10 stakes in each pit.  Destroying with grenades.
Warning Order from MAJ Harrison, S3, directs A Trp to be prepared to escort a convoy from FSB
Stuart to FSB Pershing.
SPOTREP A30 has bunker at XT513192 3' x 4' x' 3' w/overhead cover made of branches.  1 US steel helmet inside.  Blowing in place.
Convoy will consist of 6 trucks.  They will contact A Trp on A Trp's push.
A Trp located at XT512193.
A Trp is moving medics, NCS, & maintenance tracks to FSB Stuart.
A Trp has picked up the convoy and will be moving shortly.
A30 elements are at XT491220.
A Trp located at XT501249.
A Trp located at XT507260.
A Trp moving to night locations at this time.
A30 elements at XT490215.
A30 elements are at FSB & A0 elements are at CP 8.
From MAJ Harrison, S3, to CPT Moss: 2/12 IN Bn is OPCON to 3/4 Cav.  CPT Moss is to coordinate
with 2/12 IN Bn on defense of FSB Stuart.
From MAJ Harrison, S3, to A Trp:  Have the platoon at CP 8 move to CP 17 and escort LTC McGowan
& tracks to FSB Stuart.
A10 elements departing FSB Stuart to CP 17.
A10 elements at An Duc.
SPOTREP A Trp/2-12 IN has 4 flashlights at XT501194 on azimuth of 235 to 240 degrees at
450 to 500m.  Illuminating and recon by fire with small arms.
SPOTREP A30 has moving lights at XT477197 south of the MSR moving in and out of the woodline. 
3 to 5 lights, continuing to observe.
A Trp reports the lights seen at 1934 hrs are now stationary on 120 degree azimuth at 700 to 800m. 
One appears to be signaling the others.
Warning Order from MAJ Rigby, 2nd Bde 25th ID, to MAJ Harrison, S3:  One PLT of A Trp will proceed
the logistical convoy to the Big Rubber, sweep & outpost the Big Rubber at CP 3 and follow the convoy southward to FSB Stuart.
A Trp radar scan: left 0000 mils, center 0970 mils, & right 1990 mils at 1000m.
A Trp reports 2/12 IN Bn ambush patrol is located at XT497194.
A10's ambush patrol is at XT555169 & listening post on 170 degree azimuth at 100 meters.  A30's
ambush patrol is at XT483197 & listening post on 260 degrees at 50m.
Dec 13
Summary - Troop establishes 3 NDPs. 1 platoon located vicinity XT482195; 1 platoon located FSB STUART 
          vicinity XT501194; 1 platoon strong point located vicinity XT543177. At first light, platoon at
          strong point sweeps Hwy 1 from Trang Bang to 2nd Bde boundary vic XT459250.
0540    A Trp stand-to complete.
0640    A Trp at CP19.
0650    A30 elements at CP 20 and starting back.
0700    A Trp sector clear from CP8 to CP20.  A30 elements at CP 19.
0758    A20 elements at CP 19.
0810    SPOTREP A Trp needs dustoff at FSB Stuart for 1 VN child w/frag wounds in stomach.
0822    A20 elements are at XT470232 and putting dismounts on the ground.
0900    A Trp starting move to the Rubber.
0917    A20 elements at XT467235.
0919    A Trp at CP 4 and in contact w/1st Bde 25th ID.
0930    A20 elements at XT465239.
0951    A20 elements at XT460250.
1035    SPOTREP A18 found mine at XT459251 in grass at side of road.  It consisted of a 105mm rd with
          pressure release wire across road.  Blown in place.
1102    A Trp elements in the Big Rubber are to be on 4/9 IN Bn's freq.
1105    ARVNs are disorganized.  A Trp is sweeping w/them to check area.  1 ARVN wounded from a sniper.
1114    A Trp is moving on the sweep.
1130    SPOTREP A30 has 3 spider holes at XT513194 appear fresh.  Have destroyed.
1152    A Trp reports blood trails at XT514194.  Still checking area.
1155    2nd Bde 25th ID is airlifting a dog team to A Trp.
1210    A30 elements are at the western boundary of the RIF.  Making a turn on 180 degrees.
1225    A10 elements have completed sweep of Hwy QL 19.  A30 elements at XT511192.
1232    Dog team will be late.  Chopper was diverted to another unit that is in contact.
1305    Dog team on the way.  ETA is 10 to 15 mins.
1334    Dog team has arrived at A Trp's location.
1336    A10 elements are at FSB Stuart.
1420    A Trp can't go any further due to soft ground.  ARVNs are still following the dog team.
1438    A30 elements returning to road at this time.  ARVNs continuing to track VC, are cutting south
          toward road at this time.
1500    A30 elements are 400m from FSB, meeting with ARVN Advisors.  ARVNs found 1 stick of Flex-X, 1
          pair of Ho Chi Minh sandals, dogs will be picked up at the FSB at 1530 hrs.
1524    A30 elements at FSB Stuart.  A20 elements are at CP 4 held up by the convoy which has down
1528    Dog team is leaving the FSB, returning to Cu Chi on the convoy.
1555    A Trp on seep of MSR will pick up a track at the Main Gate and escort it to FSB Stuart.
          Will sweep from CP 16 to CP 20.
1620    A20 elements closed FSB except for 1 VTR towing tank and 1 tank excorting it.
1645    A Trp's 2 tanks & VTR are in FSB Stuart.
1758    A20 elements departing for night laager.
1817    A20 elements in position.
1835    A Trp ambush patrols are at XT555169 & XT478198.  3 LPs out also.
1839    MAJ Harrison, S3, to CPT Moss: A20 elements will be OPCON to B/2/12 IN who will laager at the
          same location.  A20 is to aid 2/12 with the defenses.
1900    A30 elements in position.
1955    From CPT Johnson, 3rd Bde 25th ID, the plt of A Trp that will sweep the Big & Little Rubber will
          come up on 4/9 IN Bn's push when they reach the 3rd Bde boundary.
2215    A Trp reports that they saw aircraft on an 80 to 50 degree azimuth, estimated at grid XT5620.
2230    A Trp reports that C/2/12 IN at the Trang Bang Bridge had a man drop some explosives on the
          bridge.  They have a dustoff coming in.
2245    SPOTREP C/2/12 IN has 1 WIA at the Trang Bang Bridge.  Man lit a new explosive and attempted to
          throw it into the river.  It exploded prematurely, blown off 1 arm and part of 1 leg.  Dustoff called in
          by 2/12 IN BN.
2304    Dustoff at Trang Bang bridge complete.
Dec 14
Summary: Troop establishes 3 NDPs. 1 platoon located Trang Bang football field; 1 platoon located FSB
          STUART; 1 platoon strong point vicinity XT543174. At first light, Troop A (-) swept Hwy 1 fm Cu Chi
          to 2nd Bde boundary and Hwy 19 from Trang Bang to 2nd Bde boundary, vicinity XT459250; 1 PLT
          provided escort for convoy to Dau Tieng and back. At 0920 hrs vicinity XT533195, 1st platoon tank
          detonated a 60 lb anti-tank resulting in 4 US WIA from Co C 2-12 Inf OPCON; major damage to tank.
0030    MAJ Harrison directed that B/2/12 IN & 1 PLT A/3/4 Cav sweep the high ground between XT539180
           to XT530120 after road sweep.
0532    A Trp stand-to is complete.
0553    A30 elements on the way to FSB Stuart.
0604    A30 elements at FSB Stuart.
0638    A20 elements on the move.
0641    A30 elements on the move.
0642    A30 elements at CP 7.
0653     A20 elements at CP 9.
0656    A20 elements at CP 17.
0700    A20 elements at CP 16
0725    A20 elements at CP 09.
0729    A20 elements at CP 17.
0740    A20 elements at FSB Stuart.
0806    A10 elements moving from FSB Stuart to pick up B/2/12 IN for RIF.
0812    A10 is at CP 8 coordinating w Cdr B/2/12 IN.
0832    1st PLT A/3/4 Cav is OPCON to B/2/12 IN.
0840    TM B/2/12 IN with 1/A/3/4 Cav starting RIF.
0841    3rd PLT A/3/4 Cav OPCON to 3rd Bde 25th ID.
0844    A Trp security for CP 19 & CP 20 departing.
0920    Request dustoff for 4 wounded from tank hitting mine at XT533195.  Urgent & area is secure.
          Dustoff 163 ETA 15 mins.
0920    SPOTREP A10 OPCON to B/2/12 IN had tank hit a 60 lb mine at XT533195. 4 dismounted infantry
          wounded.  Tank has damage to 3 roadwheels and 3 roadwheel     arms.
0954    Dustoff complete for A Trp.
1025    A Trp has brought tank to MSR.  Is preparing to pull it in.
1107    TM B/2/12 IN with 1/A/3/4 Cav  is at XT531200.
1120    SPOTREP A10 OPCON to B/2/12 IN has 4 old punji pits & 2 female detainnes. 1 has a mutilated ID
          card.  Will destroy punji pits.
1205    A30 elements released from OPCON.  Returned to A Trp control.  Will be going to CP 17 as security
           for the convoy.
1211    TM B/2/12 IN with 1/A/3/4 Cav  is at the pickup zone for B/2/12's airlift.
1241    A30 elements returning at this time.
1249    A Trp notified that pickup time for the RF/PFs is 1330 at CP 17.
1303    A30 elements have closed FSB Stuart.
1315    A Trp's RIF with the RF/PFs goes from CP 8 compound to the East.  Sweep this area to support
          operations to the North & East.  Mainly look for wounded from last night's contact.
1345    A20 elements picking up the RF/PFs and beginning the move.
1350    A30 elements at CP 4
1400    A20 elements at XT554174 lining up for the RIF.
1404    A10 has 2 vehicles departing FSB to outpost CP 19 & CP 20.
1435    SPOTREP A20 has 155 Arty rd on the ground at XT557176, not booby trapped or setup as a mine.
          Blowing in place.
1520    SPOTREP A20 has 105 Arty rd on the ground at XT560173, not booby trapped or setup as a mine.
          Blowing in place.  Also has a tank stuck 5m from the MSR>
1540    A Trp elements have completed the RIF.  Will have his VTR pull out the tank.
1631    A20 elements are at CP 08.
1642    A20 elements at the FSB.
1740    A10 elements departing FSB.
1741    From CPT Skiff, S3 Air, to A, B, & C Trps he must be notified before any vehicle is towed that will
          tear up the road.
1750    A Trp requests test fire from FSB Stuart on azimuth of 65 degrees.  Approved by 2nd     Bde 2th ID.
1759    A20 elements & A10 elements are in their night laagers.
1828    A10 requests test fire for his elements on azimuth of 240 degrees.  Approved by CPT     Skiff.
1940    A10 elements have completed their test fire.
2246    2nd Bde 25th ID to S3 and A Trp that there is terrorist activity reported at XT4620, 487188,
          489187, & 494192.
2247    A Trp from the Trang Bang Sector Advisor that there is contact at XT467198 w/2 VC plts, at
          XT467198 w/ 1 VC squad, & at XT473192 w/1 VC squad.
Dec 15
Summary: Troop establishes 3 NDPs. 1 platoon NDP located Trang Bang football field vicinity XT484196; 1
          platoon located FSB STUART vicinity XT501194; 1 platoon strong point vic XT550193. At first light,
          Trp A (-) swept and cleared Hwy 1 from Cu Chi to Northwest to 2nd Bde boundary.
0538    A Trp stand-to is complete.
0640    A20 elements on the move.
0645    A10 elements on the move.
0652    A20 is at CP 19 & A10 is at CP 17.
0655    A10 is at CP 9.
0700    A10 is at CP 9 & A30 is at CP 20.
0700    SPOTREP A10 has propaganda leaflets of VC origin in the Village of Bao Trai at XT615133 on about
          200m of the MSR.  Collected samples and continuing sweep.
0727    SPOTREP A10 has road obstacles of 1 log 10' long, brush, grass across MSR at XT533183 not booby
          trapped, clearing and continuing the mission.
0735    A20 elements closed the FSB
0740    A10 elements closed the FSB
0746    A30 is at XT484210.
0758    A10 elements departing FSB on the way to the Big Rubbers.
0810    A10 is at CP 19.
0815    SPOTREP A30 has road obstacle of debris, trash, leaves, & sticks width of road at XT481218 no
          booby traps.  Engrs clearing & continuing the sweep.
0830    SPOTREP A30 has a mine at XT480217 of undetermined size, green powder in plastic with 13'
          bamboo detonator powered by 2 flashlight batteries.  Blow in place, estimated at 80lb, left crater 6'
          deep.  Will need 5 tons of gravel to fill.
0917    A10 is at the Big Rubber, A20 is at CP 19, & A30 is XT4801217.
1025    A30 is at XT478219.
1145    A30 is at XT459251, just short of completing the sweep.
1244    A30 elements at FSB Stuart.
1313    Aerorifle Plt D Trp will remain in Cu Chi tonight and return to A Trp tomorrow.
1315    Aerorifle Plt D Trp completed 3 sides of cloverleaf sweep at XT650230 found fresh digging of spider
          holes and well constructed bunkers.  Trails show recent activity.  Looks like platoon size base camp.
1352    A Trp & Aerorifles entering area of aerorifle cloverleaf sweep.
1433    A Trp informed that 65th Engrs report that 2 of the 4 gravel trucks that were going to FSB Stuart
          were in an accident and will not make it.
1520    A10 closed FSB Stuart.
1525    A30 reports the 2 gravel trucks have dumped their loads and are returning to Cu Chi.
1825    A30 elements in night laager and wishes to test fire on 10 degree azimuth. Permission granted by
          2nd Bde 25th ID.
1830    A10 elements in night laager position.
2000    SPOTREP A30 has received automatic weapons fire at XT540178 of unknown number of rounds.
          Returned fire.  No fire at this time.
2029    A10's ambush patrol departing for position.
2030    A10's listening post departing for position 200 degree azimuth at 100m from night     laager.
2043    A10's LP is in position.
2115    A Trp radar scan: left 2900 mils, right 4840 mils, range 1000m.  A10's ambush patrol is in position.
2215    SPOTREP A30 has movement coming toward his position at XT545173 using starlight scope.
           Engaged with small arms.  Negative results and continuing to observe.
2340    The Commanding General is concerned about the numerous lights spoted in the Ho Bo & Boi Loi
           Woods.  Alert all personnel about this report and make sure trps in the field report & engage any
           lights or movement observed.  Believed sightings to be of a larger force in the area.
Mid Dec
          Troopers write message to VC on hootch near deserted hamlet. VC write message back to A Troop.
Dec 16
SummarY: Troop establishes 3 NDPs. 1 platoon NDP located Trang Bang football field; 1 platoon located FSB
          STUART vicinity XT501194; 1 platoon strong point FSB STUART.  At 2215 hrs vicinity XT545173
          Troop observes movement with starlight scope and engaged with SA; neg results. At 2008 hrs
          vicinity XT543178, Troop receives automatic weapons fire; returns fire; neg casualties, VC losses
0540    A Trp stand-to is complete.
0635    A30 elements sweep from night laager to CP 16 then back to FSB Stuart.  A10 elements sweep from
          night laager to CP 20 and then back to FSB Stuart.  A20 elements sweep Hwy QL 19.  A10 elements
          to conduct RIF vicinity XT5119.
0636    A10 & A30 started their sweeps.
0657    A10 elements at CP 20.
0645    SPOTREP A30 has road obstacle of brush & dirt piles on road at XT557169.  Cleared & on the move.
0707    A20 elements starting to sweep Hwy QL 19.
0713    A Trp requests dustoff for 5 VN civilians, litter patients from Arty shrapnel at XT498195.  Priority
0718    A30 elements completed sweep to CP 9.
0730    A20 elements at XT 482212 with dismounts out and beginning sweep.
0739    A30 elements swept to CP 16.
0747    A20 elements at XT480217.
0748    2nd Bde 25th ID reports Dustoff 163 ETA 7 to 10 mins.
0750    A Trp dustoff is complete.
0755    A30 elements returned to CP 8 and beginning sweep to FSB Stuart.
0758    A20 elements located at XT479219.
0800    SPOTREP A30 has road obstacle of heavy logs and brush on road at XT535182.  Cleared     road.
0829    A20 is at XT478221.  A10 is enroute to RIF area.  A30 has completed sweep to FSB     Stuart.
0838    A20 elements at XT477223.
0843    A Trp has mission to provide one plt to escort Arty convoy from FSB Stuart to FSB Pershing.  From
          MAJ Huff.
0850    SPOTREP A10 found four 8" arty rds at XT521192.  Ground was caved in and the hole had 4 rounds
          with fuses removed.  Will bring them back to FSB Stuart.
0900    A20 located at XT472227
0913     A20 at XT470232 & A10 XT520200
0920    SPTOREP A20 found 1 AK47 magazine & black PJs at XT520200 near base of tree in deserted
          village.  PJs were torn & ripped badly.  Request permission to recon by fire.
0950    A20 is at XT461248.
1039    A Trp reports Hwy QL 19 is swept.  A20 elements have returned to FSB.  A10 elements are at
1122    A10 elements have closed FSB Stuart.  A30 elements are escorting Arty convoy.
1135    A30 is at XT512261.
1145    A30 is at FSB Pershing.
1343    A30 elements are returning from FSB Pershing.
1402    A30 is at XT499243.
1410    A30 is at XT489211.
1423    A30 has reached FSB Stuart.
1540    A Trp's southern most platoon's strong point changed to XT525186 w/ambush patrol at
1625    A30's ambush patrol will be at XT4871197.  A10 & A20 elements at FSB Stuart.
1725    A Trp line-up from N to S is A20, A10, A30.
1729    A20 elements departing Main Gate.
1830    A20 elements in night laager.  A30 elements enroute to night laager.
1930    A Trp will sweep RT 6A for 2/12 IN Bn in the morning to XT499244 and provide a roving patrol until
          convoy has returned from FSB Persing. 30 RF/PFs will be available at 0730 hrs at XT490215 to
          assist.  Convoy will be on push 44.55.
1941    A30 elements have finished test firing.
2010    A30's ambush patrol is departing.  Radar scan: left 0000 mils, right 1940 mils at 1000m.
2030    SPOTREP A30 ambush patrol had 2 people at XT478187 spotted building a fire.  When people moved
          away fire died down.
2040     A20's ambush patrol is in position.
2130    SPOTREP A30 has 3 people at XT478184 in vicinity of fire reported earlier.  Engaged with unknown
2150    SPOTREP A20 has 10 lights moving west to east at XT520174 engaged with Arty & lights went out.
Dec 17
Summary: Troop establishes 3 NDPs. 1 platoon NDP located Trang Bang football field; 1 platoon located FSB
          STUART vicinity XT501194; 1 platoon strong point vicinity grid XT501194. Troop (-) sweeps Hwy 1
          from Cu Chi NW to 2nd Bde boundary. 2d Platoon moved out of FSB Stuart near Trang Bang to
          conduct morning sweep. 2nd platoon tank slews turret to the right and accidentally fires main gun.
          Young 8 year old girl walking at edge of road accidentally killed from concussion and canister round
          impacts the Buddist Temple. At 1150 hrs vicinity XT503260, 3rd platoon tank and truck detonated
          separated mines; results, 2 US WIA (dustoff), minor damage to vehicles. 3rd platoon Tank teeters
          on bridge that VC command detonated north of Trang Bang. Troop CO and 2d Platoon Leader, SFC
          Smith meeting when mortar rounds start impacting, 2d Platoon exits FSB for NDP. Trang Bang FSB
          received and returned small arms fire that night. 1st Platoon NDP receives and return small arms fire.
          3rd platoon AP observed 3 personnel and eng with SA; negative returned fire; results unknown.
0553    A Trp stand-to is complete.
0641    A10 elements will sweep RT 6A.  A20 elements will sweep from night laager to CP 16 & back to the
          FSB.  A30 elements will move from night laager to CP 20 & back to FSB.
0644    A30 elements at CP 19.
0650    A20 elements on the move.
0652    A30 is at CP 20 and turning around.
0655    A30 is at CP 19 & A20 is at CP 8.
0702    A30 is at CP 2 & A20 is at CP 17.
0706    A20 is at CP 9 & A30 is closing into FSB.
0715    A20 is at CP 16 & is turning around.
0723    A10 elements have started sweep at CP L.
0728    A20 is at CP 9.
0733    A20 is at CP 17.
0736    A20 is at CP 8.
0746    A20 elements have closed FSB Stuart.  Road is completely cleared.
0752    A10 is at XT490215.
0824    A10 is at XT491219.
0828    A20 elements departing FSB to sweep Hwy 19.
0855    A Trp reports that a main gun fired a cannister round by accident killing a VN girl. More to follow.
0900    A10 is at XT494225.
0905    A Trp report on accident: A27 tank was moving down MSR about 75m.  They had 1 cannister round
          in the chamber.  The TC went down to override w/main gun off to the right.  Rd went off and hit a
          little VN girl.  Medics went down to help but the girl died instantly.
0935    A20 at XT482212 & A10 XT495235
0948    A Trp sweep team on RT 6A has linked up w/sweep team from FSB Pershing. Outposting at this
1002    A20 is at XT477233.
1023    SPOTREP A20 reports that their A25 element has received small arms fire.
1045    A20 is at XT468235.
1102    A20 is at XT463244.
1118    A30 is at XT488206.
1123    A20 is at XT461249.
1129    A30 is at XT494230.
1138    A30 is at XT506260.
1141    A Trp reports that a vehicle escorting the convoy hit a mine.  SPOTREP to follow.
1145    SPOTREP A30 lead tank of escort hit a mine at XT508260 on the right side of the road roadwheel.
          Also 1 truck hit a mine off the side of the road.  Clearing tank to determine trafficability of bridge.
1150    SPOTREP A Trp reports 2 mines at XT508260 were hit one by a tank and one by a truck in the
          convoy.  2 US WIA, dustoff has been called.
1158    Dustoff 155 has an ETA of 20-25 mins.
1230    Warning Order to A Trp from MAJ Huff.  Be prepared to escort a VTR to CP 26 with 1 of your plts.
1245    Dustoff for A Trp is complete.
1333    SITREP A Trp: Cpt Moss feels that will need AVLB.  Also need a VTR to remove tank from bridge.
          Before the VRT or AVLB can come up the road the convoy will have to pull off to side of the road.
1417    A Trp reports the VTR has the tank off of the bridge and is preparing to tow it to FSB     Stuart.
1505    A Trp reports the bridge is passable to wheeled vehicles.
1516    A Trp requests a dustoff for 1 VN civilian.
1536    A Trp given mission by 2nd Bde 25th ID to supply 1 platoon to escort convoy to FSB     Wood.
1637    A30 elements returning to FSB Stuart at XT503242.
1706    A Trp has Arty convoy at Ambush Gate and is departing now.
1708    A Trp dustoff is complete.
1724    MAJ Rigby directs that the platoon escorting the convoy to FSB Wood will stay there     tonight.
1737    A Trp will escort the convoy to FSB Wood.
1809    Convoy w/A Trp escort is at CP 20.  A Trp going onto 3rd Bde 25th ID push.
1815    Convoy w/A Trp escort is at CP 5.
1830    A Trp reports internal security (base camp) is OK.
1831    Convoy w/A Trp escort is at CP 6.
1935    Convoy w/A Trp escort closed FSB Wood II.
2010    SPOTREP A30 reports a fire with movement at XT483197.  Spotted fire burning & could see 2 to 3
          individuals when fire flared up.  Reconning area by fire, unknown results.
2359    Base camp under rocket & mortar attack.
Dec 18
Summary: Troop establishes 3 NDPs. 1 platoon NDP located Trang Bang football field vicinity grid XT484196. 
          1 platoon located at FSB STUART vicinity grid XT501194.  1 platoon strong point vicinity grid
          XT550193. At first light Troop (-) swept Hwy 1 from Cu Chi NW to 2nd Bde boundary and Hwy 19
          from Trang Bang to 2nd Bde boundary.
0040    A Trp mission change: after road sweep they will wait for and pick-up convoy by CP 20 & proceed
          to Tay Ninh.  Go on 1st Bde 25th ID push between CP 22 & CP 33.  Pick-up Dau Tieng convoy
          between CP 25 & CP 36.  Switch to 3rd Bde 25th ID push and be under their control.  Release
          convoy in Dau Tieng, move back  & pick-up one platoon in woods, go to Go Dau Ha & then to FSB
0541    A Trp stand-to is complete.
0630    A30 elements sweep to CP 20, A10 elements sweep to CP 16.
0650    A30 is at CP 20, A10 is at CP 8.
0657    A30 is at CP 19, A10 is at CP 9.
0709    A30 returned to FSB Stuart, A10 is at CP 16.  Road is clear from CP 20 to CP 16.
0835    A30 is at CP 19.
0857    A20 elements have picked up the Dau Tieng convoy and are at CP 7.
0914    A10 at CP 20 with Dau Tieng convoy.  No commo with 3rd Bde 25th ID.
0925    A10 elements at CP 5.
0931    A30 reports lead of the convoy at CP 21.
0940    A10 is at CP 18.
1007    A10 is at CP 12.
1330    A10 lead elements at CP 18.
1337    A10 lead elements at CP 6.
1348    A10 trail elements at CP 12.  A Trp leaving 2/22 IN Bn push.
1350    A10 lead elements at CP 5.
1359    A Trp lead elements at CP 20.
1435    A30 elements at CP 4.
1437    A20 elements at CP 20.
1440    A30 elements at CP 21.
1445    A30 elements at CP 5.
1451    A20 is at CP 19 & A30 is at CP 20.
1500    A20 is at the FSB & A30 is at CP 19.
1501    A10 elements closed the FSB.
1508    A30 elements at the FSB.
1721    A20's ambush patrol is at XT478199.  A10's ambush patrol is at XT544176.
1817    A10 elements on the move.
1850    A10 elements in night laager.  Requests permission to test fire.  Granted.
1855    A Trp receiving RPGs, estimated 8 rounds, & sporatic small arms fire at FSB Stuart.  Returned fire.
1915    SPOTREP A Trp has 15 incoming rocket rounds at XT514195 from wood line Northeast of FSB Stuart.
          Returned fire with organic weapons.  Firing has ceased, observing area, & checking for casualties.
1916    Cdr A Trp believes attack was a coordinated effort.  FSB Stuart, FSB Pershing, & ARVN compound
          all hit at the same time.
2120    A Trp operation for tomorrow as follows: come up on 2/22 IN Bn push, set up strong points at
          following positions: XT460430 position C; XT464450 position B; & XT470470 position D.
2153    A Trp elements that escorts convoy contact 4/9 IN Bn as son as 1st vehicle is at CP 20.
2200    A20's ambush patrol is in position.
2210    Warning Order from LTC McGowan to A Trp: Be prepared to change with B Trp & move to Tay Ninh.
Dec 19
Summary:Troop (-) moved North to outpost Big and Little Rubber vicinity Tay Ninh. Troop (-) provides escort
          for convoy to Dau Tieng.
0541    A Trp stand-to is complete.
0605    A20 elements departing night laager going to FSB Pershing.
0625    A20 elements returned to FSB Stuart.
0642    A10 elements starting sweep.
0705    A Trp has swept from CP 7 to CP 8.
0707    A20 elements departing FSB Stuart.
0715    A20 is at CP 7.
0716    A30 is at CP 16.
0727    A20 is at CP 19 & A30 is at CP 17.
0730    A30 is at CP 8.
0740    A20 has completed sweep to CP 20.
0742    A20 is now on 4/9 IN Bn push.
0920    Lead of Dau Tieng convoy is at FSB Stuart.  A30 elements lining up to escort.
0922    A30 & convoy lead at CP 7.
0935    A Trp convoy lead at CP 17.
0938    A Trp convoy trail at CP 7 & lead at CP 20.
0945    A Trp convoy lead at CP 5 & trail at CP 19.
0956    A Trp convoy lead at CP 6.
0959    A Trp convoy trail at CP 20.
1023    A Trp convoy trail at CP 18.
1350    A Trp reports 2nd march unit is at CP 6.
1416    A Trp with convoy at CP 20, trail at CP 18.
1422    A Trp with march unit 2 is at CP 20.
1425    A Trp lead at CP 19.
1444    A10 elements closed to FSB Stuart.
1455    A30 elements closed FSB.
1525    A20 is at CP 20.
1532    A20 is at CP 19.
1541    A20 is at CP 7.
1556    A20 closed FSB Stuart.
1717    A Trp ambush patrols will be at XT525186 & 529184.
1748    A10 elements on the move.  A Trp line-up from North to South wil be A10, A20, & A30.
1818    A Trp just outside of night laager at this time.
1820    A10 elements in NDP.
1823    A Trp NDP is at XT525186 with ambush patrol at XT529184.
1829    A Trp internal security check of rear area at Cu Chi is negative.
1900    A30 elements in night laager.
1934    A30 ambush patrol in position.
1951    A10's ambush patrol location is XT479197.
2024    A10's ambush patrol is in position.
2036    A Trp radar scan: left 2500 mils, center 3470 mils, righ 1440 mils & range 1000m. Operational at this
Dec 20
Summary: Troop establishes 3 NDPs. 1 platoon NDP located Trang Bang football field; 1 platoon located at FSB
          STUART vicinity grid XT501194.  1 platoon strong point located vicinity grid XT550173. At first light
          Troop (-) sweeps Hwy 1 from Cu Chi NW to 2nd Bde boundary. 1 platoon North to outpost the Big
          and Little rubber vic Tay Ninh. Troop (-) provided escort for convoy to Dau Tieng.  Company B 1/5
          IN Inf(M) was conducting a RIF one kilometer north of the junction of Highways 1 and 7A, when at
          1355 hours, they made contact with an estimated reinforced enemy company at XT 560180.
          Company C, 1/5th(M) along with Companies B and D of the 2/12th Infantry and A/3/4 Cav were
          moved into the area of contact. Air strikes, LFTs and artillery were called in to support the forces.
          After the contact ended, a sweep of the battlefield was conducted. Enemy dead along with
          miscellaneous weapons and equipment were cleared from the battlefield.  Troop remains in Tay Ninh
          under OPCON of 1st Bde, 25th ID.  First night of operations in Tay Ninh, incident with VN SF. Troop
          requests permission to test fire, Bde Hqs approves and rescinds approval once Troop opens up.
0605    A Trp stand-to is complete.
0620    A10 on the move to the FSB.
0625    A10 is at CP 7.
0637    A30 is on the move.
0645    SPOTREP A30 has road obstacle at XT534182 consisting of debris, brush, sticks, 7 logs on road.
          Clearing at this time.
0705    A30 is at CP 16 and is turning around.
0710    A20 is starting to move, A30 is at CP 7.
0720    A20 is at CP 7 & A30 is at CP 17.
0726    A20 is at CP 19 & A30 elements are at FSB Stuart.
0740    A20 is at CP 20 & is leaving the A Trp net.
0804    A10 elements are in the Soccer Field.
0950    A Trp has march unit 1 (Dau Tieng) at FSB Stuart.
0953    A Trp is at CP 07.
1035    A Trp is at CP 18.
1431    A Trp is at CP 36 with a down tank.  Convoy has cleared.  Will hold about 30 mins to fix tank.
1843    A Trp internal security check of rear elements at Cu Chi is OK.
Dec 21
          No journal entries in 3/4 Cav Daily Staff Journals, attached to 1st Bde 25th ID at Tay Ninh.
Dec 22
          No journal entries in 3/4 Cav Daily Staff Journals, attached to 1st Bde 25th ID at Tay Ninh.
1836    A Trp internal security check of rear elements at Cu Chi. 4 EM & 1 APC for reaction force.
Dec 23
          No journal entries in 3/4 Cav Daily Staff Journals, attached to 1st Bde 25th ID at Tay Ninh
Dec 24
          No journal entries in 3/4 Cav Daily Staff Journals, attached to 1st Bde 25th ID at Tay Ninh.
Dec 25 1968, Christmas
          Troop on blocking positions around Tay Ninh. Christmas meal served at Tay Ninh base camp & in
          field. (View menu cover) (View menu) (Photo 1) (Photo 2) A10 replacement track held in Cu Chi to
          be used as a bunker next to stage for the entertainers from the Bob Hope Show. Doug Langdon,
          driver of A10 got to see the Bob Hope show. Major fireworks display in Tay Ninh area Christmas Eve.
          No journal entries in 3/4 Cav Daily Staff Journals, attached to Bde at Tay Ninh.
1919    A Trp rear element at Cu Chi has pyrotechnics locked up.
Dec 26
Summary:Troop still at Tay Ninh. Late December conducted night operations including escort of night resupply
          convoy for division. Platoons conduct NDP near Nui Ba Dinh. (photo)
          No journal entries in 3/4 Cav Daily Staff Journals, attached to 1st Bde 25th ID at Tay Ninh.
Dec 27
          No journal entries in 3/4 Cav Daily Staff Journals, attached to 1st Bde 25th ID at Tay Ninh.
1905    A Trp internal security check of rear elements at Cu Chi is negative.
Dec 28
          No journal entries in 3/4 Cav Daily Staff Journals, attached to 1st Bde 25th ID at Tay Ninh
Dec 29
          No journal entries in 3/4 Cav Daily Staff Journals, attached to 1st Bde 25th ID at Tay Ninh.
Dec 30
          No journal entries in 3/4 Cav Daily Staff Journals, attached to 1st Bde 25th ID at Tay Ninh.
Dec 31 1968, New Years Eve
          Troop in NDPs again major fireworks to celebrate the arrival of 1969
          No journal entries in 3/4 Cav Daily Staff Journals, attached to 1st Bde 25th ID at Tay Ninh.

Jan 1, 1969
Summary: Troop remains in Tay Ninh Base Camp in preparation to rejoin Squadron. Troop moves back down alt MSR,
           highway QL22. NDP at Trang Bang. (View 25th Inf Div Summary of Action for Jan 69)
1553     A Trp at CP18.
1624     A Trp at soccer field at Trang Bang.  Trail just south of CP18.
2132     A Trp has negative SITREP.
Jan 2
0032     A Trp has negative SITREP.
0525     A Trp stand-to is complete.
0719     A Trp elements have swept to CP20 and are returning.
0750     From MAJ Huff, S3, B Trp sub-element to provide security for the advance party until A Trp arrives.
1240     A Trp lead is at XT494228.
1243     A Trp lead is at XT508260.
1340     A Trp is in NDP.  B Trp released from outposting.
1645     A Trp NDP is at XT529283 (southern Boi Loi).
1800     A30's plt NDP is at XT527277.
1820     A30 element is in his NDP.
1920     A30 rpts civilians leaving villages vic XT524275.  Requests test fire of .50 cal.  Granted on 320
           degree azimuth by LT Kaiser.
2033     A Trp SITREP: no change.
Jan 3
Summanry: Troop moves North up the road to Boi Loi Woods. One tank hits mine and is evacuated.
0025     A Trp SITREP: negative.
0602     Stand-to complete in A Trp.
0615     A Trp ambush patrol has returned to the perimeter.
0700     A Trp on the move and sweeping.
0720     A10 finished sweep, negative findings.
0835     SPOTREP: LA A30; LB 1 M79 rd; LC XT516264, time 0830hrs; LD lying in road, engr can't determine
           if booby trapped; LE using grappling hooks to pick up.
1041     SPOTREP: LA A20; LB tank hit mine; LC XT541301 at 1010hrs; LD blew road wheel off and 1 track;
           LE VTR coming to get tank & fill hole in road.
1226     A Trp blowing some tunnels & moving through.
1240     A Trp is at XT530325.
1253     Squadron HQ starting to setup CP.
1905     A10 element is at XT550312.
1958     A20 listening post is in position.
2328     A Trp has a negative SITREP.
Jan 4
Summary:  Squadron S3, Maj Huff, and Chaplain  ___________with the Troop. 2 Plts establish laager to secure Squadron HQ. Trp Cmd Section and 3d Platoon conduct reconnaissance west of NDP. 1st PLT conducts independent sweep. Operations conclude late afternoon and rejoin 2d Platoons in NDP. 1st platoon mortar fires 2 illumination rounds as registration to support C Troop NDP. Division CounterMortar Radar picks up rounds and authorizes counterbattery fire mission. ARVN Arty Battery fires the mission. Rounds impact on and in front of NDP.  No injuries to personnel or damage to equipment.
0321     A Trp has negative SITREP.
0600     A10 moving out of NDP on their sweep.
0620     A10 is located at XT542303.
0630     A30 moving out of NDP on their sweep.
0800     A10 has closed on the troop NDP.
0920     A30 is at XT536288.  A Trp is at XT535287.
1040     A30 location is XT544306.
1113     A30 location is XT516265.
1225     A30 location XT494233.
1230     A30 location XT490306.
1304     SPOTREP:  LA A10; LB tunnel complex; LC XT555305 1240hrs; LD 3'x4' freshly built containing
           2 ponchos, 3 steel helmets, penicillin bottles, 40 empty sardine cans, 2 bodies (7 or 8 months old),
           & 1 sandal; LE continue to search area.
1355     A30 is located at XT489205.  A10 located at XT559319.
1400     A30's location is XT544233.
1431     A30 is at CP GIN.  A10 has closed on CP's location.
2058     A10 listening post departing position.
2103     A10 ambush patrol in position.
2123     SPOTREP: LA A20 Listening Post; LB voices & movement; LC XT564323 2120hrs; LD voices &
           movement to front of LP, size not known; LE continuing mission.
2135     A Trp is relocating LP and will recon by fire.
2152     A Trp LP relocated at 355 degrees 100 meters from perimeter.
Jan 5
0124     A Trp SITREP: A Trp had a claymore with a bad detonator and blew it accidentally.
0455     A20 & A30 listening posts have returned to NDP.
0650     A20 is escorting the squadron resupply convoy.
0857     A Trp sweep location is at XT542301.
0908     A Trp located at XT540295.
0925     A Trp sweep at XT535288.
1005     SPOTREP:  LA A20; LB booby trap; LC XT531283; LD 1 US grenade in a can with the pin pulled with
            wire stretched across the road; LE blowing in place.
1013     A Trp rpts convoy made it past the first mud hole.
1022     A Trp rpts resupply trains at XT541300.
1027     SPOTREP: LA A20; LB foot prints; LC XT530283 at 1025hrs; LD sandal prints of about 15 to 20
           people moving east down trail.
1100     A20 sweep at XT519266.
1120     A20 sweep at XT509259.
1130     A20 sweep at XT502250.
1150     A20 at XT495230.
1153     A20 at XT488205.
1156     A20 at FSB Stuart.  A30 is on the move.
1240     A30 is at XT547303.
1250     A Trp will reconn on 100 degree azimuth.
1301     A20 is on the way from FSB stuart with resupply to squadron's location.
1305     A20 is at XT488205.
1315     A20 is at XT495230.
1320     A Trp lowboy crossed into FSB Pershing.
1329     A30 is at XT562303.
1330     A Trp reports that they left one APC in FSB Stuart.  Parts will be sent out to it.
1333     A20 is at XT510260.
1340     A Trp reports that lowboy w/dozer on the move to this location.
1344     A20 is at XT529283.
1358     A10 is at XT540295.
1415     A20 is at XT555313.
1435     A30 has closed on squadron CP.
1446     A20 is at XT555313.
1515     A Trp will escort lowboy back to FSB Stuart.
1641     A20 coming from FSB stuart has picked up a VTR, lowboy, and APC on the way back to squadron
1725     A20 dropped off 2 elements, has B Trp vehicle & returning to squadron NDP.
1748     A Trp reports all elements have closed on NDP.
1750     A10 listening post 240 degrees, 100 meters.
1800     A20 listening post 45 degrees, 100 meters.  A30 listening post 100 degrees 315 meters.
1935     A Trp test fire complete.
1945     A30 listening departing perimeter.
2009     A10 & A30 listening posts in position.
2020     A20 listening post in position.
2031     A Trp has negative SITREP.
Jan 6
Summary: A Troop lost a APC when the vehicle detonated an anti-tank mine while moving along Route 6A ten kilometers northeast of TRANG BANG (XT530283).  One cavalryman riding in the vehicle was wounded the blast.
0423     A Trp has negative SITREP.
0625     A10 elements are on the move.
0710     A10 is at XT541301.
0713     VTR has arrived at the mud hole to pick up the A Trp tank that blew a jug.
0729     A10 is at XT541301.
0759     A10 is at XT533286.
0805     A20 is on the move.
0828     SPOTREP: LA A20; LB holes; LC XT557337; LD 2.5'x3.5'x3' in built up area probably used as bunkers
           by villagers.
0840     SPOTREP:  LA A20;  LB grenade; LC XT554340; LD US grenade looks like it has been there a while,
           has no safety; LE blown in place.
0843     A10 is 100 meters from FSB Pershing.
0915     SPOTREP: LA A20; LB well; LC XT557337 at 0908hrs; LD well was blown with C4; LE cont mission.
0930     A20 is at XT545338.
1002     A10 is at XT598241.
1020     SPOTREP:  LA A20; LB 1 MK82 bomb; LC XT544343 1015hrs; LD  dud on ground recent impact; LE
           do not have enough C4 to blow, leaving in place.
1035     A10 is at XT495232.
1150     SPOTREP: LA A20; LB foxholes & tunnel; LC XT553341 1050hrs; LD foxhole connects to tunnel 15'
           long, not recent; LE blown in place, cont mission.
           SPOTREP: LA A Trp; LB two deep holes; LC XT554340 1100hrs; LD 3' by 30' deep not recent; LE
           blown in place, cont mission.
           SPOTREP: LA A Trp; LB tunnel; LC XT557338 1130hrs; LD 3' diameter, 5' deep, connects to tunnel
           15' long branches to right and left for 5'. Tunnel has holes for candles, found 1 VC canteen, 1 NVA
           belt, 1 toothbrush, toothpaste with NVA writing & 1 AK round; LE blown in place, bringing in found
1230     A Trp requests dustoff: 3 litter, shrapnel in face, 1 shrapnel in leg; XT430283; APC hit mine; ETA
           15 minutes, callsign Dustoff 155.
1301     A Trp dustoff complete.
1312     SPOTREP: LA A Trp; LB mine; LC XT530283 1215hrs; LD mine tore track, road wheel, instrument
           panel, & hole in drivers' compartment of APC.  3 litter w/frag wounds; vehicle combat loss; LE
           towing vehicle to FSB Stuart & dusted off WIAs.
1845     A Trp listening posts: A10 240 degrees at 100 meters; A20 345 degrees at 100 meters; A30 150
           degrees at 100 meters.
1914     A20 LP departing perimeter.
1920     SPOTREP: LA A28; LB 2 camp fires; LC XT581322; LD observed 2 fires burning, one out, one still
           burning; LE request permission to fire/request denied.
1936     A10 & A20 LPs in position.
1939     SPOTREP: LA A25; LB 4 lights; LC XT562364; LD spotted moving west to east: LE cont to observe.
2016     SPOTREP: LA A31; LB lights; LC XT563332 2000hrs; LD moving west to east: LE cont to observe.
2025     A Trp radar; left 3872 mills, right 5740 mills range 5000 meters.  A30 LP in position.
2040     SPOTREP: LA A25; LB 11 lights; LC XT559332 2035hrs; LD moving to east: LE cont to observe.
2226     A Trp negative SITREP.
Jan 7
Summary: Troop is a reaction force for a LRRP Team 11 from F Troop, 75th Ranger. LRRPs had been inserted by helicopter into the Trapazoid.  LRRPs under attack at 1345 hrs. Troop moved to extract LRRPs and received small arms fire & RPG fire at 1445 hrs 600 meters to North of LRRP teams location at XT578294. A Troop engagement with enemy lasts until 1540 hrs. Enemy disengages at 1540 hrs and LRRP Team was picked up by A Troop at grid XT578296. (photo LRRPs in NDP) Troop moves to night postition at grid XT565318.
0110     A20 spotted movement.  Recon-by-fire with all organic weapons, negative results, & repositioned
0651     A30 moving out on sweep.
0749     A30 is at XT549318.
0812     A30 at XT542304.
0832     A30 at XT541301.
0833     From G3 to MAJ Harrison: after todays' opn we can operate in the AO as we see fit.  Expect to
           stay a minimum of 3 days.
0840     A30 at XT530296.
0905     A10 on the move to pickup downed APC.
0915     SPOTREP: LA A30; LB 105mm rd; LC XT536287 912hrs; LD lying on side of road; LE Engrs preparing
           to blow in place.
0916     Down tank and APC to be dropped off in FSB Stuart if it can't be fixed here.
0937     A30 at XT530284.
0942     A Trp Cdr has people moving around in front of his elements and is trying to get a detainee.
0953     A Trp released detainees (farmers tending cattle).
1113     A Trp has picked up downed tank and are working on the APC.
1139     A30 has met up w sweep from FSB Stuart.  Stopping to eat at this time.
1155     A Trp has the downed APC, heading back to the CP.
1310     A10 at XT541302.
1326     A Trp elements have closed the CP. (no journal pages for remainder of day)
Jan. 8
Summary: One Plt secures squadron CP and one Plt escorts squadron trains.  1st Plt A Trp inserts LRRP TM while on RIF. (photo) (view LRRP AAR)Plt then moved North in the Mushroom to conduct reconnaissance-in-force along the banks of the Saigon River (see Map). Plt encountered bunker complex and dug-in NVA with backs to river at XT575315. Received small arms, recoilless rifle and RPG fire. Plt assaulted the position three times. A16 tank takes RPG in turret. Withdraw to evacuate WIA. SSG Manny Montanez and PFC Linkins medevaced. SP4 Steve Turner treated in field. Div Cdr lands in position and ordered the 3rd assault.
0125     A Trp has negative SITREP.
0615     A Trp stand-to is complete.
0733     A20 reconnng w/M79 around a hole found in his sweep area.
0807     A10 departing for the LRRP TM 11 insertion.
0835     A10 is located at XT572303.
0848     SPOTREP:  LA A10; LB tunnel; LC XT574299 at 08:40hrs; 8' long, 6' deep' 4' wide with no sign of
           recent use, also found bamboo basket shovel, bunker 50 meters away 3'x2' w/plastic top, has
           recent tunnel leading off of it that contained 2 raincoats; LE complex will be blown.
1010     A Trp security element for resupply is leaving forward CP.
1017     SPOTREP:  LA Engr with A20; LB booby trapped grenade; LC XT52068 at 1000hrs; LD one M26
           grenade with pin pulled inserted in C-ration can; LE blown in place.
1115     A10 received one RPG round, pulling back to call in Arty.
1128     A Trp has gunships making run at XT575315.
1213     Resupply departing Trang Bang.  One tank staying to escort water trailers.
1221     A10 recieved another RPG round, 1 tank hit, FAC will bring in air strike.
1250     A Trp has a dust-off, one man with shrapnel wounds at XT569305.
1325     LRRP TM 11 inserted at XT582305.
1339     Air strikes for A Trp complete.
1340     FAC reports that there were 3 500lb duds during the strike.
1349     Air strike uncovered a bunker complex and had one secondary explosion.
1405     A Trp dust-off, 1 urgent litter at XT 569305 with head wound.
1425     A Trp dust-off complete.
1449     A Trp air strike complete with another one coming in.
1459     FAC 04 being replaced on station with FAC 21, air strike complete.
1502     A Trp is at XT569305.
1524     A Trp will get another air strike and then a CS drop.  D Trp will stand-by with ships.
1536     A Trp notified CS drop will not go in.
1555     A Trp's latest air strike has uncovered more bunkers and spider holes.
1558     FAC reports more fighters coming up in 15 mins.
1605     A10 moving back into area where they received the RPG fire the last 2 days.
1611     A10 tank has received another RPG wounding the TC and one other person. (SSG Montanez & SP4
1630     A10 elements moving back to previous area, dust-off is complete.
1637     A Trp tank was hit in turret from a different location, Arty is being put into area.
1726     A Trp air strike is complete.
1800     A10 closing NDP, A30 on move to NDP.
1809     A Trp NDP is at XT569305.
1900     FAC report on air strikes that there is possible 3 ammo fires still burning, 1 fire getting bigger.
1940     A Trp LP's: A30 at 155 degrees 70 meters; A20 40 degrees at 100 meters; A10 300 degrees at
           100 meters.  Radar: left 4200 mills, center 5177 mills, right 6255 mills at 5000 meters.
1953     A20 & A30 LPs departing NDP.
2007     A10 LP departing NDP.
2008     A30 LP is in position.
2025     SPOTREP:  LA A20 LP; LB movement; LC XT568308 at 2015hrs; LD 1 to 2 personnel moving towards
           perimeter 80 meters out, observed with starlight scope; LE engaged with organic weapons, results
           not known.
2030     SPOTREP: LA A Trp radar; LB movement; LC 5190 mills time now; LD movement of unknown size; LE
           calling Arty.
2155     SPOTREP: LA A17; LB moving lights; LC XT570297 at 2150hrs; LD lights moving southwest on 360
           degree azimuth, very far out; LE observing.
2217     SPOTREP: LA A27; LB lights; LC XT572307 at 2215hrs; LD lights moving left to right on a 50 degree
           azimuth about 350 meters out; LE continuing to observe.
Jan 9
Summary: A Trp(-) conducts sweep operations in lower Boi Loi west of the Mushroom. At 09:35 hrs troop discovers 100 lbs of rice 1 km south of the Mushroom at XT567307. MG Williamson lands in 1st Platoon area before final assault on fortified NVA position that was assaulted on the previous day. NVA position breeched and bunkers blown. Troop continued Reconnaissance-in-Force operations.
0025     A Trp has a negative SITREP.
0600     A Trp stand-to is complete.
0715     A Trp trains departing NDP for CP location.
0844     MG Williamson overhead in the squadron's area of operation.
0850     3rd Bde 25th ID gave permission for squadron to enter AO BUD.
0911     Arty prep for A Trp completed at this time.
0935     SPOTREP: LA A Trp; LB bunkers reinforced with sandbags/received one rd automatic weapons fire;
           LC XT567307; LD blowing bunkers in place; LE continuing mission.
1015     Ref: SPOTREP, also 2 bags of rice, 1 canteen, 1 gas mask, 3 bunker complexes with tunnels.
1035     A Trp found some personal letters, 2 electrical diagrams, & 1 RPG 7 rd.
1042     SPOTREP: LA A Trp; LB documents; LC XT567307 at 1040 hrs; LD pamphlet bag with 2 maps, 1
           map of Cholon, medical supplies, and assorted pamphlets; LE evacuating items to rear.
1238     SPOTREP: LA A Trp; LB bunkers; LC XT572314 at 1230 hrs; LD 6'x6'x6' with cooking room on side
           4'x5'x6'; LE destroyed in place.
1249     A Trp will move down to XT578294.
1252     A Trp at XT573307 sweeping southeast to limit of advance.
1303     SPOTREP: LA A Trp; LB bunker & tunnel; LC XT575307 at 1300 hrs; LD 1 sandal inside; LE used
           tanks to crush complex.
1333     A Trp has found 23 sandbags, 1 can of rice, & letters.
1355     A Trp still working on on track.  A10 & A20 preparing to move out.
1400     Air strike is complete.
1434     A Trp at temporary halt at tunnel containing 1 pair of pants.
1522     A Trp reports that the bunkers caved in yesterday smell like dead bodies.
1529     SPOTREP: LA A20; LB 3 tunnels; LC XT580304 at 1500 hrs; LD 3 bunkers with the tunnels, one a
           sleeping position, lots of flowers; LE continuing mission.
1554     A 10 & A20 closed squadron CP.  A30 coming up shortly.
1613     A30 closed CP.
1640     Air strikes in XT4931 now complete.
1940     A10 LP on 25 degree at 190m, A20 LP on 310 degrees at 100m, A30 LP on 220 degrees at 70m.
1948     A10 LP departed CP.
1952     A30 LP departed CP.
2004     A30 LP in position.
2033     A30 LP spotted movement 100m in front of their location.
2037     A Trp radar scanning L 360 degrees, right 100 degrees, range 1000m.
2101     A Trp spotted movement at XT561360 at 100m.
2105     A Trp sent LP out to position.
2140     A10 requested to test fire on 270 degrees azimuth.
2145     SPOTREP: LA A20; LB moving lights; LC XT553340 at 2140 hrs; LD moving about 300 meters to
           front of position; LE fired on light continuing to observe.
2151     A Trp reports no fire returned
2201     A10 completed test fire.
2320     A Trp requests dust-off for 1 litter patient who has severe stomach pains.
2328     Dust-off 48, ETAin 10 mins.
2345     Dust-off is complete.
2348     Results from G2 on documents found tentatively identify units as the recon elements of the 273rd
           VC/NVA regiment of the 101st NVA Regiment.
Jan 10
Summary: Trp operations in lower Boi Loi continue to disrupt enemy supply lines. At 09:45 hrs Trp discovers one napalm bomb, 4 bicycles, and 4,400 lbs of rice at XT498309. 1505 hrs A Trp discovers 12,600 lbs of rice.
0020     A Trp has a negative SITREP.
0602     Stand-to is complete.
0711     A Trp is on the move.
0723     A Trp trains at CP Rye.
0803     A Trp on the road in herring bone.
0850     A Trp has married up with C Trp and both have begun the sweep.
0909     A Trp & C Trp located at XT508309.
0910     LTC McGowan has an air strike on XT494314 at 0930 hrs.
1020     A Trp has an APC with transmission problems.
1023     A Trp found 25 100lbs bags of rice at XT498309.
1035     A Trp has 1000lbs of rice.
1046     A Trp has 900lbs of rice on poles above the ground.
1052     A Trp found a tunnel 4'x4'x8' with a hoe, shovel, candles, & fresh work.
1055     A Trp receiving some AK fire from the rear.
1130     A Trp at XT496303 found some bunkers, at least 4 bunkers in the area.
1235     A Trp at XT506317 has tank with thrown track, trying to walk it back on.
1243     Daisy Cutter will be dropped on XT514324.
1248     SPOTREP: LA A10; LB bunker complex; LC XT492316 at 1200 hrs. LD tunnel 8' deep, 6' wide, 6' long
           & bunker 4'x4' with 1 AK rd, 1 NVA gas mask, 1 NVA canteen, 1 pair of PJs; LE destroyed.
1258     Daisy Cutter mission completed.
1505     A Trp found 1800lbs of rice, not good and destroying in place.
1550     SPOTREP: LA A Trp; LB rice; LC XT496316 at 1500 hrs; LD 17,200lbs of rice in bad condition;
           LE burned in place.
1550     SPOTREP: LA A30/A35; LB foxholes; LC XT496316 at 1515 hrs; LD 4 foxholes 3'x4'x2' used for
           sleeping area, 4 bicycles; LE blew tunnel & continuing mission.
1550     SPOTREP: LA A20/A25; LB tunnel opening; LC XT496316 at 1515 hrs; LD tunnel 3'x3x4' sleeping
           area & 4 bicycles; LE blew tunnel & continuing mission.
1617     A Trp NDP location XT530309.
1652     A Trp trains departing squadron CP.
1825     A Trp captured 6 bikes and 18,000lbs of rice.  Can't tell exactly how much they are bringing in.
1940     A Trp LPs: A30 310 degrees at 50m, A10 25 degrees at 100m.
1959     A30 LP departing perimeter.
2003     A10 LP in position.
2008     A30 LP in position.
2013     A20 LP in position.
2014     A Trp Radar: left 2146 mills, center 3923 mills, range 5000 meters.
2328     A Trp has a negative SITREP.
Jan 11
Summary:  Trp continues operations in lower Boi Loi 7 to 10 km southeast of Bao Don at XT4932 to XT5329.
0028     A Trp has a negative SITREP.
0747     A Trp trains arrived at the squadron CP.
0806     A Trp moving out on RIF operation.
0811     A Trp trains closed squadron CP.
0851     A Trp location is XT516322.
1000     A Trp found a bunker and trench line.  Has 2 APCs stuck from falling into bunkers.
1050     FAC O4 spotted a large bunker on A Trp's right flank.
1101     A Trp reported it has already caved in.
1115     A Trp location XT522323.
1227     A Trp location XT537328.
1255     A Trp on the move again.
1325     A Trp location XT525329.
1335     SPOTREP: LA A Trp; LB rice; LC XT533329 at 1330 hrs; LD 70 100lbs bags; LE bringing in.
1408     A Trp location XT539329.
1530     A Trp has closed its NDP.
1618     A Trp trains have departed squadron CP.
1950     A Trp LP status: A10 on 99 degrees at 75m, A20 at 200 degrees at 75 m, & A30 at 318 degrees at
1959     A Trp LPs have departed perimeter.  Radar status: left 3296 mills, center 4273 mills, right 5250
           mills, & range 5000m.
2005     A20 LP in position.
2015     A10 & A30 LPs in position.
Jan 12
0033     A Trp has negative SITREP.
0311     A Trp LP returned.
0553     A Trp stand-to complete.
1055     A Trp lead vehicle hit a mine at XT523285, no casualties, blew off right front road wheel.
1109     SPOTREP: LA A27; LB hit mine; LC XT525275 at 1105 hrs; LD 50 lbs mine, blew off drivers hatch,
           detonated by claymore blaster, road wheel arm broken; LE towing to FSB Stewart.
1140     SPOTREP: LA A Trp engineer team; LB mine; LC XT522271 at 1130 hrs; LD mine on right side of
           road; LE blown in place, left a 2'x2' crater.
1151     A Trp at FSB Pershing, will sweep and move out.
1200     A Trp lead at XT568258.
1215     A Trp lead at XT495230.
1250     A Trp lead at CP18.
1320     A Trp lead at CP12.
1330     A Trp closed FSB Wood at XT4637.
1726     A Trp NDP at XT456364.
1825     A Trp NDP XT459338.
1853     A Trp LP status: A10 220 degrees at 100m, A20 95 degrees at 75m, & A30 310 degrees at 100m.
1948     A10 & A20 LP in position.
2014     A30 LP in position.
2356     A Trp has a negative SITREP.
Jan 13
0023     A Trp has negative SITREP.
0131     SPOTREP: LA A Trp; LB movement; LC XT453372 at 0105 hrs; LD light that appeared to be a
           cigarette; LE pulled in LP and reconned area by fire,  will send LP back out.
0519     A Trp stand-to complete.
0740     LTC McGowan having troops hold in place until C&C ship arrives.
0800     A Trp to move from XT475375 on 155 degree azimuth.
0815     A Trp on move for squadron RIF.
0945     Cpt Moss rpts his trp net has 3 other units on it.  Requests change of net.
0952     A Trp to use HHT net for operation.
0957     A Trp at XT488359.
1005     A Trp at XT478360.
1105     A Trp at XT478355.
1244     A Trp at XT479354.
1446     A Trp at XT478352.
1513     A Trp at XT480370.
1600     A Trp closed NDP.
1705     A Trp NDP at XT460371.
1830     A Trp LP status: A10 on 150 degrees at 75m, A20 on 260 degrees at 75m, & A30 on 340 degrees
           at 100m.
1924     A Trp requested permission to test fire one 50 cal on 150 degrees.  MSG Saylor gave permission.
1945     A20 LP in position.
2006     A Trp radar status: left 4120 mills, center 5097 mills, right 6074 mills, & range 3000m.
2015     A10 & A30 LPs in position.
2330     A Trp has a negative SITREP.
Jan 14
Summary - Troop directed to NDP to link up with Colonel McGowan. (Photo 1) Laagered a little over 3 KM from drop area. B-52 Arc Light of area north of Troop. Surreal viewing B-52 strike.  Troop prepared for first B-52 strike and totally unprepared for follow-on drop. (Photo 1) (Photo 2) First light Troop conducts BDA.
Laagered that night around old bomb craters filled with water. Troops take turns getting clean. (photo)
0940     A Trp located at XT482379.
1011     A Trp located at XT484375
1030     SPOTREP: LA A20; LB bunkers; LC XT484378 at 1020hrs; LD 3 bunkers 10'x10', 4 fighting positions,
           bunkers had heavy overhead cover; LE blown in place.
1351     A Trp has made a 180 turn and is headed back at this time.
1445     A Trp will be returning to squadron CP to refuel.
1603     A Trp trains departing squadron CP.
1820     A Trp LP status: A10 270 degrees at 100m, A20 300 degrees at 75m, & A30 100 degrees at 100m.
1920     A Trp mission for tomorrow: SP at XT484274 on an azimuth of 156 degrees.
2013     A Trp LP's in position.
2325     A Trp has a negative SITREP
Jan 15
0023     A Trp has a negative SITREP.
0344     A Trp has a dust-off, 1 litter with possible kidney trouble at XT488382, area is secure.
0357     Dust-off 46 ETA is 10 to 15 min.
0412     Dust-off is complete.
0509     Stand-to is complete.
0748     A Trp beginning RIF.  Start point is XT484374.
0915     A Trp dust-off to Dau Tieng, routine, kidney problem.
0932     Dust-off 46 out of Dau Tieng at 1200 hrs.
0951     A Trp is at phase line PEACH.
0958     Dust-off cancelled.
1217     A Trp on the move.
1234     A Trp has a APC with a thrown track.
1321     A Trp is on the move.
1550     A Trp is out of wooded area into plow area at XT471377.
1620     A Trp trains departed FSB Woods enroute to trp NDP.
1635     A Trp NDP at XT458253.
1755     A Trp LP status: A10 320 degrees at 75m, A20 90 degrees at 100m, & A30 10 degrees at 70m.
1932     A20 LP departing NDP.
1938     A20 LP in position.  A10 LP departing NDP.
2007     A20 & A30 LPs in position.
2325     A Trp has a negative SITREP.
Jan 16
Summary: 09:00 hrs troop uncovers 2 RPG-7 rds, 4 RPG-2 rds, 10 lbs of explosives, 100 rds small arms ammo, & 2 lbs of documents at XT510364. At 13:45 hrs engaged unknown number of enemy forces.  Two enemy killed and one trooper wounded during action. Troop maneuvered to the rear of Boi Loi to laager with C Troop. B Troop laagered 1 km away. Tank throws track and crew works to get it back on.  Squadron at a halt Tank track fixed, A Troop continued maneuver and had fire fight at 1820 hrs with NVA forces. At 1850 hrs a Cobra gunship supporting A Troop diverted to LRRP Team 15 in vicinity at grid XT529357.  LRRPs view enemy forces fleeing Troop. (view LRRP AAR)
0029     A Trp has a negative SITREP.
0540     A Trp stand-to is complete.
0844     Squadron has 3 daisy cutter airstrikes scheduled at 0900, 0930, & 1000 hrs.
0900     A Trp found 4 RPG rds (4 B40 & 2 B41), a canteen, a bag of explosives, & 100 rds of AK ammo
           lying on the ground.
0905     A Trp located at XT486381 holding for airstrike.
0930     2nd air strike going in.
0945     A Trp found a diary & a supply book at XT486380.
0955     2nd air strike uncovered numerous bunkers.
1000     A Trp is at XT490365.
1025     A Trp is holding in place to put track back on 2 APCs.
1041     Air strike 3 is complete.
1101     A Trp is at XT493373.
1158     A Trp requested dust-off for 1 litter case w/dislocated shoulder.
1300     A Trp requested dust-off for 1 man hit by a tree branch.
1416     A Trp at XT502368.
1437     A Trp has a tank with a thrown track.
1635     LRRP TM 15 spotted 10 VC approaching C Trp NDP.
1731     A Trp & C Trp at XT528354.
1743     LRRP TM 15 spotted 20 VC about 50 meters from his position on 340 degree azimuth.  Requests
1810     A Trp on the move again.
1950     A Trp in NDP
1905     LRRP TM 15 has movement all around his position.  Gunships scrambled by Maj Huff.
1910     Light Fire Team and flare ship enroute to TM 15's position.
1925     TM 15 requests LFT be cancelled.
2320     A Trp has a negative SITREP.
Jan 17
Summary: Squadon in major battle 8.5 km east of Bao Don. (see map) First light A & C Troop in combined laager receive RPG and SA fire. Both Troops return fire. 0730 hrs A Troop directed to reconnaissance-in-force in area of enemy firing.  C Troop remain in position awaiting LRRPs and then to maneuver abreast of A Troop.  B Troop maneuvering to vicinity. 09:05 hrs NVA force hits Troop immediately on forward movement XT527349.  Body count began, NVA in web gear and helmets. A Troop encountered fortified bunkers and NVA body count continued.  - C Troop directed to move abreast of A Troop on right flank. 1045 hrs LRRP Team 15 links up with C Troop at XT528353. NVA cause A & C Troop to cross fire on each other. Colonel McGowan directed cease fire and conduct180 to return to NDP. A Troop does a 180 degree turn, 1st and 2d Platoons respond and began maneuver.  3d Platoon had tank that could not gain traction and unable to maneuver to NDP.  Troop Hq’s and Medic track remained with 3d Platoon.  Medic track takes RPG to ACAV kit, smokes detonate and moments pass before track commander reported all OK.  3d Platoon tank gets traction and links up at original NDP. A, B & C Troop deployed in same vicinity.  C Troop Acting Cdr directed F4s in close Air support, shrapnel explodes and C Troop Acting Cdr medivaced.  (photo 1, photo 2)
Colonel McGowan directs new mission.  A Troop lead, B Troop followed by C Troop In column down edge of Bo Loi Woods.  Mission was not to stop, roll over fortifications. A Troop moved out with 1st Platoon in lead, followed by 2d and 3d plts. Dirt trail around flank of enemy was lined with NVA bunkers. Tanks fire at point blank range into bunkers. Phantoms provided napalm on right flank and artillery walked in front of A Troop.
Movement continued, NVA body count builds until 3d Platoon tank bellies up and unable to move forward.  NVA target 35 tank with RPG. Troop medics, Sp4 Jones and Sp4 O’keefe, run from track to track under heavy fire to treat wounded.  Sp4 O’Keefe arrives at 35 tank and administered life saving assistance to Sp4 Neal.  Troop Cdr, members of 3d Platoon and medics  on the ground vicinity of 35 tank now burning.  Casualties evaced to other tracks and troopers return to their tracks. A Troop commenced movement, A6 track received RPG at point blank range and skins under belly of track.  A6 Driver reported NVA in spider hole to the front.  A6 track proceeds forward and Troop Cdr drops grenade down spider hole.  NVA throws  grenade back and shrapnel strikes Troop Cdr in back. Troop Cdr reported to Colonel McGowan, one tank hit by RPGs, blowing up and he was hit.  Continued movement out of kill zone. Colonel McGowan directed after lengthy pause, “Roger that my friend, continue movement and I will have you out of there.” 3d plt tank was abandoned and on fire, B Troop sweeps around burning tank. 1st plt continued advance on squadron axis.  3d Platoon still halted and having trouble maneuvering forward.  Lt Ayres mortally wounded, Platoon Sergeant Brewington wounded and platoon rallied by ___________ to move forward with 2d Platoon following. Colonel McGowan directs left turn into clearing to form laager and evac casualties. 1st Platoon executed left turn and began establishing laager. (photo 1) 3d Platoon and 2d Platoon delayed in arriving at laager.  Finally, after much individual efforts, 3d Platoon arrived in laager followed by the 2d Platoon. Col.'s LOH started to take out casualties  Medivacs arrive and casualties  dusted off.  Lt Ayres and Sp4 Neal among the first out. A Troop regrouped in laager.  B Troop pulled past and now in lead.  C Troop having problems maneuvering past destroyed tank that was exploding. A Troop ready to move out and B Troop B6 TC reported Troop Cdr needed medivac, wounded by grenade.  Colonel McGowan immediately leaves and goes to B Troop to get the Troop Cdr out. A Troop Cdr requested permission to maneuver around B Troop to break out of the jungle. A Troop on move again. Sweeps around B Troop. By 1530 hrs the squadron had broken through the NVA position. A Troop breaks out of Boi Loi and establishes NDP at about 1700 hrs.
Colonel McGowan  came in to medivac Cpt Moss. All three Troop Cdrs medivac that day. Lt. Moore acting commander for night and next day. PSG Smith had command of both 2d and 3d plts. Sweep of area of contact by squadron results in finding 46 enemy dead, 2 AK47 rifles, 1 RPG-7 launcher, 4 RPG-7 rds, 1 machinegun, 1000 rds of small arms ammo, & numerous documents.  Squadron has 1 KIA, SP4 Neal, (photo/shadowbox) and Lt Ayers (photo/shadowbox) & 16 others wounded. (view article vol 4 no. 8 of the Tropic Lightning News)
0037     A Trp has a negative SITREP.
0531     A Trp stand-to is complete.
0602     A Trp and C Trp receiving small arms fire from the south of the NDP.
0612     Arty going in on suspected enemy location.
0910     SPOTREP: LA A Trp; LB 2 VC bodies; LC XT527349 at 0858 hrs; LD spotted VC jumping into tunnel;
           LE dropped grenades in the hole, found 2 bodies, 1 AK47, & 1 RPG launcher.  Continuing mission.
0955     Gun team scrambled in support of A & C Troops.
1012     Late entry at 0950 hrs A Trp had 1 APC hit by RPG, C Trp had 1 tank hit by RPG.  A Trp has 4 WIA.
           3 are litter and 1 ambulatory at XT527356.
1034     Air Force on station for air strike at this time.
1048     D Trp ship at A Trp location to dust-off WIA.  Picked up 2 litter & 6 ambulatory at XT527356.
1103     A Trp dust-off is complete
1111     C Trp commander wounded and dusted-off.
1115     Air Strike is complete.
1121     D Trp gun team expended 50% of basic load in area of contact.  Mission complete.
1141     Air strike is on station.
1151     Air strike is complete.
1158     A Trp on move out of contact area.
1211     A Trp & C Trp receiving fire.
1229     D Trp gun team on station.
1245     Air strike on station for XT490337.
1248     A Trp received to 2 RPGs from right flank.  2 tanks hit.
1258     1 A Trp tank is on fire.
1300     A Trp location is XT517346.
1305     Arty mission is complete.
1310     A Trp has 6 litter, urgent.
1312     Dust-off for A Trp ETA in 15 min.
1314     D Trp will handle A Trp dust-off.
1315     Update to Div G3, squadron has 24 WIA.
1317     Air strike is complete.
1319     D Trp 39 has completed dust-off for A Trp.
1340     D Trp 37 has completed dust-off for A Trp.
1410     A Trp rpts has APC with thrown track (A10).
1420     Air strike at XT529348.
1459     A Trp at XT500344.
1534     Air strike on station.
1638     A Trp closed NDP.
1642     A Trp trains arrived at NDP.
1752     A Trp NDP at XT459372.
1820     MG Williamson directed troops to have stand-down at their NDPs for maintenance effect 0700
2010     A Trp LP status: A10 on 280 degrees at 100m, A20 on 215 degrees at 75m, & A30 on 353 degrees
           at 100m.
2015     A10 LP departing NDP.  A20 LP is in position
2045     All A Trp LPs in position.
2330     A Trp has a negative SITREP.
Jan 18
Summary:  Squadron conducted resupply and maintenance. Lt. Moore acting commander. Colonel McGowan drops off A Troop Cdr after wounded were visited in hospital.
0022     A Trp has a negative SITREP.
0655     A Trp stand-to complete.
1315     Division Chaplain arrived to conduct services.
1346     There will be an awards team that will go with A Trp on tomorrows operation.
1400     A Trp will test fire M2, M60, & M16 on 90 degree azimuth.
1850     A Trp requested permission to test fire 1 M2 on 120 degree azimuth.  OK from CPT Fullerton.
1900     A Trp change in frequencies: Trp is 36.25 & A30 is 36.90.
1930     A Trp LP status: A10 on 3000 degrees at 100m, A20 on 260 degrees at 75m, & A30 on 110 degrees
           at 75m.
1945     Warning Order from Div G3: 3/4 Cav will be prepared on order to reinforce Trang Bang or Go Dau Ha
           effective immediately.  Occupy old FSB Hampton location for CP and relieve 2/12 CAV 1st Cav Div,
           Sweep TL 19.
2000     A20 LP departing NDP.
2008     A10 & A20 LPs in position.
2013     A30 LP in position.
2330     Message from G2: 3 Bns of NVA armed with AK47, RPGs, & 4 75mm recoilless rifles were reported
           passing Ben Cui III at 2300hrs at XT450450 traveling N to S into the Boi Loi Woods.
Jan 19
Summary: Squadron departs Bo Loi Woods. A Trp escorts Sqdrn CP.  A Troop leads out and same tank that hit mine going into Bo Loi, hit mine going out of Bo Loi.
0842     A Trp be prepared to have 1 platoon at FSB Hampton.
0854     A Trp commencing reconnaissance-in-force.
0927     A Trp has found 1 hooch and several bunkers (1 6x6' & 1 6'x8').  Will blow with Bangalore Torpedo.
1200     A20 element enroute to FSB Hampton.
1210     A20 at CP12.
1222     A20 cleared CP A.
1230     A20 cleared CP18.
1237     A20 cleared CP06.
????      A Trp security plt at CP05 enroute to XT413241.
????      A20 at CP20.
1453     A Trp location is XT413241.
1542     A Trp lead CP12.
1600     A Trp lead CP18.
1607     A Trp lead at CP06.
1630     A Trp lead at CP20.
1636     A Trp and Sqdn CP closed FSB Hampton at XT413241.
1755     Battery B 3rd Bn 13th Arty closed FSB Hampton.  A Trp sent a scout squad to secure remainder
           of B/3/13.
1930     A Trp has linked up with the Arty Bn trains.
2030     Late entry: A Trp and Arty closed at 19:45 hrs.
2045     A Trp mission for tomorrow: provide security for sqdn CP and trains at FSB Hampton, be prepared
           to assume responsibility for RT 1 on 21 Jan.
2330     A Trp has a negative SITREP.
Jan 20
Summary: Battery B 3rd Bn 13th Artillery at FSB Hampton is placed in DS support of 3/4 Cav. Battery B supported the Cavalry in numerous contact missions. On many occasions a preparation was fired prior to the Cavalry entering a specific area. This was used very effectively to soften the enemy' s position. The surveillance for the period of 25 January to April 3 1969 included a body count of 127 NVA KIA; 30 NVA wounded; numerous blood trails, and 15 bunkers, several machine gun positions, and 600 meters of trench line destroyed.  A Trp secures FSB Hampton.
0015     A Trp has a negative SITREP.
19??     A30 requests test fire on 40 degree azimuth.
1940     A30 LP on 60 degree azimuth at 50m.
1956     A Trp LPs in position.
2100     A Trp spotted movement 200m on 90 degree azimuth from their position.  Request permission
            to engage.  Granted by Go Dau Ha Sub-sector.
2330     A Trp has a negative SITREP.
Jan 21
0030     A Trp has a negative SITREP.
0530     A Trp stand-to is complete.
0711     A10 on the move.
0730     A10 is at CP19.
0736     A Trp is at CP07 and doing a 180 degree turn.
0737     A Trp rpts that road is clear from CP20 to Trang Bang.
0755     A10 is at CP19 doing a 180 degree turn.
0825     A10 is at CP07 headed up alternate MSR.
0831     A 10 is at CP 19.
0847     A10 is at CP20.
0900     A10 is at CP05.
0918     A10 is at CP22 making 180 degree turn.
0926     A10 is at CP04.
0932     A10 is at CP21.
0940     A10 closed squadron CP location.
1245     A Trp escorts and lowboy departing for C Trp's location.
1315     A Trp escorts and lowboy at C Trp's location.
1421     A Trp escorts and lowboy have returned to FSB Hampton.
1630     A Trp escort and lowboy have returned to FSB.
1636     A10 departing FSB Hampton enroute to its NDP (XT487347).
16??     A10 has closed its NDP.
1845     ARVN's have requested a possible reaction force for a unit in contact.
1846     A10 will return to FSB Hampton.
1915     A10 is at CP05.
1920     A10 is at CP20.
1925     A10 elements back at FSB Hampton.
2035     A Trp reports lights on 90 degree azimuth at 200m.  Continuing to observe.
2206     SPOTREP: LA A35; LB 2 individuals; LC XT417242 at 2206; LD 2 individuals walking SE to NW
            200m out near woodline; LE observing w/starlight, request permission to engage.
2216     Permission to engage with small arms only granted by ARVN.
Jan 22
Summary:  Colonel McGowan notified Troop Cdr that Lt Ayers (photo/shadowbox) died at 12th Evac from wounds received on Jan 17th.
0005     A Trp has spotted movement 200 meters out on 40 degree azimuth.  Engaging.
0034     SPOTREP: LA A30; LB movement; LC 40 degree azimuth 400 meters; LD spotted with starlight &
           engaged with 50 cal; LE continuing to observe.
0530     Stand-to is complete.
0650     A20 departing FSB Hampton.
0655     A30 on the move.
0703     A20 at CP19.
0708     A30 at CP05.
0725     A20 cleared CP07.
0800     A30 at CP18 & doing a 180 turn.
0814     A20 has met up with C Troop.
0830     LTC McGowan directs that all 45 cal pistols be put in 50 cal ammo cans that will be painted red.
           All personell 1SG & up will be authorized a 45 cal.  All others will turn theirs in and tag them.  Will be
           put at plt level.
0836     A30 at CP22.
0840     A20 at CP19.
0843     MSR swept & clear.
0849     A20 entering FSB Hampton.
0853     A30 at CP05.
1205     A Trp will conduct RIF in location that C Trp was suppose to go.
1338     SPOTREP: LA A20; LB bunker; LC XT444255 at 1335 hrs; LD blown in place; LE cont mission.
1345     SPOTREP: Saber 3; LB 5 personnel & 90 cows; LC XT441259 at 1343 hrs; LD moved northeast
           when A Trp entered the area.
1445     SPOTREP: LA A20; LB bunker & tunnel; LC XT447257 at 1435 hrs; LD 4'x4', overhead cover, 2
           bunkers connected by tunnel to 3rd & 4th bunkers, sleeping area; LE blown in place.
1610     SPOTREP: LA A20; LB booby trap; LC XT453254 at 1605 hrs; LD exploded/tripped in trees; LE     .
           negative damage.
1615     A20 on road headed to CP18.
1720     A30 at CP18.
1728     A30 at CP06.
1741     A30 at CP05.
1750     A30 had a tank hit a telephone pole at CP05.  Plt Ldr in ARVN compound at this time.
1810     Tomorrow A Trp will have 2 plts on RIF from CP18 north through the wood line, then swing SE to
1815     A30 has closed FSB Hampton.
2225     A Trp has a negative Sitrep.
Jan 23
Summary: Troop establishes NDP at FSB Hampton (photo). Troop clears road. Troop conducts a reconnaissance-in-force at XT452270.
0030     A Trp has a negative SITREP.
0555     Stand-to is complete.
0653     A10 starting road sweep at CP20.
0706     A20 at CP06.
1013     A Trp at C Trp's NDP from last night awaiting instructions.
1150     A Trp at XT452270.  Negative findings at this time.
1410     A Trp at XT468280.
1632     A Trp entering NDP at FSB Hampton.
1642     A Trp completely closed in FSB Hampton.
2025.    SPOTREP:  LA A30; LB movement; LC at 2025 hrs; LD movement around fire on 170 degree azimuth
           at 200 meters, possibly burning trash; LE continuing mission.
2319     A Trp has a negative SITREP.
Jan 24
Summary: Troop departs NDP to sweep MSR from CP22 to CP07. Troop conducts a reconnaissance-in-force at XT496927. Locates 10 tunnels & fighting positions. Destroys tunnels and returned to FSB Hampton at 1652 hrs.
0021     A Trp has a negative SITREP.
0545     Stand-to is complete.
0653     A30 departing FSB Hampton.
0658     A10 departing FSB Hampton.
0703     A30 starting road sweep at CP20.
0711     A10 starting road sweep at CP19.
0740     A10 sweep complete & returning to FSB Hampton.  A30 is 1km south of A10.
0817     A10 entering FSB Hampton.
0832     A30 entering FSB Hampton.
1126     A Trp sweep at CP05
1143     A Trp lead at CP18.
1242     A Trp at XT464272 cartwheeling left to the north, negative findings.
1320     A Trp at XT467278, negative SITREP.
1406     A Trp at XT481294.
1508     A Trp at XT487274.  Has found tunnels.
1628     A Trp at CP18.
1652     A Trp entering FSB Hampton.
2230     A Trp has a negative SITREP.
Jan 25
Summary: Troop secured road in support of B Troop operation in the Parrot's Beak.  NVA unit infiltrated a hamlet by coming across the Cambodian border and froming a  strong hold. Troop worked Southwest of Go Da Ha to secure a series of small bridges and culverts on roads to Cambodian border. Two to three tracks outpost a bridge/culvert.  Two to three more outpost next bridge/culvert. (photo) Illumination fired in support of B Troop NDP where major battle ensued.  B Troop broke the NVA stronghold.
0126     A Trp has a negative SITREP.
0545     A Trp stand-to is complete.
0735     A20 at CP05.
0752     A Trp to send 2 vehicles to pick-up AVLB at FSB Stewart.
0919     A20 is at CP18.
0951     A Trp to provide security on the Go Dau Ha Bridge.
0954     A Trp RIF elements to return to FSB Hampton for possible security mission.
1020     1 A Trp plt's tanks will secure the 3 bridges on QL1 west of Go Dau Ha.
1140     Cdr A Trp rpts 3 bridges are out posted on the route of advance.
1142     A Trp will be reaction force to B Trps operations.
1505     A Trp requested dustoff for 2 VN civilians at XT348219.  1 severe leg wound & 1 fractured leg.
           Dustoff 150 will respond.  ETA 15 min.
1555     A Trp Cdr & A10 are at CP05.
1630     SPOTREP: LA VN civilian; LB mine in road; LC XT350224 at 1610 hrs; LD 40/50 lb mine on culvert. 
           Culvert not damaged, crater 6'x4', killed one water buffalo & 1 civilian when ox cart hit mine; LE has
           been there for a while.
1640     A Trp APC hit a mine.  Need urgent dustoff for 1 litter at XT352226.  Also 2 ambulatory.
1646     A Trp dustoff is complete.  APC is combat loss.
1807     A20 in NDP XT355226.
2015     A Trp has a negative SITREP.
Jan 26
Troop continues support of B Troop operation in the Parrot's Beak.
0024     A Trp has a negative SITREP.
0556     A Trp stand-to is complete.
0735     A Trp beginning road sweep.
0748     A Trp at XT358243.
0831     A Trp at XT352225.
0848     A Trp 300 meters from Bridge.
1031     A Trp has a individual who is 28 years old and not in the Army.  Will take to the Go Dau Ha Sub
1143     A Trp has engineers and sweeping each bridge as they go.
1430     A Trp reports engineers have returned to put laterite on the road.
1542     A Trp mortar tracks have turned around and headed back to Cu Chi.
1553     Engineers with A Trp on way back to squadron CP.
1835     A Trp has picked up the convoy at this time.
1850     A Trp given permission to test fire small 2 50cal MG on a 240 degree azimuth from XT354248.
1915     SPOTREP: LA A20: LB fireing at moving vehicles; LC XT350224 at 1910 hrs; LD 1 lambretta & 1
           motorcycle moving north toward plt, were being fired on by ARVNs, they saw A Trp and turned
           back; LE continuing to observe.
2227    A Trp has a negative SITREP.
Jan 27
Troop continues support of B Troop operation in the Parrot's Beak. Squadron redeploys to MSR.
0035     A Trp has a negative SITREP.
0545     Stand-to is complete.
0802     A Trp is at XT350224.
1010     A Trp will check out wood line to their front.  Then they will hook back & prepare for move to the
1529     A Trp has linked up with B Trp and are on the way in. Squadron redeploys to the MSR.
1543     A Trp at CP05.
1603     A Trp has closed Command Post.
1852     HHT, Cmd Post, A Trp & B Trp are at FSB Hampton.
2320     A Trp has a negative SITREP.
Jan 28
Summary:  NDP in FSB Hampton.  Sweep MSR then conduct RIF in small rubber plantation with 2 plts on left flank of B Trp.
0022     A Trp has a negative SITREP.
0605     Stand-to is complete.
0709     A10 will sweep to CP07, A20 will sweep to CP18 and back to CP22.  Sweeps moving out now.
0740     A20 at CP 06, A10 at CP07 and returning.
0741     A10 at CP19.
0745     A20 at CP18.
0752     A20 at CP08.
0755     A20 at CP20.
0815     A20 at CP04.
0822     A20 at CP22.  Visibility only 50 meters.
0844     Joint operation with US Advisors/Go Dau Ha Sub-Sector.  A & B Trps depart FSB Hampton enroute
           to small rubber for RIF in plantation from Hwy QL22 in a northeast direction.  The 2 plts of A Trp
           on the left of B Trp.
0851     A10 at CP04 linking up with A20.
0901     A Trp lining up with B Trp at CP21.
1005     A Trp has come upon RF/PF forces working in the small rubber plantation.
1019     A Trp is at XT374334 and B Trp at XT378330.
1128     A Trp is sending two vehicles down to pick up squadron trains at Trang Bang.
1156     A & B Trps on line at XT385328 conducting Search & Destroy mission.
1328     A10 and A20 being sent back to FSB Hampton.
1341     A Trp lead now on the road.
1430     A Trp now in NDP.
1543     A30 departing FSB to A Trp NDP at XT452275.
1800     A Trp NDP at XT455267.  A10 at XT458252.
1810     A30 enroute to outpost culvert at XT458252.
1835     A30 closed A Trp NDP.
1945     A Trp test fire 1 50cal, 1 m60, & 1 90mm on 75 degree azimuth from XT458252.
1953     A10 test 1 50cal on 45 degree azimuth.
2018     A30 test fire complete.
2123     SPOTREP: LA A30; LB moving lights in wood line; LC 120 degree azimuth from CP B at 500 meters,
            2125 hrs; LD observed lights in wood line: LE request permission to engage with 50cal.  approved.
2155     A Trp engaged moving lights.  Negative results.
2310     SPOTREP: LA A20; LB movement; LC 245 degree azimuth at 200 meters from XT455267, 2300 hrs;
            LD 2 individuals with water buffalo herd; LE permission to engage given by ARVNs.
2330     Negative SITREP.
Jan 29
Summary:  A Trp sweeps and clears MSR. Outpost MSR & escort convoy.
0030     Negative SITREP.
0605     A Trp stand-to complete.
0654     A30 begins sweep of MSR.  A10 to sweep to CP07 and then escort convoy.  A20 to sweep to
           CP25.  A30 to outpost CPB, CPR & CP12.
0751     A Trp has a tank with thrown track near CP12.
0804     A Trp trains departing NDP for command post.
0815     A20 at CP05 and has begun outposting.
0818     A Trp sector of MSR is clear.
0838     A Trp trains entering the FSB.
0935     A10 elements with the convoy.
1252     ETA of convoy at Tay Ninh 1300 & at Dau Tieng 1320.
1630     A Trp trains departing FSB for NDP.
1646     A Trp elements have closed NDP.
1650     A Trp (A10 & A30) NDP at XT455267.  A20 at XT466251.
1831     A10 in NDP.
2105     A Trp radar is operational: L 2300 mils, R 4600 mils, center 3450 mils, range 500m.
2330     A Trp has a negative SITREP.
Jan 30
Summary:  A Trp sweeps and outposts the MSR.
0025     A Trp has a negative SITREP.
0355     A20 requested to engage movement at XT457268.  Permission granted.
0409     A20 engaged w sm arms & movement stopped.  Will check area at first light.
0643     A Trp sweeps:  A10 from NDP to CP07, A20 from NDP to CPB, & A30 from NDP to the north.
0745     SPOTREP: LA A Trp; LB road obstacles; LC XT469210 at 0740 hrs; LD soda cans & trash on road
           100 meters in length for width of road, propaganda leaflets on stakes, no booby traps; LE clearing.
0801     A Trp trains enroute to command post.
0822     Ref the road obstacle.  It included one mine electrically detonated.  Blown in place left a crater
           4' deep & 4' wide.
0830     SPOTREP: LA A Trp; LB dud 105 rd; LC XT480219 at 0830 hrs; LD blew in place; LE cont mission.
1239     A Trp element escorting AVLB reached FSB Stewart.
1347     A Trp picked up the convoy at CP36. 5 MU from Tay Ninh & 2 from Dau Tieng.
1549     A Trp trains departing FSB to NDP.
1649     A Trp NDP is at XT456263.
1830     A20 will strong point culvert at XT458252.
1913     A Trp requested permission to test fire M60, M16 on 90 degree azimuth from XT458252. Permission
1920     A Trp test firing 1 50cal. on 55 degree azimuth from NDP.  Cleared by ARVNs.
1948     A Trp radar: L 2600 mils, R 4200 mils, C 3400 mils at 500 meters.
2232     A Trp has a negative SITREP.
Jan 31
Summary:  A Trp sweeps and outposts the MSR.  Escorts convoy.
0025     A Trp has a negative SITREP.
0532     A Trp stand-to complete.
0558     A10 leaving the radio net.
0710     A30 is on the move.
0712     A Trp leaving FSB.
0714     A20 starting sweep.  A30 is at CP18.
0721     A30 is at CPA.
0728     A30 is at CPD.
0733     A30 is at CP12.
0742     A10 is at CP18.  A20 is at CPR.
0805     A Trp seeep is at CP36.  Trains are at CP06.
0822     A22 at CP07 making a 180.
0828     A Trp sweep at CP25.  A Trp prepare VTR for recovery of vehicle that has run out of diesel.
1140     A Trp outposting MSR.  A10 CP07 to CPB, A30 at CP25, & A20 with convoy.
1300     Warning Order:  From 25th ID mission change for 3/4 Cav, move & establish NDP north of Go Dau
           Ha.  Orient & protect Go Dau Ha from the N/NE/E.  Upon return of C Trp be prepared to establish
           NDP Feb 1 at XT3419.  On order mission.
1340     A Trp lead is at CP12.  Trail is at CP36.
1621     A Trp trains departing FSB for NDP.
1705     A Trp NDP at XT455266.
1905     A Trp requests permission to test fire one 50 cal, & 1 90mm on 65 degree azimuth from NDP. 
           Permission granted by ARVN.
2026     A Trp radar status: L 2100 mils, R 5200 mils, C 3650 mils, & range 500 meters.
2325     A Trp has a negative SITREP.
Feb 1
Summary: Cpt John Wells arrives at Troop and spends the day and night with Cpt Moss.  A Trp sweeps and outposts the MSR.  Escorts convoy.
0025     A Trp has a negative SITREP.
0100     SPOTREP: LA A20; LB trip flare set-off; LC XT465253; LD engages with 50cal & M73; LE no return
            fire, continuing mission.
0534     A Trp stand-to complete.
0608     A20 will oupost from CP07 to CP25.
0650     A30 starting sweep.
0706     A30 at CP05.
0711     A20 is on the move.
0715     A10 is on the move.
0724     A10 is at CPA.  A30 is at CP06.
0729     A30 is at CP18.
0733     A10 is at CPD.
0735     A Trp trains departing NDP for FSB.
0739     A10 is at CP12.
0741     A Trp trains at CP18.
0750     SPOTREP: LA A27; LB RVN flag & propaganda; LC XT467237 at 0735 hrs; LD flag torn into 25
            pieces and propaganda in English against the war; LE continuing mission.
0806     A Trp sweep at CP06.
0808     A Trp trains at FSB.
0819     A10 sweep at CP36 and completed outposting at this time.
0825     A20 at CPZ.
0830     A20 sweep complete and sector is clear.
0849     A20 outposting 2 veh at CPB & CPR.
0947     A Trp element with convoy at CP16.
0953     A Trp lead at CPA.
0955     A Trp lead at CPD.
1122     A Trp trail at CP32.
1355     Tay Ninh convoy at CP25.
1402     Dau Tieng convoy at CP12.
1625     A30 closing into NDP.
1628     A Trp(-) departing for NDP.
1635     A Trp at CP05.
1642     A20 at CP20.
1646     A20 in FSB.
1653     A Trp at CP18.
1700     A Trp(-) at XT452266.  A Trp strong point at XT458252.
1818     A10 in NDP.
1917     A Trp radar status: L 2000 mils, R 4200 mils, C 3100 mils, & range 500 meters.
2325     A Trp negative SITREP.
Feb 2
Summary: Cpt Wells assumed command. A Trp assigned to convoy duty on MSR. Major battle near the Ben Cui Plantation (see map). C Trp makes contact around 10:00AM still in major contact at 14:30. MG Williamson releases A Trp from convoy duty to reinforce right flank of C Trp. Abt 17:00 hrs B Trp release from security of FSB Hampton & moves up on the flank of A Trp. ARVN Marines reinforce C Trp. Squadron assaults online. RPG fire from bunkers across entire front. 1 A Trp & 2 B Trp APCs burning. 23 WIA, 3 KIA, & 1 man missing. Squadron withdraws across clearing to south. ARVN Marines make night assault at midnight and find enemy has withdrawn. PFC Haake, missing man, spent night hidden on enemy position. SGT Feeney, (photo/shadowbox) PFC Ramsey, (photo/shadowbox) PFC McKinnon (photo/shadowbox), & CPL Callister (B Trp) KIA.
0023     A Trp has a negative SITREP.
0535     A Trp stand-to is complete.
0700     A30 starting to sweep north of CP25 & establishing outposts.  A10 will sweeep from CPB to CP07 &
            outpost.  A20 sweeps form FSB to CP18 & escorts convoy.
0704     A20 leaving squadron net.
0706     A30 at CP18.
0711     A10 starting sweep. A30 is at CPA.
0720     A30 is at CP12.
1032     A Trp rpts that nobody is at CP12 to pickup the convoy.  Maj. Harrison says to let them turn & go
            without us.
1034     C Trp has flushed out approx 6 VC by Inf at XT388314, also found 5 foxholes.
1048     C Trp receiving fire from left flank.
1141     C Trp Dustoff, 3 ambulatory , shapnel US, XT393313.
1205     C Trp receiving RPG & SA fire.
1230     Dustoff ship hit while landing.  Pilot has light shrapnel wounds, crew chief has broken arm.  Rtn to
            Dau Tieng.
1335     Airstrike going in at this time.
1400     C trp began to sweep again, received auto fire RPG's, pulling back, 1 APC hit.
1408     Air stike going in at this time.
1428     MG Williamson gave permission to remove A Trp from convoy and put them in area of contact. A
           Trp road marches to Go Dau Ha turns off the road and travels 2 miles Northeast to link up with the
1615     C trp needs 50 cals, has 5 nonoperable, needs 4 M60s , has 10 main gun rds left, needs 2 coax
            for his tanks.
1633     A Trp arrived on line w/C trp, made contact. 
1630     Troop on reconnaissance-in-force.
1655     4th airstrike going in at this time.
1713     5th airstrike will be on station at 1730 hrs.
1744     B Trp moves up into the area of contact.
1751     FAC bringing up another strike.
1759     B Trp is moving up onto A Trp flank.
1837     A Trp has 1 man w/frag wound in head from air strike, will dustoff w/Muleskinner.
1848     A Trp has 3 Public Information Officers that would like to get out, neg.
1856     ALL Trps moving out on line. Entire squadron is on line.  Receive heavy fire from bunker complex in
           woodline. Troop assaults woodline.
1908     A Trp has an APC hit w/RPG, unk damage or casualties.
1912     RPG hit a tree to front of A Trp.
1914     A Trp reports another RPG. Sgt Finney, TC A14, (photo/shadowbox) is KIA attempting to clear M2 50 Cal.
           A14 takes 100mm recoilless rifle through the side. A14 also hit by several RPG's. 3 crewmembers were
           blown out the backdoor and wounded.  Platoon medic, "Doc" Jones, went to their aid and while
           treating one of them had his leg run over as Cpl Davenport, A14 driver, who was wounded and firing
           at the enemy tried to back the burning APC out of the kill zone.  Ammo catches on fire as A14
           burns. PFC McKinnon is KIA (photo/shadowbox) when M60 mg he is holding is hit by an RPG and explodes
           while he is on the backdeck of A35. A 2nd RPG hits vehicle and blast knocks PFC Roger Haake off the
           backdeck.  PFC Haake initially unaccounted for and spent night in enemy positions.  See Tropic
           Lightning News Article March 3, 1969.  PFC Ramsey is KIA (photo/shadowbox). Troop begins to withdraw
           back across clearing with casualties and establishes night defensive position. A14 continues to burn most
           of the night.
1915     B Trp has 2 tracks burning.
1920     A Trp while pulling back receiving fire from right flank.
1939     ALL Sqdn elements located West of N-S road vic XT391301 (Northeast of HWY QL22).
1940     At 1950 hrs a "spookie" (C130 gunship) will be on station.
1951     A Trp has 4 litter w/ C Trp for dustoff purpose.
2008     B Trp has 2 wounded, D Trp 73 w/dustoff.
2019     D trp 73 dustoff: 4 litter enroute to Cu Chi 2 ambulatory.
2020     FAC 14 up for air strike.
2038     D Trp 63 picked up 6 personnel, 5 ambulatory & 1 litter.
2106     "Leatherneck" will make an assault to the south.
2142     ISSUE 16 FAC up on station, Spooky 76 on station, 1 - 2 hrs stations.
2150     Total 30 BC KBA.
2225     Air strikes are rolling in at this time.
2245     Air strikes complete, another on station.
2353     ARVN Marines moving out on assault to south.
2356     A Trp has 5-6 machine guns they need in field, also get request for vital resupply in morning, have
           slicks on call from D Trp.
2400     Journal closed.
Feb 3
Summary: Continuation/wrap-up of major battle near the Ben Cui Plantation  (see map). A Trp evacuates KIAs from previous day. Returns to FSB Hampton.
0046     ARVN Marines make night assault on enemy positions.
0050     ARVN Marines have crossed the north/south road.  Are 30 meters from area of contact.  Negative
0120     Leatherneck 1 reports 4 BC and 1 captured ChiCom machine gun.
0135     A Trp has a negative SITREP.
0546     A Trp has completed stand-to.
0830     3/4 Cav elements, 3/22 Inf & ARVN Marines starting to sweep area of contact.
0843     A Trp has 1 KIA to evacuate and 1 urgent dustoff of man with shrapnel in leg.  D
           Trp will provide dustoff in 15 mins.
0922     ARVN Marines have found 12 VC KIA & 1 VC POW along with 1 50cal, 1 BAR, 1 75mm Recoilless rifle.
1022     D Trp enroute to A Trp to pickup 2 KIAs.
1200     All Trps moving back to FSB Hampton.
1234     A Trp closing on FSB Hampton.
1530     A Trp departing FSB.  Lead at CP05.
1535     A Trp trains departing FSB.
1547     A Trp reports that between CP05 and CP06 people openly hostile. Also wearing blue shirts & pants
           similar to those in last night's contact.  Several people in Go Dau Ha w/possible frag wounds.  A Trp
           lead at CP06.
1555     A Trp at CP 18 & trains at CP6
1600     A Trp trains at CP18.
1605     A Trp trains in NDP.
1637     A20 & Trp HQ at NDP, XT455270.
1919     A Trp has a strong point at XT458252.
2049     A Trp radar status: L 2400 mils, R 5400 mils, C 3900 mils at 500 meters.
2127     A Trp has a negative SITREP.
Feb 4
Summary: A Trp sweeps and outposts the MSR.  Escorts convoy.
0030     A Trp has a negative SITREP.
0431     A Trp stand-to is complete.
0654     A10 sweeps from FSB to CP05 to CP18.  Then A10 escorts the convoy.  A20 sweeps from NDP to
           CP 25 & outposts MSR.  A30 sweeps from CPB to CP07 & outposts the MSR.
0743     A Trp trains departing NDP for FSB.
0812     A Trp trains at FSB.
0902     A20 completed sweep & outposted.
0932     A Trp observed a burial at XT443273.  Several indiv rode away on bicycles & others entered
1018     SPOTREP: LA A Trp; LB picutre; LC XT422266 at 1018 hrs; LD picture against hootch of chopper
           being shot down.
1100     A Trp reports MP jeep turned over 600 meters east of CP25.  1 man medevaced through other
1105     SPOTREP: LA A25; LB convoy vehicle hit mine; LC XT348448 at 1040 hrs; LD flat bed hit mine,
           slight damage, blown tires, 3x3' crater from 30/35 Lbs mine; LE changing tires.
1436     A Trp reports vehicle hit mine.  Negative damage, vehicle is able to move on.
1547     A Trp trains departing FSB.
1615     A Trp trains entering NDP.
1655     A10 & Trp Hq NDP at XT453263.  A20 will be at CPB & A30 will be in FSB.
1800     A20 departing for strong point.
2330     A Trp has a negative SITREP.
Feb 5
Summary: A Trp sweeps and outposts the MSR.  Escorts convoy.  Returns to Cu Chi for stand-down.
0028     A Trp has a negative SITREP.
0535     A30 stand-to is complete.
0544     A Trp(-) stand-to is complete.
0640     A20 sweeps from NDP to CP07 & outposts to CPR/CPB.  A10 seeeps from NDP to CP25 & outposts.
           A30 sweeps from FSB to CP18 & picks up/escorts convoy.
0744     A Trp trains departing NDP for Cu Chi.
0920     A10 has completed sweep and is outposted.
0928     Centaur 42 rpts 2 personal lying in culvert at XT362432.  A Trp will check it out.
1017     A Trp rpts a cemetery at XT400393 where they are burying 8 people.
1131     A Trp trains have entered Cu Chi.
1633     A Trp departing for move back to Cu Chi.
1748     A Trp entering motor pool at Cu Chi.
2400     Journal closed.
Feb 6
Summary: A Trp at Cu Chi for stand-down.
Feb 7
Summary: A Trp at Cu Chi for stand-down.
Feb 8
Summary: A Trp at Cu Chi for stand-down. Formation and awards ceremony at squadron HQ.
Feb 9
Summary: A Trp at Cu Chi for stand-down.
Feb 10
Summary: A Trp returns to operations along MSR. Conducts RIF & outposts MSR. Troop operates without M48 Tanks that have been turned in at Cu Chi for the replacement M551 Sheridan tanks. Tank crewmen given option to transfer to other tracks or troop headquarters.  M551 trained crewmen had started arriving as replacements as early as the last week of December 68. M551s picked up at Long Bhin and driven back to A Troop motor pool (photo). Tank crews undergo M551 training for several weeks.  See article on M551 training in the Mar 3 issue of the Tropic Lightning News. See a crew qualification patch here given to crews on completion of training. Crews modifiy vehicles to meet the needs of the field. Bustle racks were constructed and welded on with help of 65th Engineers.  Additional racks added to hold more 50cal ammo. Mounts for a 2nd M2 50 cal machinegun added to the front of the loaders hatch. The TC's overide was moved up to the right side of the TC hatch so it could be fired by the TC riding high in the cupoloa. The gunners rode in the bustle rack on the outside and guarded the rear of the vehicle with an M16 rifle.
xxxx      A Trp departs Cu Chi.
0907     A Trp holding at CP18.
1045     A Trp conducts RIF vic XT442265.
1055     A Trp has found some spider holes, placing dismounts out.
1125     A Trp at XT445255.
1150     SPOTREP:  LA A Trp; LB tunnel & bunker; LC XT438254 at 1150 hrs; LD 1 tunnel 10' long, 2 bunkers
           2x2'x3', 1 bunker 4'x4'x4'.  All about 3 or 4 months old; LE checking area/will destroy.
1153     A Trp found a dud 155mm rnd at XT438253, no fuse.  Blown in place.
1219     SPOTREP:  LA B23; LB 2 bunkers; LC XT439251 at 1215 hrs; LD 4'x4x2' about 3 to 3 wks old, no
           activity: LE blown in place/continued mission.
1310     A Trp at CP B and outposting.
1320     A Trp will take right flank of MSR and make sweep through at CPB.
1534     A Trp trains departing FSB for NDP.
1603     A Trp(-) NDP at XT453263.  Plt strong point at XT459253.
1623     A Trp trains entering NDP.
2030     A Trp radar status: L 5600 mils, R 2400 mils, C 800 mils at 750 meters.
2135     A Trp has spooted 5 personnel moving in woodline at XT450262.  Permission to fire given by Bao
           Don advisor.
2330     A Trp has a negative SITREP.
Feb 11
Summary: A Trp sweeps and outposts the MSR.  Escorts convoy.  Establishes NDP and 1 platoon strongpoint.
0021     A Trp has a negative SITREP.
0535     A Trp stand-to is complete.
0654     A30 is on the move.  Will sweep CPB to CP07.
0704     A10 is on the move.  Will sweep CP18 to CP36.
0713     A10 is at CP18.
0723     A30 is at CP R.  A10 is at CP A.
0727     A10 is at CP D.
0733     A10 is at CP12.
0745     A30 is outposting.  Has a roving partol of 2 vehicles from CP O to CP S.
0958     A20 is escorting convoy.  Lead is at CP O.
1158     A Trp section has been extended to CP25.
1222     A Trp has relieved unit at CP25 at this time.
1638     A Trp(-) NDP at XT455266.  1 platoon strong point at XT458252.
2055     SPOTREP 1: LA A Trp; LB automatic weapons fire; LC XT440238 at 2040 hrs; LD 2 trip flairs and 50 
           cal fire to the NW; LE continuing to observe
           SPOTREP 2: LA A Trp; LB automatic weapons fire; LC XT456252 at 2040 hrs; LD 2 unk cal fire on a
           170 degree azimuth; LE continuing to observe.
2316     A Trp has a negative SITREP.
Feb 12
Summary: 1st platoon sets up night laager at XT458235 (CP B) near X. Gia Binh on Hwy LTL 19. Lt Moore, SGT Blanding, SP4 Harwell, and platoon medic wounded by booby trap set off in position. SGT Hawley calls in medevac.
0020     A Trp neg sit rep
0115     Perimeter neg sit rep, A Trp neg sit rep
0220     SPOTREP: LA-A Trp, LB-Movement LC-470308, LD-Spotted movement on 180º 2 O’Clock set off a
           starcluster Neg further
0225     A and C Trp neg sit rep
0430     A and B Neg sit rep
0525     B & C Trp neg sit rep
0535     A Trp stand to complete
0705     A Trp moving out on sweep
1400     A Trp 10 element at CP 18
1410     A Trp 10 element at CP B
1420     A Trp 20 lead of convoy at CP A
1433     A Trp 20 and convoy at CP06
1441     A Trp 20 w/convoy at CP05
1521     A Trp and convoy trail cleared CP18
1533     A Trp and convoy trail at CP 05
1540     Be Prepared Mission Rec’d From Div be prepared to assemble SQdn via Go Da Hau as a reactionary
           force, no move req’d at this time.  Note: US Aircraft has gone down on Cambodian side of Angel’s
           Wing A/C orbiting the area.  Tracked trying to receive permission to extract A/C and crew.  If
           necessary, ¾ Cav will be ground reaction force. A/C Location XT274229.
1540     A Trp trains at CP06 A 30 trail at CP 18
1550     A Trp trail at CP 06
1553     A Trp trains at CP 18
1557     A Trp trains at NDP
1600     A Trp convoy trail CP20
1605     A Trp 20 & 30 NDP XT455266 A10 at CP B No LP’s
1646     3rd Plt, Trp A closed NDP
1741     A Trp had a booby trap go off at CP B 4 litter 1 amb.  Centaur 63 is dustoff
1745     Centaur 63 off Hampton to CP B.
1750     Dustoff complete for A Trp 2 litter
1751     Cent 73 enroute to CP B for 1 litter
1753     A Trp sending 1 reinforce sub element out to CP B
1758     Dustoff complete for A Trp at 1805 from Maj Harrison to stand down the Aerorifles
1813     SPOTREP: LA A Trp 10; LB booby trap; LC CP B T1715; LD 1 M79 rnd in base of tree results 4 litter
           1 amb; LE Did not know how it was set off.
1820     A Trp 30 CP R
1832     A Trp 1 reinforce element is at CP B. Engineers down checking out area.
1940     Cent 63 off from CP B
2028     A, B, C, Neg sit rep
2125     A, C, perimeter neg sit rep also C Trp LP’s in pos.
2223     A, C Trp Neg sit rep
2231     Perimeter neg sit rep
2327     A, C Trp neg sit rep
Feb 13
0030     All elements sit rep neg
0045     From rear toc at 122350 ARVNs at Trang Bang reported 500 VC vic XT505155 w/1 122 mm rocket
           and launcher and 62 mm mortars Mission to attack Trang Bang 12 Feb. ARVNs C-2 rating Us F-6
0535     A Trp stand to complete
0700     A Trp sweep from CP B CP07, A20 sweep from N DP CP12, A30 sweep from NDP- CP-18-CP05-CP22
0703     A Trp 20 beginning sweep
0704     A Trp 30 sweep started
0709     A Trp 20 at CP18
07--       C Trp 20 departing FSB
07--       ?? at CP18
07--       A10 on move
07--       A Trp 30 at CP06
07--       A Trp 30 at CP04
0735     SPOTREP: A, ATrp 10, B road obstacle C XT463243, 0735, D stretched 100M along road with
           road contained straw covering road with bamboo poles and flags
0745     A Trp 30 at CP22
0755     A Trp 10 has found 2 mines in road obstacle
0807     A Trp trains departed NDP for FSB Hampton
0828     A Trp 30 at NDP from last night waiting on convoy
0830     C Trp eng team to be lifted out to C Trp 20 position at CP07 to help A Trp
0835     A Trp trains closed FSB Hampton
0845     A Trp sweep at CP486207
0851     A Trp sweep at XT480218
0855     A Trp closing on CP07
0931     A Trp reports there is a lot of interfering traffic on their freq
0932     A Trp convoy CP S
0937     A Trp convoy at CP 18
0941     A Trp convoy at CP A
0943     R 50: A Trp reports according to Hoi Chan the traffic is VC also ref to yesterday’s booby trap it
           was a dud M 79 rnd that was picked up by a cook
0946     A Trp convoy at CP D
0955     A Trp convoy at CP 12
1528     A Trp trains departing FSB for NDP
1538     Trains cleared CP 05
1538     Trail convoy CP Z
1541     Trail convoy CP 07
1552     Trains cleared CP 18
1600     A Trp trains reached NDP
1611     A Trp NDP XT455267, 20 vic CP B
2005     SPOTREP: A Trp 20, moving lights XT458244 T, 2000 3 lights 900 meters south in wood line, 50 to
           100 meters apart can be seen by naked eye, Requesting arty
2130     Trp A, Trp C perimeter base camp neg sit rep
2230     All elements sit rep negative
Feb 14
Summary: A Trp outposts the MSR.  Conducts sweep in Boi Loi Woods.  M551s arrive in Cu Chi.
0030     A Trp has a negative SITREP.
0535     A Trp stand-to is complete.
0726     A30 departing NDP.  A10 at CP18.
0755     A10 & A20 cleared CP18 enroute to pick up Bao Don elements.
0757     A30 at CP05.
0802     A Trp trains departing NDP for FSB.
0808     SPOTREP: LA Engr with A Trp; LB propaganda; LC XT468236 at 0800 hrs; LD propaganda letters
           from POW to GIs saying don't shoot civilians or burn homes; LE bringing in & continuing mission.
0811     Sweep finished and beginning to outpost.
0829     A Trp trains entering FSB.
1045     A Trp sweep at XT493302
1137     A Trp found 105mm dud at XT441251, blown in place.
1158     A Trp at XT441247
1834     A Trp Sheridans have arrived.  Training starts in the morning.
2207     A Trp reports lights at XT467256, requests clearance to engage.  Permission granted.
2319     A Trp has a negative SITREP.
Feb 15
Summary: A Trp outposts the MSR.  Conducts sweep in Boi Loi Woods. Establishes NDP and platoon strong point. M551 training starts in Cu Chi. (Last page of journal missing)
0040     SPOTREP:  LA A Trp; LB movement; LC XT450261 at 0030 hrs; LD radar detected one individual
           moving on 180 degree azimuth towards perimeter until 25 meters away; LE sent up flare, saw
           nothing, got another radar hit of personnel moving back to woodline, called in Arty w/unknown
           results, got a third sighting of personal moving towards perimeter engaged with M79, unknown
0535     A Trp stand-to complete.
0714     A20 & A30 starting sweep.
0749     SPOTREP: LA A20 Engr Tm; LB propaganda; LC XT467236 at 0730 hrs; LD paper VC flags & leaflets
           with statements from GIs about war on poles along road; LE picking up/continuing mission.
0821     A Trp trains departing NDP for FSB.
0936     A20 at CP S.  Waiting to escort convoy.
1005     A20 at CP18 picking up convoy.
1039     A Trp requests a LFT VR on area west of road vic Little Rubber at about 1330 hrs.
1245     A Trp rpts that in the village at XT360312 the people have been forced to carry ammo and cart
           wounded from the battle at the ARVN outpost last night (XT433272).
1519     A Trp trains departing FSB for NDP.
1615     A Trp NDP at XT455266 and platoon strong point at CP B.
Feb 16
Summary: M551 crews train in Cu Chi.  A Trp sweeps & outposts MSR.  A Trp escorts convoy.
0030     A Trp has a negative SITREP.
0552     A Trp stand-to complete.
0749     A Trp departing NDP.
0824     SPOTREP: LA A13; LB 30 paper flags and 2 banners; LC At XT466240 at 0810 hrs; LD policing up;
           LE continuing mission.
0836     A Trp trains on the move.
0925     A10 has completed sweep and is outposting.
1000     A Trp elements have picked up lead of the convoy.
1749     A Trp has closed on FSB Hampton.
2330     A Trp has a negative SITREP.
Feb 17
Summary: M551 crews train in Cu Chi.  A Troop conducts RIF in southern Boi Loi Woods.
0020     A Trp has a negative SITREP.
0540     A Trp stand-to is complete.
0712     A20 is on the move.
0715     A30 is on the move.
0817     A10 is on the move.
0857     A10 is at CP18 with HQ7 & VTR departing for B Trp NDP.
0919     VTR has been delivered to B Trp.
1030     A Trp starting RIF.
1039     A Trp at XT475294 heading on a 160 degree azimuth.
1141     A Trp at XT477268.
1210     A Trp at XT462276.
2333     A Trp has a negative SITREP.
Feb 18
Summary: M551 crews train in Cu Chi. A Trp sweeps MSR. A Troop conducts RIF in southern Boi Loi Woods.
Establishes NDP and a platoon strongpoint.
0013     A Trp has a negative SITREP.
0545     A Trp stand-to is complete.
0710     A Trp beginning sweep.
0730     Road swept from CP07 to CP05.
0750     Road swept from CP05 to CP22.
0933     A Trp on the move to today's operation at XT4628, XT4728, & XT4729.
1003     A Trp at XT474289.
1150     SPOTREP: LA A20; LB graves; LC XT475288 at 1100 hrs; LD 2 old graves & 1 fresh one. One with
           a 5 star symbol; LE checking them out.
1217     A Trp at XT475272, on the move.
1251     SPOTREP: LA A Trp; 1 Chicom grenade, 1 butterfly bomb; LC XT474265 at 1220 hrs; LD 9 holes
           in L shape about 1 or 2 months old; LE Engrs will blow in place.
1305     A Trp completed search of area, found 4 old buners and L shaped position.  No recent activity.
1326     Ref bunkers, also had pieces of paper with names and numbers on them.  Will bring to FSB.
1418     SPOTREP: LA A Trp; LB possible base camp; LC XT472258 at 1350 hrs; LD 4 bunkers, fighting
           fighting positions, trench line, 4 rice pots, 2 ponchos, indications of booby traps; LE signs of use in
           last 5 days.
1447     SPOTREP: LA A Trp; LB punji pit and tunnel; LC XT472252 at 1445 hrs; LD old tunnel w/4 punji pits;
           LE blowing in place.
1331     A Trp trains at CP18.  A Trp has relived C Trp of the road at CP B.
1336     A Trp trains at NDP at XT458265.
1620     A Trp(-) NDP at XT455266 and A10 platoon strongpoint at XT458253.
2030     A Trp radar status: R 58 mils, L 1800 mils, at 500 meters.
2100     SPOTREP: LA A20; LB fires; LC XT455285 at 2100  hrs; LD observed ARVN patrol that popped a
           star cluster which started the fire; LE continuing to observe.
2324     A Trp has a negative SITREP.
Feb 19
Summary: M551 crews train in Cu Chi. A Trp sweeps & outposts MSR. Escorts convoy. Establishes NDP.
0020     A Trp has a negative SITREP.
0600     A Trp stand-to is complete.
0702     A30 is on the move.
0822     A Trp cleared MSR from CP18 to CP Z.
0828     A Trp trains departing NDP.
0829     A10 outposting at CP R.
0853     SPOTREP: LA A14; LB bamboo detonator; LC XT458254 at 0800 hrs; LD observed on side of road;
           LE blown in place, continuing mission.
0909     A Trp trains entering FSB.
0930     A Trp elements escorting convoy.
1515     A Trp resupply departed FSB Hampton enroute to XT456266.
1545     A Trp resupply at NDP.
2105     A Trp radar status: L 5800 mils, C 700 mils, R 1600 mils at 500 meters.
2300     A Trp has a negative SITREP.
Feb 20
Summary: M551 crews train in Cu Chi. A Trp sweeps & outposts MSR. Escorts convoy. Establishes NDP and a platoon strong point.
0030     A Trp has a negative SITREP.
0540     A Trp stand-to complete.
0710     A20 sweeps south from XT459252.  A10 sweeps north from XT455267.
0824     A & B Trps cleared MSR from CP07 to CP22.
0837     A Trp trains departing NDP.
0926     A Trp elements escorting convoy.
1605     A Trp NDP at XT 459252 & 1 element at XT455265.
2353     A Trp has a negative SITREP.
Feb 21
Summary: M551 crews train in Cu Chi. A Trp sweeps & outposts MSR. Escorts convoy. Establishes NDP.
0035     A Trp has a negative SITREP
0535     A Trp stand-to is complete.
0700     A Trp beginning sweep.
1032     A Trp trains on the move.
1603     A Trp rpts that a man in the convoy driving vehicle #49 BP 155 #6 was drunk & running
           Vietnamese people off of the road.
2045     A Trp radar status: 4000 mils, 2400 mils, at 7500 meters.
2225     A Trp perimeter negative SITREP.
Feb 22
Summary: M551 crews train in Cu Chi. A Trp sweeps & outposts MSR. Escorts convoy. Establishes NDP and a platoon strong point.  Battery B 3rd Bn 13th Artillery in DS of 3/4 Cav moves to FSB Woods II to support 3/4 Cav operations in the Boi Loi Woods.
0030     A Trp has a negative SITREP.
0645     A20 is on the move.
0707     A10 on the move to CP22.
0732     A10 at XT470323.
0740     A Trp trains departing NDP
0803     A Trp trains at FSB.
0806     A20 outposting at CP07.
0945     A10 outposting from XT459252 to XT488204.  A30 will escort convoy.
1520     A Trp trains departing FSB.
1555     A Trp(-) NDP at XT452262.  A20 platoon strong point at CP B.
1950     A Trp radar status:  1400 mils, 4500 mils, 2950 mils at 500 meters.
2020     A Trp LPs:  Charlotte at XT459265, Judy at XT452258.
2105     A Trp rpts LP Judy in position.
2230     A Trp has a negative SITREP.